### v1

# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

[…]
Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design
[…]
siebel@math.uni-heidelberg.de
Heidelberg University
{adjustwidth}
50pt50pt We consider the statistical inverse problem of recovering a parameter $\theta\in H^{\alpha}$ from data arising from the Gaussian regression problem
$Y=\mathscr{G}(\theta)(Z)+\varepsilon$
with nonlinear forward map $\mathscr{G}:\mathbb{L}^{2}\to\mathbb{L}^{2}$, random design points $Z$ and Gaussian noise $\varepsilon$. The estimation strategy is based on a least squares approach under $\left\|\cdot\right\|_{H^{\alpha}}$-constraints. We establish the existence of a least squares estimator $\hat{\theta}$ as a maximizer for a given functional under Lipschitz-type assumptions on the forward map $\mathscr{G}$. A general concentration result is shown, which is used to prove consistency and upper bounds for the prediction error. The corresponding rates of convergence reflect not only the smoothness of the parameter of interest but also the ill-posedness of the underlying inverse problem. We apply the general model to the Darcy problem, where the recovery of an unknown coefficient function $f$ of a PDE is of interest. For this example, we also provide corresponding rates of convergence for the prediction and estimation errors. Additionally, we briefly discuss the applicability of the general model to other problems.
1   1Introduction
Overview: In this paper, we consider a nonlinear Gaussian regression model with random design, i.e., for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma\in\mathbb{R}_{>0}$, we have access to data $(Y_{i},X_{i})_{i=1}^{N}$ arising from
$Y_{i}=u_{f}(X_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.1)
where $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim N(0,1)$. Here, $X_{1},\dots,X_{N}$ are i.i.d. copies of $X$, which is drawn independently of $\varepsilon$ and follows the uniform distribution over a bounded domain $\mathcal{O}\subseteq\mathbb{R}^{d}$ with smooth boundary $\partial\mathcal{O}$. The regression function $u_{f}:\mathcal{O}\to\mathbb{R}$ denotes the solution of an (elliptic) partial differential equation (PDE), which depends nonlinearly on an unknown coefficient function $f$. The goal is to present an estimation strategy for $f$ by using the data set $(Y_{1},X_{1}),\dots,(Y_{N},X_{N})$, which is an example of a nonlinear statistical inverse problem as $f$ is not observed directly, see, e.g. Nickl [2023].
Many real-world phenomena, like physical and biological laws, fit into this framework as they are often described by (non-)linear dynamical systems and partial differential equations which exhibit spatial and/or temporal dependencies. Examples range from the description of cell movements (see Tokihiro [2021]) and geophysical laws describing the atmosphere (see Kalnay [2002]), to ocean dynamics (see Bennett [2002]), turbulence (see Majda and Harlim [2012]), and fluid flows (see Cotter et al. [2009]). Moreover, this setting aligns with various frameworks from the perspective of inverse problems: examples include electrical impedance tomography (see Calderón [2006] and Isaacson et al. [2004]), photoacoustic tomography and other hybrid imaging problems (see Kuchment and Kunyansky [2015], Bal [2014], and particularly Bal and Uhlmann [2010] for a general framework), inverse scattering (see Colton and Kress [2019] and Hohage and Weidling [2015]), interacting particle models (see Nickl et al. [2024]), and reaction diffusion equations (see Nickl [2024]). In this work, we consider the so-called Darcy Problem (see, for instance Stuart [2010]), where we refer to Bonito et al. [2017] and Hohage and Pricop [2008], and the references therein, for concrete application examples. In this situation $u_{f}:\mathcal{O}\to\mathbb{R}$ describes a steady-state heat flow, i.e. the unique solution to the elliptic PDE
$\begin{dcases}\hfill\operatorname{div}(f\cdot\nabla u)=g,\text{ on $\mathcal{O}$},\\ \hfill u=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (1.2)
with $\nabla$ denoting the gradient of a function and $\operatorname{div}$ the divergence operator applied to a vector field, respectively. Here $g>0$ is a heat source function on $\mathcal{O}$, which is known in advance, and $f:\mathcal{O}\to(0,\infty)$ describes an unknown conductivity (or diffusion) coefficient function. In real-world observation schemes, idealized observations that exactly match the mathematical formalism of a PDE are typically unavailable, as measurement errors introduce statistical noise, motivating the observation scheme Eq. 1.1 in this work. More precisely, the data consists of two parts - the exact solution $u_{f}$ of the Darcy problem Eq. 1.2, corrupted by the second part - the Gaussian noise. Nonetheless, the goal remains the same: to reconstruct the underlying unknown coefficient function $f$ that fits the pattern of Eq. 1.2.
In the present work, we follow the approaches in Monard et al. [2020] and Nickl et al. [2020] and embed the regression model Eq. 1.1 into a more generic setting. Specifically, given a probability space $(\mathcal{Z},\mathscr{Z},\zeta)$ and a possibly nonlinear operator (forward map) $\mathscr{G}:\mathbb{L}^{2}(\mathcal{O})\to\mathbb{L}_{\zeta}^{2}(\mathcal{Z})$, we assume to have access to observations $(Y_{i},Z_{i})_{i=1}^{N}$, which arise from the equation
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.3)
where, in this situation, $Z_{1},\dots,Z_{N}$ are i.i.d. drawn from $\zeta$. The goal is now to provide an estimation strategy for $\theta$, which defines an object in a Sobolev space $H^{\alpha}(\mathcal{O})$. Note that Eq. 1.1 corresponds to the special case where $\mathscr{G}$ is the nonlinear solution map $f\mapsto u_{f}$ of the Darcy problem. The reconstruction of parameters $\theta$ following an indirect and noisy observation scheme as in Eq. 1.3 has a long mathematical history. For an extensive overview of models with deterministic noise and a linear relationship $\theta\mapsto\mathscr{G}(\theta)$, we refer to Engl et al. [1996]. In the last twenty years a wide field has been established to also consider nonlinear and statistical inverse problems. We refer to Benning and Burger [2018], Kaltenbacher et al. [2008], Dashti et al. [2013], Dashti and Stuart [2015], Stuart [2010], Nickl [2023], Arridge et al. [2019], Bissantz et al. [2007], and Kaipio and Somersalo [2005], as well as to the references therein. For a novel proof in the Bayesian approach within this PDE-constrained regression setting, we refer to Monard et al. [2020] and Monard et al. [2021], where statistical guarantees, especially for the posterior mean, are derived under Gaussian process priors on the parameter of interest $\theta$ and analytical assumptions on the forward map $\mathscr{G}$. The applicability to PDE models is illustrated not only for the inversion of non-Abelian X-ray transforms but also for a model as in Eq. 1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the stationary Schrödinger equation. Giordano and Nickl [2020] adopt these ideas for the Darcy problem Eq. 1.2 and provide results on posterior convergence. General assumptions to show posterior convergence are formulated in Nickl [2023]. In Nickl et al. [2020], the authors propose a (generalized) penalized least squares estimator $\hat{\theta}$ for $\theta$ in a Gaussian white noise model, which can be interpreted as minimizer of the quantity
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\left\|Y-\mathscr{G}(\theta)\right\|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|^{2}_{H^{\alpha}(\mathcal{O})}\in\mathbb{R}_{\geq 0}.$
They make use of M-estimation techniques to derive statistical guarantees for its prediction loss. The results are then applied for the special case of considering Eq. 1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the Darcy problem Eq. 1.2 and also for the time-independent Schrödinger equation. In particular, they provide upper bounds for the prediction loss and show that these are in some situations minimax-optimal and coincide with rates known for linear operators. For the proof, the mathematical formulation of the observation scheme via a white noise process is very suitable, however, this scheme is rather infeasible for real-world applications. Furthermore, the authors discuss the relationship between $\hat{\theta}$ and the Bayesian approach, where $\hat{\theta}$ can also be interpreted as maximizer of the likelihood $\pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$ of the posterior distribution $\Pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$, under a Gaussian process priors $\Pi(\cdot)$ on $\theta$. This also results in the naming of $\hat{\theta}$ as maximum a posteriori (MAP) estimator. In contrast, we aim to prove results in the more realistic i.i.d. regression model Eq. 1.3.
For random design regression models described by Eq. 1.3, Bohr and Nickl [2023] and Nickl and Wang [2024] provide statistical and computational guarantees for a version of the maximum a posteriori estimator (MAP) $\hat{\theta}$, defined as the $D$-dimensional minimizer of the quantity
$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\frac{1}{N}\sum_{i=1}^{N}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}^{2}\in\mathbb{R}_{>0}.$ (1.4)
Under fairly general assumptions on the forward map $\mathscr{G}$, the authors demonstrate the consistency and stability of $\hat{\theta}$, provided that $\theta$ can be sufficiently well approximated by its orthogonal projection onto a finite-dimensional subspace. The key of the proof is an empirical process argument, which implicitly requires the forward map $\mathscr{G}$ to satisfy the ’global’ stability estimates
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\mathbb{L}^{2}}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\mathbb{L}^{2}}\text{ and }\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty}$ (1.5)
uniformly over the whole parameter space. While the solution map of the time-independent Schrödinger equation meets these requirements, there is no known analogous result for the second condition for the solution map of the Darcy problem, which renders the theory of Bohr and Nickl [2023] and Nickl and Wang [2024] inapplicable. In addition, the first requirement leads to slower convergence rates, which we improve by taking the ill-posedness of the forward map $\mathscr{G}$ into account (see Nickl et al. [2020] for corresponding results in the white noise model). In section 2.2. of Nickl and Wang [2024], the authors propose conditions under which the stochastic gradient algorithm
$\vartheta_{0}:=\theta_{\operatorname{init}},\hskip 4.30554pt\vartheta_{k+1}=\vartheta_{k}+\rho\nabla\log\tilde{\pi}(\vartheta_{k}|(Y_{i},Z_{i})_{i=1}^{N})$ (1.6)
converges geometrically to the minimizer $\hat{\theta}$ of Eq. 1.4 (and hence to the ground truth $\theta$ after sufficiently many iterations) in polynomial time, if the initial value $\theta_{\operatorname{init}}\in\mathbb{R}^{D}$ is chosen appropriately. We refer to section 2.2.5. (and particularly Theorem 2.8) in Nickl and Wang [2024]. The results are then applied to $\mathscr{G}$ being the solution map $f\mapsto u_{f}$ of the stationary Schrödinger equation, which meets global stability conditions Eq. 1.5. Under more general assumptions, Nickl [2023] discusses situations in which the analysis of the local concavity of the associated log-likelihood to Eq. 1.3 is sufficient to specify a stochastic gradient algorithm, for which $\hat{\theta}$ is computable in polynomial time, if the underlying parameter space is finte-dimensional. Under slightly technical changes, the methods also seem to apply to the Darcy problem. However, we refrain from discussing the computational aspects in more detail in this work due to the similarities with Nickl [2023] and Nickl and Wang [2024].
A different approach of modeling real-world phenomena utilizes the description of dynamical (time-dependent) systems via stochastic partial differential equations (SPDEs), which inherently take into account the noise in the system dynamics. The SPDE framework describes uncertainty by embedding stochastic processes directly into the model equations, thereby offering an alternative for capturing the evolution of systems influenced by random fluctuations. For an overview of recent work of SPDEs we refer to Altmeyer et al. [2023], Altmeyer and Reiß [2021] and Gaudlitz and Reiß [2023] as well as the references therein. We emphasize that SPDEs are not the focus of the present work.
In the present paper, we follow the approach of Monard et al. [2020], Monard et al. [2021] and Nickl et al. [2020] and consider the general nonlinear Gaussian regression model with random design (as Eq. 1.3) and define a maximum a posteriori estimator, similar to Eq. 1.4. We introduce conditions on the forward map $\mathscr{G}$, generalizing Eq. 1.5 by replacing the right-hand sides in both stability estimates with $(H^{\kappa})^{*}$- and $C^{\eta}$-norms, respectively, while also allowing polynomial growth. We present a general concentration result, where we compare the prediction loss $\mathfrak{d}^{2}(\hat{\theta},\theta)$ with some proxy $\mathfrak{d}^{2}(\tilde{\theta},\theta)$ (interpretable as approximation error) and obtain a corresponding consistency result for the MAP estimate, whenever the proxy approximates the ground truth sufficiently well. This result extends the findings of Bohr and Nickl [2023] and Nickl and Wang [2024], as we allow a generic reference function $\tilde{\theta}$, rather than only a finite-dimensional approximation of the ground truth and provide faster convergence rates for the squared prediction loss, which are now upper-bounded by $N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}$, where $\kappa\in\mathbb{R}_{>0}$ indicates the degree of ill-posedness of the underlying inverse problem. Under an additional inverse stability assumption on the forward map $\mathscr{G}$, an analogous result for the estimation loss of $\hat{\theta}$ is established. For the sake of completeness, we provide an existence result for the MAP-estimator in this random design regression model by exploiting the direct method in the calculus of variations, where we adapt a corresponding proof of Nickl et al. [2020] for the white noise model. The general theory is applied on the Darcy problem. Specifically, we demonstrate that the corresponding solution map satisfies the new stability assumption with $\eta=1$. In this particular situation, we provide consistency for the MAP estimator, as well as corresponding convergence rate for prediction and estimation errors, which fills a gap in existing theory.
Organisation of the paper: In Section 2 we introduce the most essential notation and the statistical model in detail. Further, we introduce and discuss the assumption made on the forward map $\mathscr{G}$. We define the MAP-estimator, for which we provide an existence result. The main results - consisting of the concentration inequality, an upper bound for the prediction error, as well as consistency for the MAP-estimator - are presented. In Section 3 we apply the main result to the special case of the Darcy problem. We provide analogously consistency for the MAP-estimator, as well as upper bounds for the prediction and estimation risk. In Appendix A, we provide detailed proofs for the results in the main part and further helping results.
Acknowledgement: This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1-39090098 (the Heidelberg STRUCTURES Cluster of Excellence). The author thanks Richard Nickl and Jan Johannes for their time, patience, and fruitful discussions, as well as for the visit to Richard Nickl’s research group in Cambridge in November 2023, which marked the beginning of this work.
2   2A general statistical inverse Problem
Notation § 2.1 (Preliminary).
We start with some basic notation that we use throughout the whole paper.
  1. i)
    For all $\mathbb{K}\in\{\mathbb{N},\mathbb{R}\}$ and $a\in\mathbb{R}$ set $\mathbb{K}_{\geq a}:=\mathbb{K}\cap[a,\infty)$ and $\mathbb{K}_{>a}:=\mathbb{K}\cap(a,\infty)$. Analogously, $\mathbb{K}_{\leq a}$ and $\mathbb{K}_{<a}$ are defined. In particular, we set $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
  2. ii)
    For all $a,b\in\mathbb{Z}$ with $a\leq b$ set $\left\llbracket a,b\right\rrbracket:=[a,b]\cap\mathbb{Z}$ and further $\left\llbracket n\right\rrbracket:=\left\llbracket 1,n\right\rrbracket$ for any $n\in\mathbb{N}$.
  3. iii)
    For all $a,b\in\mathbb{R}$ define $a\lor b:=\max\{a,b\}$ and $a\land b:=\min\{a,b\}$.
  4. iv)
    For a topological space $(S,\mathcal{T})$ we denote by $\mathscr{B}_{S}$ the Borel-$\sigma$-field of $S$ generated by the topology $\mathcal{T}$. We write $S^{d}:=\bigtimes_{i\in\left\llbracket d\right\rrbracket}S$ for some $d\in\mathbb{N}$ and denote by $\mathscr{B}^{\otimes d}_{S}:=\bigotimes_{i\in\left\llbracket d\right\rrbracket}\mathscr{B}_{S}$ the product $\sigma$-algebra on $S^{d}$. The topological dual of $S$ is denoted by $S^{*}$ and consists of all linear and bounded functionals $L:S\to\mathbb{R}$. If $S$ is equipped with a metric $d_{S}$, we write $\mathbb{B}_{S}(s_{o},r):=\{s\in S:\,d_{S}(s,s_{o})\leq r\}$ for the closed ball of radius $r\in\mathbb{R}_{>0}$ around $s_{o}\in S$ in $S$.
  5. v)
    For two normed spaces $(S_{1},\left\|\cdot\right\|_{S_{1}})$ and $(S_{2},\left\|\cdot\right\|_{S_{2}})$, we write $S_{1}\hookrightarrow S_{2}$, if $S_{1}\subseteq S_{2}$ is continuously embedded into $S_{2}$.
  6. vi)
    For $d\in\mathbb{N}$ denote by $\left\|\cdot\right\|_{\mathbb{R}^{d}}$ the usual Euclidean norm on $\mathbb{R}^{d}$.
  7. vii)
    For a function $f:\mathcal{S}\to\mathbb{R}$ defined on an arbitrary space $\mathcal{S}$, we denote by $\left\|f\right\|_{\infty}:=\sup_{s\in\mathcal{S}}\left\{|f(s)|\right\}$ the supremum norm of $f$ over $\mathcal{S}$. We do not highlight the underlying space, if it is clear from the context.
  8. viii)
    Universal constants are denoted by $c$. If a constant depends on a family of objects $A$, we write $c(A)$. Constants arising from model assumptions are denoted by $C$ with some further specification. In each statement and also in the corresponding proofs, we will number constants consecutively, always starting anew. If not further mentioned, the value of a constant can change from line to line.
   Hölder and Sobolev spaces
Throughout let $(\mathcal{Z},\mathscr{Z},\zeta)$ be a measure space. For $p\in[1,\infty]$ we define the usual space $\mathbb{L}_{\zeta}^{p}(\mathcal{Z}):=\mathbb{L}^{p}(\mathcal{Z},\mathscr{Z},\zeta)$ of (equivalence classes of) functions $f:(\mathcal{Z},\mathscr{Z})\to(\mathbb{R},\mathscr{B}_{\mathbb{R}})$ endowed with the norm $\left\|\cdot\right\|_{\mathbb{L}^{p}_{\zeta}(\mathcal{Z})}$. Let $(V,\left|\cdot\right|_{V})$ be a finite-dimensional normed $\mathbb{R}$-vector space with dimension $d_{V}:=\operatorname{dim}_{\mathbb{R}}(V)$. By identifying $V$ canonically with $\mathbb{R}^{d_{V}}$, we define the space of all vector fields $f=(f_{1},\dots,f_{d_{V}}):\mathcal{Z}\to V$ with the integrable components $f_{i}\in\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$, $i\in\left\llbracket d_{V}\right\rrbracket$, as the tensor space $\mathbb{L}^{p}_{\zeta}(\mathcal{Z},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$. Recalling that $\mathbb{L}^{2}_{\zeta}(\mathcal{Z})$ defines a Hilbert space with inner product
$\langle f,g\rangle_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}:=\int_{\mathcal{Z}}f(x)g(x)\mathrm{d}\zeta(x)$
for $f,g\in\mathbb{L}^{2}_{\zeta}(\mathcal{Z})$, the space $\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ also inherits a Hilbert space structure with an inner product given by
$\langle\cdot,\cdot\rangle_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})},$
where $\cdot_{i}$ indicates the $i$-th component. Throughout let $d\in\mathbb{N}$. We denote by $\mathcal{O}\subseteq\mathbb{R}^{d}$ a bounded domain with smooth boundary $\partial\mathcal{O}$. For any $s\in\mathbb{N}$ denote by $C^{s}(\mathcal{O})$ the space of $s$-times continuously differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with bounded derivatives. By convention, we denote by $C^{0}(\mathcal{O})$ the space of continuous functions on $\mathcal{O}$. We extend the definition for any $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by defining $C^{s}(\mathcal{O})$ as the function space of all maps $f:\mathcal{O}\to\mathbb{R}$, such that for all multi-indices $\bm{\beta}\in\mathbb{N}_{0}^{d}$ with $|\bm{\beta}|\leq\lfloor s\rfloor$ the integer part of $s$ the derivative $D^{\bm{\beta}}f$ exists and is $s-\lfloor s\rfloor$-Hölder continuous. A norm on $C^{s}(\mathcal{O})$ is given by
$\left\|f\right\|_{C^{s}(\mathcal{O})}:=\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|\leq\lfloor s\rfloor}\left\|D^{\bm{\beta}}f\right\|_{\infty}+\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|=\lfloor s\rfloor}\sup_{x,y\in\mathcal{O}:x\neq y}\left\{\frac{|D^{\bm{\beta}}f(x)-D^{\bm{\beta}}f(y)|}{\left\|x-y\right\|_{\mathbb{R}^{d}}^{s-\lfloor s\rfloor}}\right\}.$
A function $f:\mathcal{O}\to\mathbb{R}$ is called smooth, if it belongs to $C^{\infty}(\mathcal{O}):=\bigcap_{s\in\mathbb{R}_{>0}}C^{s}(\mathcal{O})$. We further denote by $C^{\infty}_{c}(\mathcal{O})$ its subspace of all smooth functions having compact support. Further let $\mathcal{L}^{d}$ be the usual $d$-dimensional Lebesgue measure on $(\mathbb{R}^{d},\mathscr{B}_{\mathbb{R}}^{\otimes d})$ and denote by $\mathcal{L}_{\mathcal{O}}^{d}$ its restriction on $\mathcal{O}$. To shorten notation, set $\mathbb{L}^{2}(\mathcal{O}):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O})$ and analogously $\mathbb{L}^{2}(\mathcal{O},V):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O},V)$. For $s\in\mathbb{N}$, we denote by $H^{s}(\mathcal{O}):=H^{s}(\mathcal{O},\mathbb{R})$ the usual order-$s$ Sobolev space of weakly differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with square integrable weak derivatives up to order $s$. By $\left\|\cdot\right\|_{H^{s}(\mathcal{O})}$, we denote its usual norm. This definition extends to positive $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by interpolation, see Triebel [1983]. By convention, set $H^{0}(\mathcal{O}):=\mathbb{L}^{2}(\mathcal{O})$. Further define $H^{s}_{c}(\mathcal{O}):=\overline{C_{c}^{\infty}(\mathcal{O})}^{\left\|\cdot\right\|_{H^{s}(\mathcal{O})}}$ as well as $\bar{H}^{s}(\mathcal{O})$ to be the space of all functions in $H^{s}(\mathcal{O})$, whose support is $\bar{\mathcal{O}}$. If $s\neq k+\frac{1}{2}$ for some $k\in\mathbb{N}$, we have $\bar{H}^{s}(\mathcal{O})=H^{s}_{c}(\mathcal{O})$. Negative-order Sobolev spaces are defined as the topological dual of $\bar{H}^{s}(\mathcal{O})$, i.e., $H^{-s}(\mathcal{O}):=H^{-s}(\mathcal{O},\mathbb{R}):=\left(\bar{H}^{s}(\mathcal{O})\right)^{*}$. Denoting by $\operatorname{tr}_{s}:H^{s}(\mathcal{O})\to\bigtimes_{l=0}^{s-1}H^{s-l-\frac{1}{2}}(\mathcal{O})$ the higher-order trace operator for any $s\in\mathbb{N}$, we have $H^{s}_{0}(\mathcal{O}):=\left\{f\in H^{s}(\mathcal{O}):\,\operatorname{tr}_{s}[f]=0\right\}=H^{s}_{c}(\mathcal{O})$, where we refer to Lions and Magenes [1972] for detailed definitions and properties. All of the previous spaces of functions defined on $\mathcal{O}$ with values in $\mathbb{R}$ can be extended to spaces of the corresponding vector fields $f:\mathcal{O}\to V$ in the same manner as before. Indeed if $F(\mathcal{O})$ is one of the above function spaces, we set $F(\mathcal{O},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}F(\mathcal{O})$. If $F(\mathcal{O})$ is equipped with some norm $\left\|\cdot\right\|_{F(\mathcal{O})}$, a norm on $F(\mathcal{O},V)$ is given by $\left\|\cdot\right\|_{F(\mathcal{O},V)}^{2}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\left\|\cdot_{i}\right\|_{F(\mathcal{O})}^{2}$. Similarly, if $F(\mathcal{O})$ is endowed with a Hilbert space inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O})}$, $F(\mathcal{O},V)$ inherits the Hilbert space structure with the inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O},V)}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{F(\mathcal{O})}$. Further, for simplicity, we write for centered balls with radius $r\in\mathbb{R}_{>0}$ in function spaces $F(\mathcal{O})$ and $F(\mathcal{O},V)$ $F(\mathcal{O},r):=\mathbb{B}_{F(\mathcal{O})}(0,r)$ and $F(\mathcal{O},V,r):=\mathbb{B}_{F(\mathcal{O},V)}(0,r)$, respectively.
   The measurement model
Throughout let $(W,\left|\cdot\right|_{W})$ be another finite-dimensional and normed $\mathbb{R}$-vector space with dimension $d_{W}:=\operatorname{dim}_{\mathbb{R}}(W)\in\mathbb{N}$, such that we canonically identify $W$ with $\mathbb{R}^{d_{W}}$. Assuming that $(\mathcal{Z},\mathscr{Z},\zeta)$ is a probability space, we are interested in a (possibly nonlinear) forward map
$\mathscr{G}:\tilde{\Theta}\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V),$
where $\tilde{\Theta}\subseteq\mathbb{L}^{2}(\mathcal{O},W)$ denotes a parameter space. We emphasize that in this paper we are only concerned with forward maps $\mathscr{G}$ that have for each $\theta\in\tilde{\Theta}$ a unique continuous representative in $\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$, so that point evaluations of the form $\mathscr{G}(\theta)(z)$ are well-defined for each $z\in\mathcal{Z}$. We assume to have access to measurements of a random design Gaussian regression model, i.e. for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma^{2}\in\mathbb{R}_{>0}$, we observe $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from the equation
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i\in\left\llbracket N\right\rrbracket,\,\theta\in\tilde{\Theta},$ (2.1)
where $Z_{1},\dots,Z_{N}$ are i.i.d copies of $Z$, which is drawn from $\zeta$. The error terms $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim\operatorname{N}(0,\operatorname{Id}_{V})$ with $\operatorname{Id}_{V}$ denoting the identity operator on the vector space $V$. The design variable $Z$ and the error $\varepsilon$ are assumed to be independent. The distribution of a generic observation $D:=(Y,Z)$ arising from Eq. 2.1 is denoted by $\mathbb{P}_{\theta}$. The joint product law of $D_{N}$ on $(V\times\mathcal{Z})^{N}$ is denoted as $\mathbb{P}_{\theta}^{N}=\bigotimes_{i\in\left\llbracket N\right\rrbracket}\mathbb{P}_{\theta}$. The corresponding expectation operators are denoted by $\mathbb{E}_{\theta}$ and $\mathbb{E}_{\theta}^{N}$, respectively. Keeping the identification of $V$ with $\mathbb{R}^{d_{V}}$ in mind, $\mathbb{P}_{\theta}$ has a probability density function $p_{\theta}$ with respect to the product measure $\mathcal{L}^{d_{V}}\otimes\zeta$, namely
$\forall y\in V,z\in\mathcal{Z}:\,p_{\theta}(y,z)=(2\pi\sigma^{2})^{-\frac{d_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\left|y-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right).$
Consequently, $\mathbb{P}_{\theta}^{N}$ also has a probability density function with respect to $\bigotimes_{i\in\left\llbracket N\right\rrbracket}(\mathcal{L}^{d_{V}}\otimes\zeta)$ denoted by $p_{\theta}^{N}$, which reads for all $(y,z)\in(V\times\mathcal{Z})^{N}$ as
$p_{\theta}^{N}(y,z)=\prod_{i\in\left\llbracket N\right\rrbracket}p_{\theta}(y_{i},z_{i})=(2\pi\sigma^{2})^{-\frac{Nd_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|y_{i}-\mathscr{G}(\theta)(z_{i})\right|_{V}^{2}\right).$
Given the data vector $D_{N}$, a shifted log-likelihood function $\ell_{N}:\tilde{\Theta}\to\mathbb{R}$ is defined by
$\ell_{N}(\theta):=\ell_{N}(\theta|D_{N}):=\log p_{\theta}^{N}(D_{N})+\frac{Nd_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
for each $\theta\in\tilde{\Theta}$. Analogously, given a single datum $D=(Y,Z)$, we define
$\ell(\theta):=\ell(\theta|D):=\log p_{\theta}(D)+\frac{d_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\left|Y-\mathscr{G}(\theta)(Z)\right|_{V}^{2}.$
The results obtained below hold under the frequentist assumption that the data $D_{N}$ from Eq. 2.1 is generated from a fixed ground truth $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ including the law $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$.
   Maximum a posteriori estimator
We are now introducing the maximum a posteriori estimator, whose statistical guarantees are presented in the following sections. Given the data set $D_{N}=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from Eq. 2.1, fix some $\alpha,\mathfrak{r}\in\mathbb{R}_{\geq 0}$. We define a penalized least squares functional by
$\mathscr{J}_{\mathfrak{r},N}:\tilde{\Theta}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\text{ }\mathscr{J}_{\mathfrak{r},N}[\theta]:=-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$ (2.2)
and set
$\hat{\theta}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right\}.$ (2.3)
We call $\hat{\theta}_{\textsc{MAP}}$ maximum a posteriori estimator, or shortly MAP estimator.
Remark § 2.2 (Maximum a posteriori estimator).
  1. i)
    Up to now, we do not have formulated any specific conditions on $\tilde{\Theta}$, except of $\tilde{\Theta}\subseteq\mathbb{L}^{2}(\mathcal{O},W)$. In Proposition 2.5, a restriction of $\tilde{\Theta}$ is neccessary in order to provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer of the optimization problem Eq. 2.3. However, the results of the statistical analysis hold true for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$. From the application perspective, the definition of $\hat{\theta}_{\textsc{MAP}}$ seems unfeasible as the optimization problem is formulated over a possibly infinite-dimensional space. The theory shown in Nickl [2023] and Nickl and Wang [2024] provide general conditions and algorithms, under which $\hat{\theta}_{\textsc{MAP}}$ can be computed in polynomial time, if $\tilde{\Theta}$ is finite dimensional.
  2. ii)
    The MAP estimator $\hat{\theta}_{\textsc{MAP}}$ introduced in Eq. 2.3 has two different interpretations: On the one hand, $\hat{\theta}_{\textsc{MAP}}$ is strongly related to the classical theory of Tikhonov-regularized estimators - we refer to Engl et al. [1996] for a detailed analysis of the regularization of ill-posed deterministic inverse problems. Indeed, $\hat{\theta}_{\textsc{MAP}}$ minimizes the empirical squared distance. To exlude unnecessary complex solutions of the problem, the regularity is penalized in terms of the $H^{\alpha}(\mathcal{O},W)$-norm. On the other hand, $\hat{\theta}_{\textsc{MAP}}$ corresponds to the framework of Bayesian inverse problem techniques. Under Gaussian process priors on the parameter space $\tilde{\Theta}$, the corresponding likelihood of the a posteriori distribution is then maximised by $\hat{\theta}_{\textsc{MAP}}$. We refer here to the comments in Nickl et al. [2020] and Bohr and Nickl [2023], and further to Dashti et al. [2013], Helin and Burger [2015] for more details.
   A convergence rate result
The main results of this work - Proposition 2.5, Theorem 2.6 and Corollary 2.7 - are presented in the following. Firtsly, we provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a solution to the optimization problem Eq. 2.3 over a suitable subset $\Theta\subseteq\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$. Moreover, we are showing a concentration result controlling the prediction error and the regularity of $\hat{\theta}_{\textsc{MAP}}$ simultanously under $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability. Therefore, we define for every $\theta_{1}\in\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$, $\theta_{2}\in\tilde{\Theta}$, and $\mathfrak{r}\in\mathbb{R}_{\geq 0}$ the functional
$\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}):=\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
We start with analytical assumption on the forward map $\mathscr{G}$.
Condition § 2.3 (Forward Map I).
  1. C1)
    $\mathbb{L}^{2}$-Lipschitz continuity: Let $\alpha,\gamma_{1},\kappa\in\mathbb{R}_{\geq 0}$. Set $\mathscr{H}=H^{\alpha}_{c}(\mathcal{O},W),\text{ if $\kappa\geq\frac{1}{2}$ }$ and $\mathscr{H}=H^{\alpha}(\mathcal{O},W),\text{ if $\kappa<\frac{1}{2}$}$. The map $\mathscr{G}$ is $(\alpha,\gamma_{1},\kappa)$-regular in $\mathbb{L}^{2}$-norm, if there exists a constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in\tilde{\Theta}\cap\mathscr{H}$
    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
  2. C2)
    Uniform boundedness: There exists a constant $C_{\mathscr{G},\operatorname{B}}\in\mathbb{R}_{\geq 1}$, such that
    $\sup_{\theta\in\tilde{\Theta},z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq C_{\mathscr{G},\operatorname{B}}.$
  3. C3)
    $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha,\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, such that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. The map $\mathscr{G}$ is $(\alpha,\gamma_{2},\kappa)$-regular in $\mathbb{L}^{\infty}$-norm, if there exists a constant $C_{\operatorname{Lip},\infty}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O},W)$
    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Remark § 2.4 ($\mathbb{L}^{\infty}$-Lipschitz continuity).
The effect that stronger assumptions on the statistical model are needed to prove consistency for the maximum likelihood estimator than to show corresponding concentration results in the Bayesian setting (see Le Cam [1986]) is also apparent in the nonlinear random regression setting considered here: While in Nickl [2023] it suffices to assume condition C1) and C2) (as special case with $\gamma_{1}=0$) to obtain posterior convergence results, in Bohr and Nickl [2023] consistency of the MAP estimator necessitates a $\mathbb{L}^{\infty}$-stability estimate of the forward map $\mathscr{G}$, i.e.,
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty},$
uniformly over the whole parameter space. C3) in 2.3 generalizes this condition by replacing the uniform norm by the stronger $\left\|\cdot\right\|_{C^{\eta_{2}}}$-norm. Moreover, we now allow a polynomial increase of order $\gamma_{2}$. While the conditions formulated in Bohr and Nickl [2023] (which correspond to the special case $\eta_{1}=\eta_{2}=\gamma_{2}=0$ in the present setting!) are insufficient for the Darcy problem, we see in Section 3 and Appendix A that the associated solution map fulfills C3) with $\eta_{1}=\eta_{2}=1$ and $\gamma_{2}=4$. Note that the condition $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$ in C3) ensures that all norms are well-defined as in this situation the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{i}}(\mathcal{O},W)$, $i\in\left\llbracket 2\right\rrbracket$, holds true.
The first main results addresses the open question of existence of $\hat{\theta}_{\textsc{MAP}}$ as the maximizer of the penalized Tikhonov-regularized functional Eq. 2.2. For the white noise model, the authors in Nickl et al. [2020] provide a proof using the white noise formulation of the regression model. We adapt the corresponding proof to the random design regression model, also exploiting the Direct Method of the Calculus of Variations, making reference to Dacorogna [2008]. The proof can be found in Appendix A.
Proposition § 2.5 (Existence of $\hat{\theta}_{\textsc{MAP}}$).
Suppose that $\mathscr{G}:\tilde{\Theta}\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},W)$ satisfies the conditions C1) - C3) imposed in 2.3 for some $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$, let $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$ be closed in the weak topology of the Hilbert space $\mathscr{H}$. Then for all $\mathfrak{r}\in\mathbb{R}_{>0}$, almost surely under $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, there exists a maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, satisfying
$\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\mathscr{J}_{\mathfrak{r},N}[\hat{\theta}_{\textsc{MAP}}].$
Let $R,\mathfrak{r}\in\mathbb{R}_{>0}$ be fixed. To shorten notation, we introduce the event
$\Xi_{\mathfrak{r},R}:=\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}$
for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, some reference parameter $\tilde{\theta}\in\Theta$ (with $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$) and some $c_{1}\in\mathbb{R}_{\geq 1}$. To shorten notation, we define the family of symbols $\texttt{P}:=\left\{\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma\right\}$ and write $c(\texttt{P})$ for a constant to emphasize the dependency.
Theorem § 2.6 (Concentration inequality).
Suppose that $\mathscr{G}:\tilde{\Theta}\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfies C1)-C3) in 2.3 with some $\alpha>\max\left\{\eta_{1}\lor\eta_{2}+d\lor\frac{d}{2}(1+\gamma_{2}),\frac{d}{2}\gamma_{1}-\kappa\right\}$. Let $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ from Eq. 2.1 with some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$. Then the following hold.
  1. i)
    Let $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$. Then for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\bar{c},\texttt{P})\in\mathbb{R}_{>0}$ sufficiently large, such that for all $\mathfrak{r}\in\mathbb{R}_{>0}$ and $R\in\mathbb{R}_{>0}$, with $R\geq\mathfrak{r}\geq N^{-\frac{1}{2}}$, satisfying
    $R^{-1+\frac{d\gamma_{1}}{2(\alpha+\kappa)}}\mathfrak{r}^{-\frac{d(\gamma_{1}+1)}{2(\alpha+\kappa)}}\leq N^{\frac{1}{2}}$ (C4)
    and
    $R^{-2+\frac{d(\gamma_{2}+1)}{\alpha-\eta_{2}}}\mathfrak{r}^{-\frac{d(\gamma_{2}+1)}{\alpha-\eta_{2}}}\leq N,$ (C5)
    any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ and $\tilde{\theta}\in\Theta$ satisfies
    $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{2}\cdot\exp(-\bar{c}\cdot NR^{2})$ (2.4)
    with $c_{2}=c_{2}(d_{V})\in\mathbb{R}_{>0}$.
  2. ii)
    Let $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$. Then, there exists a constant $c_{3}=c_{3}(\texttt{P})\in\mathbb{R}_{>0}$, such that for all $\mathfrak{r}\in\mathbb{R}_{>0}$ satisfying
    $\mathfrak{r}^{-1}\lor\mathfrak{r}^{-1-\frac{d}{2(\alpha+\kappa)}}\leq N^{\frac{1}{2}},$ (C6)
    we have for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$
    $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}\left[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\right]\leq c_{3}\cdot\inf_{\tilde{\theta}\in\Theta}\left\{\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}(1+\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O})}^{2})+\frac{1}{N}\right\}.$ (2.5)
Corollary § 2.7 (Consistency).
Under the assumptions of Theorem 2.6, assume $\theta_{\mathfrak{o}}\in\Theta$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}\in\mathbb{R}_{>0}$. Let $\mathfrak{r}\in\mathbb{R}_{>0}$ be given by
$\mathfrak{r}:=\mathfrak{r}_{N}:=N^{-\frac{\alpha+\kappa}{2(\alpha+\kappa)+d}}.$
Then, for all $\bar{c}\in\mathbb{R}_{\geq 1}$, there exists $c_{4}=c_{4}(\bar{c},\operatorname{L},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{4}\mathfrak{r}^{2}_{N}\right\}\right)\leq c_{5}\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
with $c_{5}=c_{5}(d_{V})\in\mathbb{R}_{>0}$. In particular, there is $c_{6}=c_{6}(d_{V},\operatorname{L})\in\mathbb{R}_{>0}$, such that
$\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right]\right\}\leq\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta)]\right\}\leq c_{6}\cdot\mathfrak{r}_{N}^{2}.$
Remark § 2.8 (Consistency and prediction error).
  1. i)
    Note, that Theorem 2.6 does not require $\theta_{\mathfrak{o}}\in\Theta$ as long as it can be approximated sufficiently by some $\tilde{\theta}\in\Theta$. Therefore, the assumption $\theta_{\mathfrak{o}}\in\Theta$ in Corollary 2.7 can also be weakened to the effect that $\mathscr{G}(\tilde{\theta})$ approximates $\mathscr{G}(\theta_{\mathfrak{o}})$ in some sense (see equation 2.23 in Bohr and Nickl [2023] for instance).
  2. ii)
    Evidently, Corollary 2.7 provides the consistency of the MAP estimator $\hat{\theta}_{\textsc{MAP}}$. The rate of convergence for the prediction loss is known to be optimal for $\kappa$-smoothing linear forward maps $\mathscr{G}$.
Without any further assumptions, there is no hope to obtain an analogous upper bound for the estimation risk as well. However, the next corollary demonstrates stability estimates under a inverse continuity modulus of $\mathscr{G}^{-1}$.
Condition § 2.9 (Forward Map II).
  1. C7)
    Inverse continuity modulus: Additionally to 2.3, define for any $\delta\in\mathbb{R}_{>0}$ the set
    $\Lambda_{\delta}:=\left\{(\theta_{1},\theta_{2})\in\tilde{\Theta}^{2}:\,\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{M},\,\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq\delta\right\}$
    for some universal constant $\operatorname{M}\in\mathbb{R}_{>0}$. There exists a constant $C_{\operatorname{M}}\in\mathbb{R}_{>0}$ and some $0<\tau\leq 1$, such that for $\delta$ small enough
    $\sup\left\{\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}:\,(\theta,\theta^{\prime})\in\Lambda_{\delta}\right\}\leq C_{\operatorname{M}}\cdot\delta^{\tau}.$
Corollary § 2.10 (Stability).
Under 2.3 and 2.9 and under the conditions formulated in Theorem 2.6, fory every $\bar{c}\in\mathbb{R}_{>0}$, there exists $c_{7}=c_{7}(C_{\operatorname{M}},\bar{c},\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c_{7}\cdot\mathfrak{r}_{N}^{\tau},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{7}\right\}\right)\geq 1-c_{8}\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
for some $c_{8}=c_{8}(d_{V})\in\mathbb{R}_{>0}$.
3   3Application: PDE constrained regression models
We apply the results from Section 2 on the inverse problems described by the PDE Eq. 1.2 and derive corresponding convergence rates for the parameter $f$. We further discuss the applicability of the setup for other PDE models. Throughout this section, we are interested in estimating the unknown scalar function $f:\mathcal{O}\to\mathbb{R}$ and corresponding solutions $u=u_{f}:\mathcal{O}\to\mathbb{R}$ of the Darcy problem. For any $d\in\mathbb{N}$, integer $\alpha>\frac{d}{2}$, a fixed constant $f_{\min}\in\mathbb{R}_{>0}$, and denoting by $\partial_{n}$ the outward pointing normal derivative, define the parameter space
$\mathcal{F}:=\mathcal{F}_{\alpha,f_{\min}}:=\left\{f\in H^{\alpha}(\mathcal{O}):\,f>f_{\min}\text{ on $\mathcal{O}$},f|_{\partial\mathcal{O}}\equiv 1,\partial^{j}_{n}f\equiv 0,j\in\left\llbracket\alpha-1\right\rrbracket\right\},$ (3.1)
and its subclass $\mathcal{F}(\operatorname{L}):=\mathcal{F}\cap H^{\alpha}(\mathcal{O},\operatorname{L})$, $\operatorname{L}\in\mathbb{R}_{>0}$. Let $g\in C^{\infty}(\mathcal{O})$ be a known heat source function. If $f\in\mathcal{F}_{\alpha,f_{\min}}$ for some $\alpha>\frac{d}{2}+1$ and $f_{\min}\in\mathbb{R}_{>0}$, then well-known PDE theory (see Gilbarg and Trudinger [2001] for instance) provides the existence and uniquess of a classical solution $u\in C^{0}(\overline{\mathcal{O}})\cap C^{2+\nu}(\mathcal{O})$ of the Dirichlet boundary value problem
$\begin{dcases}\operatorname{div}(f(x)\nabla u(x))&=g(x)\text{ on $\mathcal{O}$,}\\ \hfill u(x)&=0\text{ on $\partial\mathcal{O}$,}\end{dcases}$ (3.2)
for some (small) $\nu\in\mathbb{R}_{>0}$. In other words, the solution operator $\mathcal{F}\ni f\mapsto G^{D}(f)=:u_{f}\in\mathbb{L}^{2}(\mathcal{O})$ is always well-defined. The nonlinearity of $G^{D}$ becomes more clear when considering the Feynmann-Kac formula, which reads for each $x\in\mathcal{O}$ as
$u_{f}(x)=-\frac{1}{2}\mathbb{E}^{x}\left[\int_{0}^{\tau_{\mathcal{O}}}g(X_{s}^{f})\mathrm{d}s\right],$
where $(X_{t}^{f})_{t\in\mathbb{R}_{\geq 0}}$ is a Markov process starting at $x$ with associated exit time $\tau_{\mathcal{O}}$. We refer to Bass [1998] for an extensive overview and more details. Following Eq. 2.1, we then observe
$Y_{i}=G^{D}(f_{\mathfrak{o}})(X_{i})+\sigma\varepsilon_{i},i\in\left\llbracket N\right\rrbracket,$ (3.3)
where $X_{1},\dots,X_{N}$ are drawn from the uniform law on the domain $\mathcal{O}$, where we assume for simplicity that $\operatorname{vol}(\mathcal{O})=1$. The inverse problem consists of recovering the ground truth $f_{\mathfrak{o}}\in\mathcal{F}$ by using a data set $D_{N}={(Y_{i},X_{i})}_{i\in\left\llbracket N\right\rrbracket}\in(\mathbb{R}\times\mathbb{R})^{N}$, whose joint product law is denoted by $\mathbb{P}_{f_{\mathfrak{o}}}^{N}$. We apply the results of Section 2 to a suitable re-parametrization $\Theta$ of $\mathcal{F}$, which then defines a linear subspace of $H^{\alpha}(\mathcal{O})$. Therefore, we introduce link functions $\Psi$.
Definition § 3.1 ((Regular) Link Function).
Let $f_{\min}\in\mathbb{R}_{>0}$.
  1. i)
    A function $\Psi:\mathbb{R}\to\mathbb{R}_{>f_{\min}}$ is called a link function, if $\Psi$ is a smooth, strictly increasing bijective map with $\Psi(0)=1$ and $\Psi^{\prime}(x)>0$ on $\mathbb{R}$.
  2. ii)
    A link function $\Psi$ is called regular, if and only if all their derivatives are bounded, i.e.
    $\sup_{x\in\mathbb{R}}\left\{\left|\Psi^{(k)}(x)\right|\right\}<\infty,$
    for all $k\in\mathbb{N}$.
For examples and typical choices of (regular) link functions in practise, we refer to Nickl et al. [2020] and Nickl and Wang [2024]. In the notation of Section 2, throughout this section we set $\mathcal{Z}=\mathcal{O}$, $\mathscr{Z}:=\mathscr{B}_{\mathcal{O}}$ and $\zeta=\mathcal{L}_{\mathcal{O}}$ (recalling that $\operatorname{vol}(\mathcal{O})=1$!). We set $\tilde{\Theta}:=\Theta:=\left\{\Psi^{-1}\circ f:\,f\in\mathcal{F}\right\}=H^{\alpha}_{c}(\mathcal{O})$ (the pulled-back) parameter space, and
$\mathscr{G}^{D}:\Theta\to\mathbb{L}^{2}(\mathcal{O}),\,\mathscr{G}(\theta):=G^{D}(\Psi\circ\theta).$
In other words, for each $\theta\in\Theta$, we identify some $f\in\mathcal{F}$ by $f=\Psi\circ\theta$. To emphasize the identification, we sometimes also write $f=f_{\theta}$. Given the data set $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$, we define a Tikhonov-regularized functional
$J_{\mathfrak{r},N}:\mathcal{F}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\,f\mapsto-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-G^{D}(f)(X_{i})|^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\Psi^{-1}\circ f\right\|_{H^{\alpha}(\mathcal{O})}^{2}$ (3.4)
for any $\mathfrak{r}\in\mathbb{R}_{>0}$. In the notation of Eq. 2.2, we have $J_{\mathfrak{r},N}[f]=\mathscr{J}_{\mathfrak{r},N}[\theta]$ for any $f=f_{\theta}\in\mathcal{F}$, such that maximizing $J_{\mathfrak{r},N}$ over $\mathcal{F}$ is equivalent to maximizing $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta=H^{\alpha}_{c}(\mathcal{O})$. Any pairs of maximizer will be denoted by
$\hat{f}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{f\in\mathcal{F}}\left\{J_{\mathfrak{r},N}[f]\right\}\text{ and }\hat{\theta}_{\textsc{MAP}}=\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\in\operatorname*{arg\,max}_{\theta\in H^{\alpha}_{c}(\mathcal{O})}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}.$
For a link function $\Psi$, define for $f_{1},f_{2}\in\mathcal{F}$
$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{1},f_{2}):=\left\|G(f_{1})-G(f_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}^{2}+\mathfrak{r}^{2}\left\|\Psi^{-1}\circ f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2}.$
Identifying $f_{\theta_{1}}$ and $f_{\theta_{2}}$ with $\theta_{1},\theta_{2}\in H^{\alpha}_{c}(\mathcal{O})$, respectively, we see
$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{\theta_{1}},f_{\theta_{2}})=\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}).$
Theorem § 3.2 (Consistency and stability - Darcy problem).
Let $\mathcal{F}$ be as in Eq. 3.1 for some $\alpha>1+\frac{5}{2}d$ and $f_{\min}\in\mathbb{R}_{>0}$. Let $G^{D}(f)=u_{f}$ denote the unique solution of Eq. 3.2 and let $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$ from Eq. 3.3 for some $f_{\mathfrak{o}}\in\mathcal{F}$. Let $\Psi$ be a regular link function. Let $J_{\mathfrak{r}_{N},N}$ be as in Eq. 3.4 with
$\mathfrak{r}_{N}:=N^{-\frac{\alpha+1}{2(\alpha+1)+d}}.$
Then the following hold.
  1. i)
    For each $f_{\mathfrak{o}}\in\mathcal{F}$ and, almost surely under $\mathbb{P}_{f_{\mathfrak{o}}}^{N}$, there exists a maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$.
  2. ii)
    Let $\operatorname{L}\in\mathbb{R}_{>0}$. Then there exist constants $c_{1}=c_{1}(d,\sigma,\mathcal{O},f_{\min},g,\Psi,\alpha,\operatorname{L}),c_{2},c_{3}\in\mathbb{R}_{>0}$, such that for any maximizer $\hat{f}_{\textsc{MAP}}$, we have for all $\operatorname{M}\geq c_{1}$
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{P}_{f}^{N}\left(\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f)\geq\operatorname{M}\mathfrak{r}_{N}^{2}\right\}\right)\right\}\leq c_{2}\exp\left(-c_{3}\operatorname{M}^{2}N\mathfrak{r}_{N}^{2}\right).$ (3.5)
  3. iii)
    Assume additionally that $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\bar{c}\in\mathbb{R}_{>0}$ exists $c_{2}=c_{2}(\alpha,d,\mathcal{O},f_{\min},g_{\min},f_{\mathfrak{o}},\sigma)$ large enough and $c_{3}\in\mathbb{R}_{>0}$, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\left\{\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{2}\mathfrak{r}_{N}^{\frac{\alpha-1}{\alpha+1}},\,\left\|\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{2}\right\}\right)\geq 1-c_{3}\exp(-c_{1}N\mathfrak{r}_{N}^{2}).$
Theorem § 3.3 (Prediction and estimation risk).
Grant the conditions formulated in Theorem 3.2.
  1. i)
    (Prediction Error) For every $\operatorname{L}\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$, there exists a finite constant $c_{4}=c_{4}(\alpha,\beta,d,\mathcal{O},f_{\min},L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f)\right\|_{H^{\beta}(\mathcal{O})}\right]\right\}\leq c_{4}\cdot N^{-\frac{\alpha+1-\beta}{2(\alpha+1)+d}}.$
  2. ii)
    (Estimation Error) Let $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\operatorname{L}\in\mathbb{R}_{>0}$, there exists $c_{5}=c_{5}(g_{\min},f_{\min},\mathcal{O},d,\alpha,L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\right\}\leq c_{5}\cdot N^{-\frac{\alpha-1}{2(\alpha+1)+d}}.$
Remark § 3.4 (Rate-optimality).
The convergence rates for the prediction risk presented in Theorem 3.3 correspond to the usual interpretation of statistical inverse problems: the smoothness of the parameter of interest $f$ is represented by the parameter $\alpha$. The ill-posedness of the underlying problem by $\kappa=1$, which corresponds to the smoothing property of $G^{D}$ (Indeed, Lemma 23 in Nickl et al. [2020] shows that if $f$ is $\alpha$-smooth, then $G^{D}(f)$ is $\alpha+1$-smooth). The prediction risk is considered with respect to the $H^{\beta}$-norm (which can be interpreted as differentiating $\beta$ times), which yields the characteristic numerator $\alpha+1-\beta$ of the rate. The rates of convergence here coincide with the rates obtained in Nickl et al. [2020] for the Darcy problem. There it is also shown that the rate of convergence of the prediction risk is minimax-optimal for the choice $\beta=2$, which also holds true in the random design model, as the proof can be easily adopted.
The proofs of Theorem 3.2 and Theorem 3.3 are given in detail in Appendix A. They are based on verifying that the forward map $\mathscr{G}^{D}$ satisfies C1)-C4) formulated in 2.3 and 2.9 for suitable parameters $\alpha,\kappa,\gamma_{1},\gamma_{2},\eta_{1}$ and $\eta_{2}$, with $V=W=\mathbb{R}$, and an application of the results in Section 2.
Comments on further models
The general framework presented in Section 2 does not only apply to the Darcy problem, it promises to consider various PDE-related problems. As already brielfly discussed in Remark 2.4, the results in Bohr and Nickl [2023] concerning the MAP estimate are special cases by setting $\eta_{1}=\eta_{2}=\gamma_{2}=0$. Hence, the solution map of the time-independent Schrödinger equation satisfies also meets the requirements of 2.3 and 2.9 taking the results of Nickl et al. [2020] also into account. Moreover, very analogous and elementary computations show that the solution map of the heat equation studied in Kekkonen [2022] also satisfies the conditions in 2.3 and 2.9 with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
In very recent work, Nickl and Titi [2024] are considering a non-linear Bayesian data assimilation model for the periodic $d=2$-dimensional Navier Stokes equation with unknown initial condition. In this setting $u=u_{\theta}:\mathcal{O}\times(0,T)\to\mathbb{R}^{2}$ solves the non-linear evolution equation
$\begin{dcases}\frac{\mathrm{d}u}{\mathrm{d}t}+\nu Au+B(u,u)&=f,\\ \hfill u(0)&=\theta,\end{dcases}$
where $\theta$ is the initial condition of the system, $f$ a forcing term and $\nu\in\mathbb{R}_{>0}$ a viscosity parameter. For this type of equation in two dimensions a complete solution theory already exists. In particular, if the initial condition $\theta$ is contained in some specific function class $\tilde{\mathcal{F}}$, the solution operator ~$\tilde{\mathcal{F}}\ni\theta\mapsto G^{NS}(\theta)=u_{f}\in\mathbb{L}^{2}(\mathcal{O}\times(0,T))$ is always well-defined. Based on a data set $D_{N}\in(\mathbb{R}^{2}\times\mathcal{O}\times(0,T))^{N}$ arising from Eq. 2.1, it seems feasible to apply the theory from Section 2. Indeed, using the theory presented in Nickl and Titi [2024], elementary computations show that the solution map is $(\alpha,0,0)$-regular in $\mathbb{L}^{2}$-norm and uniformly bounded, thus satisfying C1) and C2) from 2.3. C7) from 2.9 can also be shown with $\tau=1$. If C3) of 2.3 is additionally satisfied, one would obtain rates of the form $N^{-\frac{\alpha}{\alpha+2}}$ for estimating $\theta$ both in prediction and estimation error.
Appendix A Appendix
   Proofs of Section 2
In this section, we provide detailed proofs of the main results presented in Section 2. We start with the proof of Proposition 2.5 - the existence result for the MAP-estimator. To do so, fix $N\in\mathbb{N}$ as well as $\mathfrak{r}\in\mathbb{R}_{>0}$ throughout. The proof follows the approach of Nickl et al. [2020], where the existence of $\hat{\theta}_{\textsc{MAP}}$ is shown for the white noise formulation of the model. The authors make use of the direct method of the calculus of variations (see Dacorogna [2008] for instance) and split the proof into two parts: in one part it is shown that $\hat{\theta}_{\textsc{MAP}}$ exists as a maximizer of the functional $\mathscr{J}_{\mathfrak{r},N}$ over bounded subsets of $\Theta$ (local existence). In the other part it is shown that this already suffices to conclude the existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer over the whole parameter space $\Theta$ (global existence). In the present work, we adapt the proof idea to the random design regression setting. We need the functional $\mathscr{J}_{\mathfrak{r},N}$ to be lower semicontinuous with respect to the weak topology of Sobolev spaces. Let $\tilde{\mathscr{H}}$ be either $H^{\alpha}(\mathcal{O})$ or $H^{\alpha}_{c}(\mathcal{O})$. Let $\mathcal{T}_{w}$ denote the weak topology of $\tilde{\mathscr{H}}$, i.e., the coarsest topology with respect to which all bounded linear functionals $L:\tilde{\mathscr{H}}\to\mathbb{R}$ are continuous. With a slight abuse of notation, the subspace topology on subsets of $\tilde{\mathscr{H}}$ is also denoted by $\mathcal{T}_{w}$. By standard theory of functional analysis (see, for instance, Megginson [1998], Theorem 2.6.23), on any closed ball ~$\tilde{\mathscr{H}}(r):=\{\theta\in\tilde{\mathscr{H}}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}\leq r\}$ of radius $r\in\mathbb{R}_{>0}$, the weak topology $\mathcal{T}_{w}$ is metrizable by some metric $\tilde{d}$. Defining analogously :=×$\mathscr{H}:=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\tilde{\mathscr{H}}$, the weak topology $\mathcal{T}_{w}^{W}$ on $\mathscr{H}$ equals the product ×$\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\mathcal{T}_{w}$ (see proof of Theorem 3.10 in Brezis [2011]). By a standard fact of functional analysis, on each closed ball $\mathscr{H}(r):=\{\theta\in\mathscr{H}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq r\}$, the weak topology $\mathcal{T}_{w}^{W}$ is then also metrizable by some metric $d$. In the following we claim that $-\mathscr{J}_{\mathfrak{r},N}:\Theta\cap\mathscr{H}(2^{j})\to\mathbb{R}\cup\{\infty\}$ is lower semicontinous with respect to $\mathcal{T}_{w}^{W}$ for any $j\in\mathbb{N}$. To prove this, we need the following general result from functional analysis.
Lemma § A.1 (Compact operators and weak convergence).
Let $(X,\left\|\cdot\right\|_{X})$ and $(Y,\left\|\cdot\right\|_{Y})$ be two Banach spaces. Let $T:X\to Y$ be a linear compact operator. Assume that $(x_{n})_{n\in\mathbb{N}}\subseteq X$ is a weakly convergent sequence with limit $x\in X$. Then, $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, i.e. $\lim_{n\to\infty}\left\|T[x_{n}]-T[x]\right\|_{Y}=0$.
Proof of 1 (Lemma A.1).
Let $(x_{n})_{n\in\mathbb{N}}\subseteq X$ be a weakly convergent sequence with limit $x\in X$, i.e., for all $L\in X^{*}$, we have $\lim_{n\to\infty}L[x_{n}]=L[x]$. By an application of the uniform boundness principle, there exists $B\in\mathbb{R}_{>0}$, such that $\left\|x_{n}\right\|_{X}\leq B$ for all $n\in\mathbb{N}$. By hypothesis on $T$ being compact, the set $T((x_{n})_{n\in\mathbb{N}})$ is relative compact in $Y$, i.e. $\overline{T((x_{n})_{n\in\mathbb{N}})}$ is compact in $Y$. Hence, there exists a subsequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ with $\lim_{m\to\infty}\left\|T[x_{n_{m}}]-v\right\|_{Y}=0$ for some $v\in Y$. Further, for any $A\in Y^{*}$, we have that $A\circ T\in X^{*}$. Thus, $\lim_{n\to\infty}A[T[x_{n}]]=A[T[x]]$, such that the sequence $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges weakly in $Y$ with limit $T[x]\in Y$. By uniquess of the weak limit, we deduce $v=T[x]$. Overall, we have that the sequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$. Now, we see that each subsequence $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, which already proves the claim.
Remark § A.2 (Embeddings and weak convergence).
For $d\in\mathbb{N}$ let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be a bounded domain with smooth boundary. Denoting by $B_{pq}^{s}(\mathcal{O})$ the ususal Besov space with parameters $p,q\in\mathbb{R}_{>0}\cup\{\infty\}$ and $s\in\mathbb{R}$, we have that the embedding
$B_{p_{1}q_{1}}^{s_{1}}(\mathcal{O})\hookrightarrow B_{p_{2}q_{2}}^{s_{2}}(\mathcal{O})$
is compact if $p_{1},p_{2},q_{1},q_{2}\in\mathbb{R}_{>0}\cup\{\infty\}$, $-\infty<s_{2}<s_{1}<\infty$ and $v:=s_{1}-s_{2}-d\max\{\frac{1}{p_{1}}-\frac{1}{p_{2}},0\}\in\mathbb{R}_{>0}$ (see Edmunds and Triebel [1996] for definitions and properties). Let $\alpha\in\mathbb{R}_{>0}$ and $\eta\in\mathbb{R}_{\geq 0}\setminus\mathbb{N}$. Identifying $H^{\alpha}(\mathcal{O})=B_{22}^{\alpha}(\mathcal{O})$, $\mathbb{L}^{2}(\mathcal{O})=B_{22}^{0}(\mathcal{O})$ and $C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$, we have in particular, that the embeddings $H^{\alpha}(\mathcal{O})\hookrightarrow\mathbb{L}^{2}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})$ for $\alpha>\eta+\frac{d}{2}$ are compact. Further, following Proposition 4.5 in Triebel [2006], we have for any $k\in\mathbb{N}$ the continous embedding $B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\overline{\mathcal{O}})$. Making use of above results, we thus have that the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is also compact, provided that $\alpha>k+\frac{d}{2}$. If $(W,\left|\cdot\right|_{W})$ denotes a $d_{W}$-dimensional $\mathbb{R}$-vector space, the same results apply, if we replace the function spaces $F(\mathcal{O})$ by their tensor products $F(\mathcal{O},W)$. Now let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\mathscr{H}$ be a sequence, such that $\theta_{n}\to\theta\in\mathscr{H}$ with respect to $\mathcal{T}_{w}^{W}$. Then, an application of Lemma A.1 yields that $(\theta_{n})_{n\in\mathbb{N}}$ converges strongly in $\mathbb{L}^{2}(\mathcal{O},W)$ as well as in $C^{\eta}(\mathcal{O},W)$, if $\alpha>\eta+\frac{d}{2}$ for $\eta\in\mathbb{R}_{\geq 0}$.
We derive that the forward map $\mathscr{G}$ is continuous, which is necessary to prove the lower semicontinuity of $-\mathscr{J}_{\mathfrak{r},N}$.
Lemma § A.3 (Continuity of $\mathscr{G}$).
For $\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$ let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C3) of 2.3. Let $\Theta_{o}\subseteq\Theta$ be a bounded subset of $\mathscr{H}$ with $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$. Let $(\Omega,\mathcal{A},\mathbb{P})$ be the underlying probability space of $Z:\Omega\to\mathcal{Z}$. Then, the map
$(\Theta_{o},d)\ni\theta\mapsto\mathscr{G}(\theta)(Z)\in(V,\left|\cdot\right|_{V})$
is continuous $\mathbb{P}$-almost surely.
Proof of 2 (Lemma A.3).
Fix some $\omega\in\Omega$ (outside of a $\mathbb{P}$-null-set). Let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta_{o}$ be a convergent sequence with weak limit $\theta\in\Theta_{o}$, i.e., $\lim_{n\to\infty}d(\theta_{n},\theta)=0$. Note that $\left\|\theta_{n}\right\|_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq B$ for some $B\in\mathbb{R}_{>0}$ uniformly over all $n\in\mathbb{N}$. The claim is proven, if $(\varphi_{n}(Z(\omega)))_{n\in\mathbb{N}}\subseteq\mathbb{R}$, defined by
$\varphi_{n}(Z(\omega)):=\left|\mathscr{G}(\theta_{n})(Z(\omega))-\mathscr{G}(\theta)(Z(\omega))\right|_{V}$
is a null-sequence. By hypothesis C3) and the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$, there exists a universal constant $c\in\mathbb{R}_{>0}$, such that
$\varphi_{n}(Z(\omega))\leq\left\|\mathscr{G}(\theta_{n})-\mathscr{G}(\theta)\right\|_{\infty}$ $\leq c(1+B^{\gamma_{2}})\left\|\theta_{n}-\theta\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Exploiting Remark A.2, the right hand side converges to zero, which proves the claim.
Lemma § A.4 (Uniform continuity of $\mathscr{G}$).
Let $\mathscr{G}$ satisfy C2) and C3) of 2.3. Let $D_{N}=((Y_{i},Z_{i}))_{i\in\left\llbracket N\right\rrbracket}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for $N\in\mathbb{N}$. Let $r\in\mathbb{R}_{>0}$. Then, the map
$E:(\Theta\cap\mathscr{H}(r),d)\to(\mathbb{R},|\cdot|),\,\theta\mapsto E[\theta]:=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
is almost surely uniformly continuous.
Proof of 3 (Lemma A.4).
For any $\delta\in\mathbb{R}_{>0}$, let $M_{\delta}$ be the modulus of continuity of $E$, i.e.,
$M_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|E[\theta]-E[\tilde{\theta}]|\right\},$
which defines a random variable. If $(\Omega,\mathcal{A},\mathbb{P})$ denotes the underlying probability space supporting the law of $Z$ and $\varepsilon$, respectivly, we define the measurable event
$A:=\left\{\omega\in\Omega:\,\lim_{\delta\to 0}M_{\delta}=0\right\}.$
We aim to show $\mathbb{P}(A)=1$. As $M_{\delta}$ is decreasing in $\delta$, it suffices to verify that $\lim_{\delta\to 0}\mathbb{E}[M_{\delta}]=0$. Indeed, elementary computations show that
$|E[\theta]-E[\tilde{\theta}]|$ $\leq\frac{2C_{\mathscr{G},\operatorname{B}}}{N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}+\frac{1}{N\sigma}\sum_{i\in\left\llbracket N\right\rrbracket}|\langle\varepsilon_{i},\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\rangle_{V}|.$
Analogously, define for each $\delta\in\mathbb{R}_{>0}$ and $Z\sim\zeta$ the modulus of continuity of $\mathscr{G}$ by
$R_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{\left|\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}\right\},$
which in turns defines a random variable. In Lemma A.3, we have shown that $\mathscr{G}$ is almost sure continuous as a mapping from $\Theta\cap\mathscr{H}(r)$ to $V$. As $\Theta\cap\mathscr{H}(r)$ is a compact metric space, $\mathscr{G}$ is in fact almost surely uniformly continuous, and hence $\lim_{\delta\to 0}R_{\delta}=0$ almost surely. Moreover, since $Z$ and $\varepsilon\sim N(0,\operatorname{Id}_{V})$ are independent by hypothesis, we have
$\mathbb{E}\left[\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\langle\varepsilon,\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\rangle_{V}|\right\}\right]$ $\leq\mathbb{E}[\left|\varepsilon\right|_{V}R_{\delta}]=\sqrt{d_{V}}\cdot\mathbb{E}[R_{\delta}].$
Now by exploiting C2), $(R_{\delta})_{\delta\in\mathbb{R}_{>0}}$ is uniformly integrable and hence, the last expectation is converging to zero for $\delta\to 0$. Overall, we obtain
$\mathbb{E}[M_{\delta}]\leq\left(\frac{2C_{\mathscr{G},\operatorname{B}}}{\sigma^{2}}+\frac{\sqrt{d_{V}}}{\sigma}\right)\mathbb{E}[R_{\delta}]\overset{\delta\to 0}{\longrightarrow}0,$
which proves the claim.
We are now providing the proof of Proposition 2.5.
Proof of 4 (Proposition 2.5).
Grant the assumptions of Proposition 2.5. We follow along the lines of Nickl et al. [2020] and split the proof of existence in two parts.
  1. i)
    Local Existence: Let $j\in\mathbb{N}$ be arbitrarily. We show now that $\mathscr{J}_{\mathfrak{r},N}$ has a maximizer over $\Theta\cap\mathscr{H}(2^{j})$. As $\Theta$ is assumed to be weakly closed and $\mathscr{H}(2^{j})$ is weakly sequentially compact, due to the separability and reflexivity of Sobolev spaces and by an applycation of the Banach-Alaoglu theorem, any sequence $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta\cap\mathscr{H}(2^{j})$ has a weakly convergent sub-sequence $(\theta_{n_{k}})_{k\in\mathbb{N}}$ with weak limit $\theta\in\Theta\cap\mathscr{H}(2^{j})$. Thus, following the Direct Method of the Calculus of Variations, it suffices to show that $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ is almost surely weakly lower semicontinuous. Rewrite $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ as
    $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}(\theta)=E_{1}(\theta)+E_{2}(\theta)$ (A.1)
    with
    $E_{1}(\theta):=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\text{ and }E_{2}(\theta):=\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
    $E_{2}$ is weakly lower semicontinuous, which is a standard result in functional analysis concerning norms on Banach spaces. In Lemma A.4, we have derived that $E_{1}$ is almost surely uniformly continuous as a map between the metric spaces $(\Theta\cap\mathscr{H}(2^{j}),d)$ and $(\mathbb{R},|\cdot|)$. In particular, $E_{1}$ is weakly lower semicontinuous. All in all, local existence of $\hat{\theta}_{\textsc{MAP}}$ follows immediately by an application of the Direct Methods of Calculus of Variations (see Dacorogna [2008]).
  2. ii)
    Global Existence: For each $j\in\mathbb{N}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$, we define the event
    $A_{j}:=\left\{\omega\in\Omega:\,\hat{\theta}_{\textsc{MAP}}(\omega)\notin\Theta\cap\mathscr{H}(2^{j})\right\}.$
    Then, $(A_{j})_{j\in\mathbb{N}}$ satisfies $\lim_{j\to\infty}\mathbb{P}[A_{j}]=0$ according to Theorem 2.6. Thus, choosing $j\in\mathbb{N}$ large enough, we have
    $\sup_{\theta\in\Theta\cap\mathscr{H}(2^{j})}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}=\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}$
    with probability as close to one as desired.
We now approaching the proof of Theorem 2.6. Therefore, we need some notation and further helping lemma.
Notation § A.5 (Empirical Measures).
For $N\in\mathbb{N}$ let $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ be a data set. For any $(y,z)\in V\times\mathcal{Z}$ denote by $\delta_{(y,z)}(\cdot)$ the Dirac measure on $V\times\mathcal{Z}$. The empirical measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ based on $D_{N}$ is then defined by
^$\widehat{\mathbb{P}}_{N}(A):=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}\delta_{(Y_{i},Z_{i})}(A)$
for any $A\in\mathscr{B}_{V\times\mathcal{Z}}$. Evidently, we have for any measurable function $h:V\times\mathcal{Z}\to\mathbb{R}$
$\int_{V\times\mathcal{Z}}h(y,z)\mathrm{d}\widehat{\mathbb{P}}_{N}(y,z)=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}h(Y_{i},Z_{i}).$
The following helping lemma provides M-estimation techniques, which are used to prove Theorem 2.6. The proof follows along the lines of van de Geer [2001], Nickl et al. [2020] and Bohr and Nickl [2023].
Lemma § A.6 (M-estimation techniques).
With the notation of Theorem 2.6 and the lines above it, we have for any $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ and $\tilde{\theta}\in\Theta$, $\mathfrak{r}\in\mathbb{R}_{>0}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$ (A.2)
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-almost surely. Moreover, for any $R,\mathfrak{r}\in\mathbb{R}_{>0}$ and some fixed constant $c_{0}\in\mathbb{R}_{\geq 2}$ define the measurable event
$\Xi_{\mathfrak{r},R}:=\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{0}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}.$
On $\Xi_{\mathfrak{r},R}$, we have further
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}\cdot R^{2}$ (A.3)
for some $c_{1}\in\mathbb{R}_{\geq 1}$, as well as
$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})-\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})\geq\frac{1}{6}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}).$ (A.4)
Proof of 5 (Lemma A.6).
The proof of Eq. A.2 is elementary and follows along the lines of Lemma 4.10 in Nickl and Wang [2024] (replacing the finite-dimensional approximation of $\theta_{\mathfrak{o}}$ by some arbitrary $\tilde{\theta}\in\Theta$). We continue with the proof of Eq. A.3. On the event $\Xi_{\mathfrak{r},R}$, we have
$c_{0}\cdot R^{2}$ $\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})-c_{0}\cdot\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$
$\leq 2\left(\frac{1}{2}\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\ -\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right)+\mathfrak{r}^{2}\left(\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right)$
$\leq 2\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}-\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right)^{2}+2\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\mathfrak{r}^{2}\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$
$\leq 2\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}),$
where we have used $\frac{1}{2}a^{2}-b^{2}\leq(a-b)^{2}$ for two real numbers $a,b\in\mathbb{R}$ as well as the reversed triangle inequality. Rearranging, the last inequality shows Eq. A.3. Eq. A.4 follows by the same elementary computations as in van de Geer [2001] (see pp. 3-4).
Proof of 6 (Theorem 2.6).
The proof follows the key ideas of Bohr and Nickl [2023]. Since we are allowing a more general $\tilde{\theta}\in\Theta$, we also need to adapt the ideas of Nickl et al. [2020]. We start with the exponential inequality Eq. 2.4. Fix $\tilde{\theta}\in\Theta$. According to Lemma A.6, it sufficies to controll the $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability of the event
$\tilde{\Xi}_{\mathfrak{r},R}:=\Xi_{\mathfrak{r},R}\cap\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})\right\}.$
For any $\theta\in\Theta$, we define the empirical process
$\mathcal{W}_{N}(\theta):=\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\theta}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)$
and obtain with Eq. A.3 and Eq. A.4 from Lemma A.6
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\tilde{\Xi}_{\mathfrak{r},R})\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}^{2}R^{2},\,\frac{1}{12}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\leq|\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})|\right\}\right).$
for some $c_{1}\in\mathbb{R}_{\geq 1}$. In the following step, we apply a so-called peeling device, where we refer to van de Geer [2000] for a generic explanatation. This method was already used in a similar way in Nickl et al. [2020] and Bohr and Nickl [2023] to controll the process $\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})$ over disjoint slices of a subset of $\Theta$. In detail, for all $l\in\mathbb{N}$ and $c_{1}$ sufficiently large, define the family of slices
[…]
where the inclusion follows from the definition of $\mathfrak{d}_{\mathfrak{r}}^{2}$. Note that $\{\theta\in\Theta:\,\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\geq R^{2}c_{1}^{2}\}=\bigcup_{l\in\mathbb{N}}\Theta_{l}$. We obtain
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right).$
We follow the approach of Bohr and Nickl [2023] and decompose $\mathcal{W}_{N}$ into two empirical processes. Indeed, for each $\theta\in\Theta$, simple computations show
$N\mathcal{W}_{N}(\theta)$ $=\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})],$
which evidently defines a centered process. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, we further decompose the log-likelihood difference of the last display by
$\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})$ $=\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|Y_{i}-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}$
[…]
Defining for all $\theta\in\Theta$
$T_{1}(\theta):=\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V},$ (A.5)
and
$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\},$ (A.6)
respectively, we have that $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{1}(\theta)]=0$, as $Z$ and $\varepsilon$ are independent and $\varepsilon$ is centered. In other words, we have derived so far
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)$ $\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right)$
$\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)+\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right).$
We will discuss the two different probabilties separately and starting with the the multiplier-type process defined in Eq. A.5. To apply Lemma A.7, we denote for each $j\in\left\llbracket d_{V}\right\rrbracket$ by $\mathscr{G}(\cdot)_{j}:\Theta\to\mathbb{L}^{2}(\mathcal{O})$ the $j$-th coordinate of the forward map $\mathscr{G}:\Theta\to\mathbb{L}^{2}(\mathcal{O},V)$, where we again identify $V$ with $\mathbb{R}^{d_{V}}$. We define a class of measurable functions by
$\mathcal{H}_{l,j}:=\left\{\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}:\mathcal{Z}\to\mathbb{R}:\,\theta\in\Theta_{l}\right\}.$ (A.7)
To shorten notation in the subsequent, we write $h_{\theta,j}:=\mathscr{G}(\theta_{\mathfrak{o}})_{j}-\mathscr{G}(\theta)_{j}$. Thus, for any $\theta\in\Theta$, we have
[…]
where $\varepsilon_{ij}\overset{i.i.d.}{\sim}\operatorname{N}(0,1)$ for $(i,j)\in\left\llbracket N\right\rrbracket\times\left\llbracket d_{V}\right\rrbracket$. Conclusevly, we see
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|T_{1}(\theta)|\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)\leq d_{V}\max_{j\in\left\llbracket d_{V}\right\rrbracket}\left\{\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{2^{2l}}{96d_{V}}R^{2}c_{1}^{2}\sigma\right)\right\}.$
We now verify that the function class $\mathcal{H}_{l,j}$ defined in Eq. A.7 satisfies the assumptions formulated in Lemma A.7. Indeed, by C2) in 2.3, we have uniformly for all $\theta\in\Theta_{l}$
$\sup_{z\in\mathcal{Z}}\left\{|h_{\theta,j}(z)|\right\}=\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})_{j}(z)-\mathscr{G}(\theta)_{j}(z)|\right\}\leq\sup_{z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\tilde{\theta})(z)-\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq 2C_{\mathscr{G},\operatorname{B}}.$
Moreover, for $Z\sim\zeta$ and all $\theta\in\Theta_{l}$, $l\in\mathbb{N}$, we have
$\mathbb{E}^{Z}[h_{\theta,j}^{2}(Z)]\leq\sum_{j\in\left\llbracket d_{V}\right\rrbracket}\int_{\mathcal{Z}}\left|\mathscr{G}(\theta)_{j}(Z)-\mathscr{G}(\tilde{\theta})_{j}(Z)\right|_{V}^{2}\mathrm{d}\zeta(z)=\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}.$
Observing that $\theta\in\Theta_{l}$, the last expression is further upper bounded by
$\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\leq 2^{2l}R^{2}c_{1}^{2}=:v_{l}^{2}.$
Additionally, we need to upper bound the $\mathbb{L}^{2}$- and $\mathbb{L}^{\infty}$-metric entropy integrals $\mathcal{J}_{2}$ and $\mathcal{J}_{\infty}$. Thus, we proceed by upper bounding the respective metric entropies
$\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right),$
where $d_{2}$ and $d_{\infty}$ are defined according to Lemma A.7. Starting with the first one, let $\theta_{1},\theta_{2}\in\Theta_{l}$ be arbitrarily chosen. Exploiting C1) in 2.3, we have
$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta_{2}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
By definition of $\Theta_{l}$, $\theta_{1}$, $\theta_{2}$ are contained in a $H^{\alpha}$-ball, such that the polynomial growth in the last display can be upper bounded. It suffices to find a $\rho$-covering of $\Theta_{l}$ with respect to the dual $(H^{\kappa}(\mathcal{O},W))^{*}$-norm. Applying Lemma A.9, there exists a constant $c_{2}=c_{2}(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that for all $\rho\in\mathbb{R}_{>0}$
$N(\Theta_{l},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}\right).$
Consequently, for every $\rho\in\mathbb{R}_{>0}$ there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with $N$ upper bounded by the line above. Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{2}(\theta,\theta^{(i_{\circ})})=\left\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)\cdot\rho=:m_{1}\cdot\rho.$
Thus,
$\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)\leq c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}m_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}=c_{3}\cdot\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}$
for some constant $c_{3}=c_{3}(\alpha,\kappa,d,d_{W})\in\mathbb{R}_{>0}$. An analogous argument yields a similar upper bound for the $d_{\infty}$-metric entropy. Indeed, for any $\theta_{1},\theta_{2}\in\Theta_{l}$, exploiting C3) in 2.3, we have
$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
Using that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$, the polynomial growth can be bounded due to the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$. Thus, it suffices to find an covering for $\Theta_{l}$ with respect to the $C^{\eta_{2}}$-norm. Indeed, by an application of Lemma A.9, we have
$N(\Theta_{l},\left\|\cdot\right\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)\leq\exp\left(c(\alpha,\eta_{2},d)\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\right).$
Now, for $\rho\in\mathbb{R}_{>0}$ fixed, let $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ be a covering of $\Theta_{l}$ w.r.t. $\left\|\cdot\right\|_{C^{\eta_{2}}(\mathcal{O},W)}$ with $N$ bounded by the line above. Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
$d_{\infty}(\theta,\theta^{(i_{\circ})})\leq c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right)\cdot\rho.$
Thus, there exists a constant $c_{4}=c_{4}(\alpha,\eta_{2},d,d_{W})\in\mathbb{R}_{>0}$, such that
$\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\leq c_{4}\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta}}\text{ with}m_{2}:=c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right).$
We are now able to upper bound the corresponding entropy integrals. Indeed, as $2(\alpha+\kappa)>d$, we have
$\mathcal{J}_{2}(\mathcal{H}_{l,j})=\int_{0}^{4v_{l}}\sqrt{\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)}\mathrm{d}\rho\leq\sqrt{c_{3}}\cdot\int_{0}^{4v_{l}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{2(\alpha+\kappa)}}\mathrm{d}\rho=\sqrt{c_{3}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}},$
where the numerical value of constant $c_{3}$ changed without any further dependencies. To shorten notation in the further analysis write $\tau_{1}:=\frac{d}{2(\alpha+\kappa)}\in(0,1)$. By assumption $R\geq\mathfrak{r}$. We have
$\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}}$ $=c_{1}^{\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)^{\tau_{1}}\cdot 2^{l\tau_{1}}\cdot R^{\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}}\cdot v_{l}^{-1-\tau_{1}}$
$\leq c_{1}^{\tau_{1}+\gamma_{1}\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot 2^{\tau_{1}}\cdot 2^{l\tau_{1}+l\tau_{1}\gamma_{1}}\cdot R^{\tau_{1}+\gamma_{1}\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}-\tau_{1}\gamma_{1}}\cdot v_{l}^{-1-\tau_{1}}$
$=c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot 2^{\tau_{1}}\cdot 2^{-l(1-\tau_{1}\gamma_{1})}\cdot R^{-1+\gamma_{1}\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}(+\tau_{1}\gamma_{1})}$
$\leq c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot 2^{\tau_{1}}\cdot\sqrt{N},$
where we have used Eq. C4 and the fact that $\alpha+\kappa>\frac{d}{2}\gamma_{1}$. In particular, the last bound implies the existence of a constant $c_{5}=c_{5}(\alpha,\gamma_{1},\kappa,d,d_{W},C_{\operatorname{Lip},2})\in\mathbb{R}_{>0}$, such that
$\mathcal{J}_{2}(\mathcal{H}_{l,j})\leq c_{5}\cdot c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot v_{l}^{2}\cdot\sqrt{N}.$
Evidently, the upper bound of $\mathcal{J}_{2}(\mathcal{H}_{l,j})$ can be as small as desired by choosing $c_{1}$ accordingly large. Analogously, as $\alpha-\eta>d$, we have
$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})=\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\mathrm{d}\rho\leq c_{4}\cdot\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\mathrm{d}\rho\leq c(\alpha,\eta_{2},d,d_{W},C_{\mathscr{G},\operatorname{B}})\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{\alpha-\eta_{2}}}.$
Using Eq. C5 with similar arguments as above, one can verify analogously verify
$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})\leq c_{6}v_{l}^{2}N$
with some $c_{6}=c_{6}(\alpha,\eta_{1},\eta_{2},\gamma_{2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},c_{1})\in\mathbb{R}_{>0}$, which can be also as small as desired by choosing $c_{1}$ sufficiently large. Now in the setting of Lemma A.7, for arbitrary $\bar{c}\in\mathbb{R}_{>0}$, such that $x:=\bar{c}N2^{2l}R^{2}\in\mathbb{R}_{\geq 1}$, we have for some constant $L\in\mathbb{R}_{>0}$
$L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\leq L\left(c_{5}\cdot v_{l}^{2}\sqrt{N}+v_{l}\sqrt{\bar{c}N2^{2l}R^{2}}+N^{-\frac{1}{2}}(c_{6}\cdot 2^{2l}R^{2}N+8C_{\mathscr{G},\operatorname{B}}\bar{c}N2^{2l}R^{2})\right)$
$\leq\frac{\sqrt{N}v_{l}^{2}}{96\cdot d_{V}}\left(c_{5}\cdot c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot d_{V}\cdot 96\cdot L+L\cdot 96\cdot d_{V}\frac{\sqrt{\bar{c}}}{\sqrt{c_{1}}}+c_{6}\cdot 96\cdot d_{V}\cdot L+\frac{8C_{\operatorname{Lip},2}\bar{c}}{c_{1}^{2}}\cdot 96\cdot d_{V}\cdot L\right)\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}},$
which is attainable for $c_{1}=c_{1}(\alpha,\eta_{2},\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ large enough. Hence an applycation of Lemma A.7 yields
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{v_{l}^{2}\sigma}{96d_{V}}\right)$ $\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\right)$
$\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
As the last upper bound is independent of the choice of $j\in\left\llbracket d_{V}\right\rrbracket$, we have overall
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{v_{l}^{2}}{96}\right)\leq 2d_{V}\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
It remains to study the process $T_{2}$. By writing
$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}=:\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i}),$
we need to controll
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{\left|\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[h_{\theta}(Z_{i})])\right|\right\}\geq\sqrt{N}\frac{2^{2l}}{48}R^{2}c_{1}^{2}\sigma^{2}\right\}\right).$
Similiar as before, we define for each $l\in\mathbb{N}$ a class of measurable functions
$\mathcal{H}_{l}:=\left\{h_{\theta}:=\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)\right|_{V}^{2}:\mathcal{Z}\to\mathbb{R}\,:\,\theta\in\Theta_{l}\right\},$
and verify the assumptions of Lemma A.7. Indeed, note that for each $\theta\in\Theta_{l}$, we have
$\sup_{z\in\mathcal{Z}}\left\{\left|h_{\theta}(z)\right|\right\}$ $=\sup_{z\in\mathcal{Z}}\left\{\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\tilde{\theta})(z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right|\right\}\leq 8C_{\mathscr{G},\operatorname{B}}$
due to C2) in 2.3. Moreover, for $Z\sim\zeta$ and all $\theta\in\Theta_{l}$, $l\in\mathbb{N}$, we have
$\mathbb{E}^{Z}[h_{\theta}^{2}(Z)]$ $=\mathbb{E}^{Z}\left[\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\tilde{\theta})(Z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}^{2}\right|^{2}\right]\leq 16C_{\mathscr{G},\operatorname{B}}^{2}\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq 16C_{\mathscr{G},\operatorname{B}}^{2}v_{l}^{2},$
where we have applied C2) from 2.3 and used the notation as before. In order to bound the entropy integrals, observe that for any $\theta_{1},\theta_{2}\in\Theta_{l}$
$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)},$
as well as
$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\infty}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\infty}$ $\leq 4C_{\mathscr{G},\operatorname{B}}\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}.$
Evidently, we obtain the same upper bounds for the metric entropies
$\log N\left(\mathcal{H}_{l},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l},d_{\infty},\rho\right),$
as already seen for the process $T_{1}$, by a change of constant. Thus, following the argumentation as before, we can choose $c_{1}:=(\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ sufficiently large to obtain
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
Overall, we have derived the existence of constants $c_{7}=c_{7}(d_{V})\in\mathbb{R}_{>0}$, such that
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{7}\cdot\sum_{l\in\mathbb{N}}\exp(-\bar{c}N2^{2l}R^{2})\leq c_{7}\exp(-\bar{c}NR^{2}),$
if $c_{1}\in\mathbb{R}_{\geq 1}$ in the definition of $\Xi_{\mathfrak{r},R}$ is choosen sufficiently large, observing that the numerical values of the constants may changed again in the last line without any further dependencies. This shows the exponential inequality in i). Now we verify the upper bound of the expectation, i.e. ii). Let $\tilde{\theta}\in\Theta$ be fixed. For some appropriate $\mathfrak{m}\in\mathbb{R}_{>0}$ specified later, we have
$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+\int_{(\mathfrak{m},\infty)}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq t)\mathrm{d}\mathcal{L}(t).$
Let $\mathfrak{r}\in\mathbb{R}_{>0}$ satisfy Eq. C6 and fix some $\bar{c}\in\mathbb{R}_{\geq 1}$. We define the map $T:(\mathfrak{r}^{2},\infty)\to\mathbb{R}_{>0},\,t\mapsto c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)$ and set $\mathfrak{m}:=c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+\mathfrak{r}^{2})$, with $c_{1}$ from the first part of the theorem. Applying the transformation theorem, we obtain
$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot\int_{(\mathfrak{r}^{2},\infty)}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)\right)\mathrm{d}\mathcal{L}(t).$
Now observe that $t\in(\mathfrak{r}^{2},\infty)$ satisfies $\sqrt{t}\geq\mathfrak{r}$ and hence also Eq. C4 and Eq. C5 by an application of Eq. C6. Thus, by an application of the first part of the theorem
$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot c(d_{V})\cdot\int_{(\mathfrak{r}^{2},\infty)}\exp(-\bar{c}Nt)\mathrm{d}\mathcal{L}(t)\leq\mathfrak{m}+\frac{c_{1}\cdot c(d_{V})}{\bar{c}N},$
where we have chosen $c_{1}=c_{1}(\bar{c},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma)$ sufficiently large. This shows the claim.
Proof of 7 (Corollary 2.7).
To prove the claim, we aim to apply Theorem 2.6. Therefore choose $\tilde{\theta}=\theta_{\mathfrak{o}}\in\Theta\cap\mathscr{H}(\operatorname{L})$. Let $\mathfrak{r}=\mathfrak{r}_{N}$ be as in the hypothesis. Elementary computations show that Eq. C4, Eq. C5 and Eq. C6 are verified for the choice $R=\mathfrak{r}_{N}$. Moreover, observe for any $c_{1}\in\mathbb{R}_{>0}$ that
$c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+R^{2})=c_{1}\cdot(\mathfrak{r}^{2}\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}+\mathfrak{r}^{2})\leq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}.$
Thus, for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\bar{c},\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ sufficiently large, that
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}\right)\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq 2(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
This shows the first part of the claim. Moreover, according to the second part of Theorem 2.6, we have
$\mathbb{E}_{\theta_{0}}^{N}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|^{2}_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right]$ $\leq c_{3}\cdot\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{3}\cdot\left((\operatorname{L}+1)\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{4}\cdot\mathfrak{r}_{N}^{2}$
for some constant $c_{4}=c_{4}(\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{>0}$. This proves the claim.
Proof of 8 (Corollary 2.10).
In Corollary 2.7, we have seen that for each $\bar{c}\in\mathbb{R}_{\geq 1}$ we have
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot\mathfrak{r}^{2}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}),$
for a constant $c_{1}=c_{1}(\bar{c},,\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ large enough, which is equivalent with
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}_{N}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|^{2}_{H^{\alpha}(\mathcal{O},W)}\leq c_{1}\mathfrak{r}_{N}^{2}\right)\geq 1-c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
Thus, to prove the claim, it suffices to restrict ourselves on the measurable event
$T_{N}:=\left\{\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\mathfrak{r}_{N}^{2},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}\right\}.$
As $\theta_{\mathfrak{o}}\in H^{\alpha}(\mathcal{O},W)$ is fixed, there exists some constant $\operatorname{L}=\operatorname{L}(\theta_{\mathfrak{o}})\in\mathbb{R}_{>0}$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}$ and thus some $c_{2}=c_{2}(c_{1},\operatorname{L})\in\mathbb{R}_{>0}$, such that on the event $T_{N}$, we have
$\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq c_{2}\cdot\mathfrak{r}_{N}^{2}\text{ as well as }\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{2}.$
By an application of C7) in 2.9, we have on the event $T_{N}$
$\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c(C_{\operatorname{M}},c_{2})\cdot\mathfrak{r}_{N}^{\tau},$
which shows the claim.
   Proofs of Section 3
[…]
   Darcy Problem - Proofs
Let $f\in C^{1}(\bar{\mathcal{O}})$ with $f\geq f_{\min}\in\mathbb{R}_{>0}$. We define the differential operator
$L^{D}_{f}:H_{0}^{2}(\mathcal{O})\to\mathbb{L}^{2}(\mathcal{O})\text{, }u\mapsto L_{f}^{D}[u]:=\operatorname{div}(f\nabla u),$
where $\nabla$ denotes the usual gradient and $\operatorname{div}$ the divergence operator. Refering on Evans [2010], it is well known from elliptic PDE theory, that $L^{D}$ is a linear and bounded operator, which posses a linear inverse, denoted by
$V_{f}^{D}:\mathbb{L}^{2}(\mathcal{O})\to H^{2}_{0}(\mathcal{O})\text{, }h\mapsto V_{f}^{D}[h].$
Thus, for any $h\in\mathbb{L}^{2}(\mathcal{O})$, there exists a unique weak solution $w_{f,h}:=V^{D}_{f}[h]\in H^{2}_{0}(\mathcal{O})$ solving the Dirichlet boundary value problem
$\begin{dcases}L_{f}[w_{f,h}]&=h\text{ on $\mathcal{O}$,}\\ \hfill w_{f,h}&=0\text{ on $\partial\mathcal{O}$.}\end{dcases}$
Utilizing the Feynman-Kac formula, Lemma 20 in Nickl et al. [2020] demonstrates that
$\left\|V^{D}_{f}[h]\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|h\right\|_{\infty},$ (A.8)
for all $h\in C^{\eta}(\mathcal{O})$, $\eta\in\mathbb{R}_{>0}$, if additionally $f\in C^{2}(\mathcal{O})$. Keeping the notation of Section 3 in mind, we now verify that the forward map $\mathscr{G}^{D}$ satisfies the conditions C1) - C3) formulated in 2.3. As already suggested in the previous section, C3) is the cruicial part for the Darcy problem, as it fails to satisfy the assumption with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
  1. C1)
    $\mathbb{L}^{2}$-Lipschitz continuity: Let $\alpha>2+\frac{d}{2}$. The forward map $\mathscr{G}^{D}:H^{\alpha}_{c}(\mathcal{O})\to\mathbb{L}^{2}(\mathcal{O})$ is $(\alpha,4,1)$-regular, that is for all $\theta_{1},\theta_{2}\in H^{\alpha}_{c}(\mathcal{O})$, we have
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq C_{\operatorname{Lip},2}\cdot(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{4})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{1}(\mathcal{O}))^{*}}$
    with some universal constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$. The proof can be found in Nickl et al. [2020].
  2. C1)
    Uniform boundedness: As described in Section 3, we identify with each $\theta\in H^{\alpha}_{c}(\mathcal{O})$ some $f=f_{\theta}\in\mathcal{F}$, such that $\mathscr{G}^{D}(\theta)=G^{D}(f_{\theta})=V_{f_{\theta}}^{D}[g]$. Thus, with Eq. A.8, we obtain
    $\sup_{\theta\in H^{\alpha}_{c}(\mathcal{O}),x\in\mathcal{O}}\left\{|\mathscr{G}^{D}(\theta)(x)|\right\}$ $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|G^{D}(f)(x)|\right\}=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|V_{f}^{D}[g](x)|\right\}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|g\right\|_{\infty}.$
  3. C1)
    $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha>2+\frac{d}{2}$. Let $\theta_{1},\theta_{2}\in H^{\alpha}_{c}(\mathcal{O})$ and identify $f_{\theta_{1}}=\Psi(\theta_{1}),f_{\theta_{2}}=\Psi(\theta_{2})\in\mathcal{F}$. With above notation, we observe
    $L^{D}_{f_{\theta_{1}}}[\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})]=L^{D}_{f_{\theta_{1}}}[u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}]=\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}),$
    where the right-hand side defines a function in $C^{\nu}(\mathcal{O})$ for some small $\nu\in\mathbb{R}_{>0}$. In other words, applying the linear inverse $V_{f_{\theta_{1}}}^{D}$ on both sides of the equation, we obtain
    $u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}=V_{f}^{D}[\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})].$
    Thus, by an applycation of Eq. A.8
    $\left\|u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}.$
    Moreover, by an applycation of the product rule
    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}$ $\leq\left\|\nabla(f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}\right\|_{\infty}+\left\|(f_{\theta_{2}}-f_{\theta_{1}})\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$
    $\leq\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}\cdot\left(\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}+\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\right).$
    We study the two remaining terms in the last display seperately, starting with the second one. We have
    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$ $=\left\|f_{\theta_{2}}^{-1}\cdot(g-\nabla u_{f_{\theta_{2}}}\cdot\nabla f_{\theta_{2}})\right\|_{\infty}\leq\left\|f_{\theta_{2}}^{-1}\right\|_{\infty}\left(\left\|g\right\|_{\infty}+\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}\right).$
[…]
    Thus, elementary computation show that
    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3}).$
    Hence, by an further application of Lemma 22 in Nickl et al. [2020], we have
    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
    Overall, we have
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}^{3}\lor\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
    By an application of Lemma 29 in Nickl et al. [2020] and Lemma A.10, we obtain finally
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})}^{4})\cdot\left\|\theta_{2}-\theta_{1}\right\|_{C^{1}(\mathcal{O})}.$
    Thus, the forward map $\mathscr{G}^{D}$ satisfies C3) in 2.3 with $\gamma_{2}=4$, and $\eta_{1}=\eta_{2}=1$.
After prooving the conditions C1) - C3) for the forward map $\mathscr{G}^{D}$ formulated in 2.3, we are showing an analogous inverse continuity modulus for the solution operator $G^{D}$ of the Darcy Problem as assumed in C7) in 2.9. The proof is based on a interpolation inequality for Sobolev spaces, which reads for all $\alpha\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$
$\left\|u\right\|_{H^{\beta}(\mathcal{O})}\leq c(\alpha,\beta,\mathcal{O})\cdot\left\|u\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|u\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$ (A.9)
  1. C’7)
    Inverse Continuity Modulus: Let $f_{1},f_{2}\in\mathcal{F}$ for some $\alpha>2+\frac{d}{2}$. Then there exists a constant $\operatorname{B}\in\mathbb{R}_{>0}$, such that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq\operatorname{B}$. Lemma 24 in Nickl et al. [2020] provides the existence of a finite constant $c_{1}=c_{1}(g_{\min},f_{\min},\operatorname{B},\mathcal{O},d)\in\mathbb{R}_{>0}$, such that
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{1}\cdot\left\|f_{2}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}.$
    By an application of Eq. A.9, we have further
    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}\leq c_{2}(\alpha,\mathcal{O})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
    Moreover, by an application of Lemma 23 in Nickl et al. [2020], we have
    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\leq c_{3}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})$
    with a finite constant $c_{3}=c_{3}(\alpha,d,\mathcal{O},f_{\min},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. Thus, we have
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{4}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}$
    for a constant $c_{4}=c_{4}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{B},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. Now, let $\operatorname{M}\in\mathbb{R}_{>0}$ be arbitrary and assume that $\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}+\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}\leq\operatorname{M}$, such that $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq\delta$ for $\delta\in\mathbb{R}_{>0}$ small enough. By observing that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq c(\operatorname{M},\alpha,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})$, we obtain
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{5}\cdot\delta^{\frac{\alpha-1}{\alpha+1}}$
    with $c_{5}=c_{5}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{M},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. This shows the claim with $\tau=\frac{\alpha-1}{\alpha+1}\in(0,1)$.
Proof of 9 (Theorem 3.2).
Firstly, note that the Darcy problem is formulated as special case of Section 2 with $V=W=\mathbb{R}$.
  1. i)
    The existence of a maximizer $\hat{f}_{\textsc{MAP}}$ follows directly by the equivalent formulation presented in Section 3 and an application of Theorem 2.6 under considerations that the forward map $\mathscr{G}^{D}$ satisfies all assumption of Theorem 2.6.
  2. i)
    As the forward map $\mathscr{G}^{D}$ satisfies C1) - C3) in 2.3, we apply Theorem 2.6 directly with $\tilde{\theta}=\theta_{\mathfrak{o}}$ by observing that
    $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})=\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}}).$
    Thus, for every $\bar{c}\in\mathbb{R}_{>0}$ there exists a constant $c_{1}=c_{1}(\alpha,f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O},\sigma)\in\mathbb{R}_{>0}$ sufficiently large, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\right)\leq c_{2}\exp\leq\left(-\bar{c}m^{2}N\right)$
    for all $m\geq\mathfrak{r}_{N}$ satisfying Eq. C4 and Eq. C5. We identify $f_{\mathfrak{o}}\in\mathcal{F}$ with some $\theta_{\mathfrak{o}}=\Psi^{-1}\circ f_{\mathfrak{o}}\in H^{\alpha}_{c}(\mathcal{O})$ and conclude $\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{3}(\Psi,\operatorname{L})\in\mathbb{R}_{>0}$ uniformly over all $f_{\mathfrak{o}}\in\mathcal{F}(\operatorname{L})$. Choosing $\operatorname{M}\geq 1\lor c_{3}$ and $m:=\operatorname{M}\mathfrak{r}_{N}\geq\mathfrak{r}_{N}$, such that $c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\leq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}$. Thus,
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}\right)\leq c_{2}\exp\leq\left(-\bar{c}M^{2}N\mathfrak{r}_{N}^{2}\right).$
    Updating constants yields the claim.
  3. i)
    As we have shown an analogoue inverse continuity modulus C’7), the claim follows directly following along the lines of the proof of Corollary 2.10 and is hence omitted.
Proof of 10 (Theorem 3.3).
Observing that we have shown the analogue results to the random design regression model, the proof of this theorem follows exactly along the lines of corresponding proofs in Nickl et al. [2020]. Hence, we are here only providing a scetch of the proof and ommit the details, which are let for the interested reader.
  1. i)
    By an applycatio of Eq. A.9, we obtain
    $\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\leq c_{1}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$
    While the second factor can be upper bounded by regularity estimates for $G^{D}$, the first factor is already indicating the correct exponent. Indeed, the first factor can be upper bounded by $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$. Decomposing the underlying probability space into slices
    $A_{0}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})<\delta_{N}\right\}\text{ and }A_{j}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\in(2^{j-1}\delta_{N},2^{j}\delta_{N}]\right\}$
    for some $\delta_{N}\approx\mathfrak{r}_{N}$, the corresponding expectation can be upper bounded by
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\sum_{j\in\mathbb{N}_{0}}\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\mathds{1}_{A_{j}}\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\frac{\alpha+1-\beta}{\alpha+1}}\left(1+\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\alpha^{2}\beta}\right)\right].$
    Using the upper bound of $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$ on each event $A_{j}$, as well as the concentration inequality of Theorem 3.2, we obtain
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
    which shows the claim.
  2. i)
    Similiarly, as in the proof of C’7) before, we can upper bound
    $\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\lesssim\left\|\hat{f}_{\textsc{MAP}}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}.$
    Thus, by an application of Eq. A.9 with $\beta=2$, we obtain
    $\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}\lesssim\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{1+\alpha}(\mathcal{O})}^{\frac{2}{1+\alpha}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
    The proof now follows the same idea as for the prediction error, to obtain an upper bound of the form
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha-1}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
    which proves the claim.
   Auxiliary results
The following lemma was first proven in Nickl and Wang [2024] (see Lemma 3.12) and is a key tool in the proof of Theorem 2.6.
Lemma § A.7 (Chaining for an empirical process).
Let $\Theta$ be a countable set. Let $(\Omega,\mathcal{A},\mathbb{P}^{Z})$ be a probability space. Suppose a class of real-valued measurable functions
:=$\mathcal{H}:=\left\{h_{\theta}:\Omega\to\mathbb{R}:\,\theta\in\Theta\right\}.$
Assume that $\mathcal{H}$ is uniformly bounded and has a variance envelope, namely
  1. i)
    there exists $C_{\operatorname{B}}\in\mathbb{R}_{>0}$, such that
    $\sup_{\theta\in\Theta}\left\{\sup_{\omega\in\Omega}\left\{|h_{\theta}(\omega)|\right\}\right\}\leq C_{\operatorname{B}},$
  2. i)
    there exists $\sigma^{2}\in\mathbb{R}_{>0}$, such that
    $\sup_{\theta\in\Theta}\left\{\mathbb{E}^{Z}[h_{\theta}(Z)^{2}]\right\}\leq\sigma^{2},$
    for some $Z\sim\mathbb{P}^{Z}$, where $\mathbb{E}^{Z}$ denotes the expectation operator associated to $\mathbb{P}^{Z}$.
We define the metric entropy integrals
$\mathcal{J}_{2}(\mathcal{H}):=\int_{0}^{4\sigma}\sqrt{\log N(\mathcal{H},d_{2},\rho)}\mathrm{d}\rho\text{ and }\mathcal{J}_{\infty}(\mathcal{H}):=\int_{0}^{4C_{\operatorname{B}}}\log N(\mathcal{H},d_{\infty},\rho)\mathrm{d}\rho$
with $\mathbb{L}^{2}$- and $\mathbb{L}^{\infty}$-structure given by
$d_{2}(\theta,\theta^{\prime}):=\sqrt{\mathbb{E}^{Z}[h_{\theta}(Z)-h_{\theta^{\prime}}(Z)]^{2}}\text{ and }d_{\infty}(\theta,\theta^{\prime}):=\sup_{\omega\in\Omega}\left\{|h_{\theta}(\omega)-h_{\theta^{\prime}}(\omega)|\right\},$
respectively. Let $N\in\mathbb{N}$. For i.i.d. copies $Z_{1},\dots,Z_{N}$ of $X\sim\mathbb{P}^{Z}$ and $\varepsilon_{1},\dots\varepsilon_{N}$ of $\varepsilon\sim\operatorname{N}(0,1)$ independent of $Z$, consider empirical processes either of the form
$T_{N}(\theta):=\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i})\varepsilon_{i},\,\theta\in\Theta$
or
$T_{N}(\theta):=\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}[h_{\theta}(Z)]),\,\theta\in\Theta.$
Then, there exists a universal constant $L\in\mathbb{R}_{>0}$, such that for all $x\in\mathbb{R}_{\geq 1}$, we have
$\mathbb{P}\left(\sup_{\theta\in\Theta}\left\{|T_{N}(\theta)|\right\}\geq L\left(\mathcal{J}_{2}(\mathcal{H})+\sigma\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H})+C_{\operatorname{B}}x)\right)\right)\leq 2\exp(-x).$
Similarly to Nickl and Wang [2024], whenever we apply Lemma A.7 in this work with an uncountable set $\Theta$, one can verify that the supremum can be realized as one over a countable subset of it.
Lemma § A.8 (Covering Numbers of Product Spaces).
Let $(\tilde{S},\left\|\cdot\right\|_{\tilde{S}})$ be a normed space. Assume there exists a function $\psi:\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to\mathbb{R}$, such that for any $r,\rho\in\mathbb{R}_{>0}$ the covering number $N(\mathbb{B}_{\tilde{S}}(0,r),\left\|\cdot\right\|_{\tilde{S}},\rho)=\psi(r,\rho)$ is finite. For $K\in\mathbb{N}$ define $S:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\tilde{S}$ equipped with the norm $\left\|s\right\|_{S}:=\left(\sum_{i\in\left\llbracket K\right\rrbracket}\left\|s_{i}\right\|_{\tilde{S}}^{2}\right)^{\frac{1}{2}}$ for any $s=(s_{1},\dots,s_{K})\in S$. We then have
$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\rho)\leq\psi\left(r,\frac{\rho}{\sqrt{K}}\right)^{K}.$
Proof of 11 (Lemma A.8).
By hypothesis, for all $\rho,r\in\mathbb{R}_{>0}$ there exists a covering $\mathcal{C}_{\tilde{S}}:=\{\tilde{s}^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}\subseteq\tilde{S}$ of $\mathbb{B}_{\tilde{S}}(0,r)$ with $N:=\psi(r,\rho)$, i.e. for all $\tilde{s}\in\mathbb{B}_{\tilde{S}}(0,r)$ there exists $i_{\circ}=i_{\circ}(\tilde{s})\in\left\llbracket N\right\rrbracket$ with
$\left\|\tilde{s}-\tilde{s}^{(i_{\circ})}\right\|_{\tilde{S}}\leq\rho.$
Now let $s=(s_{1},\dots,s_{K})\in\mathbb{B}_{S}(0,r)$ be arbitrary chosen, such that
$\left\|s\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}\right\|_{\tilde{S}}^{2}\leq r^{2}.$
Thus, for all $j\in\left\llbracket K\right\rrbracket$, we have $s_{j}\in\mathbb{B}_{\tilde{S}}(0,r)$. Hence, for each $s_{j}$, $j\in\left\llbracket K\right\rrbracket$, there exists $i=i_{j}\in\left\llbracket N\right\rrbracket$, such that $s_{j}\in\mathbb{B}_{\tilde{S}}(\tilde{s}^{(i_{j})},\rho)$. Defining $s^{(i_{1},\dots,i_{K})}:=(\tilde{s}^{(i_{1})},\dots,\tilde{s}^{(i_{K})})\in S$, we have
$\left\|s-s^{(i_{1},\dots,i_{K})}\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}-\tilde{s}^{(i_{j})}\right\|_{\tilde{S}}^{2}<K\rho^{2}.$
Hence, $\mathcal{C}_{S}:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\mathcal{C}_{\tilde{S}}$ forms a $\sqrt{K}\varepsilon$-covering of $\mathbb{B}_{S}(0,r)$ and thus
$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\sqrt{K}\rho)\leq N^{K}=\psi(r,\rho)^{K}.$
This shows the claim.
In the following lemma we use Lemma A.8 to derive an upper bound for the covering number of Sobolev balls.
Lemma § A.9 (Covering Numbers of Sobolev Spaces).
Let $d\in\mathbb{N}$. Let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be a bounded domain with smooth boundary $\partial\mathcal{O}$. Let $(W,\left|\cdot\right|_{W})$ be a finite-dimensional normed vector space with $\operatorname{dim}_{\mathbb{R}}W=d_{W}$.
  1. i)
    Let $\kappa\in\mathbb{R}_{\geq 0}$, $r\in\mathbb{R}_{>0}$. Assume that $\mathcal{F}$ is a subspace of $H^{\alpha}(\mathcal{O},W,r)$, if $\kappa<\frac{1}{2}$ and of $H^{\alpha}_{c}(\mathcal{O},W,r)$, otherwise. Then there exists a constant $c=c(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that
    $N(\mathcal{F},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha+\kappa}}\right)$
    for any $\rho\in\mathbb{R}_{>0}$.
  2. i)
    Let $\eta\in\mathbb{R}_{\geq 0}$ , $r\in\mathbb{R}_{>0}$. Assume that $\mathcal{F}$ is a subspace of $H^{\alpha}(\mathcal{O},W,R)$ for $\alpha>\eta+\frac{d}{2}$. Then there exists a constant $c=c(\alpha,\eta,d)\in\mathbb{R}_{>0}$, such that
    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O},W)},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha-\eta}}\right)$
    for any $\rho\in\mathbb{R}_{>0}$.
Proof of 12 (Lemma A.9).
  1. i)
    Restricting ourselfes to the case $d_{W}=1$, the claim is already shown in Nickl et al. [2020]. For general $d_{W}\in\mathbb{N}$, the claim follows by an application of Lemma A.8.
  2. i)
    We restrict ourselves to the case $d_{W}=1$. The general claim follows then by Lemma A.8. First, let $\eta=0$, such that $\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})}=\left\|\cdot\right\|_{\infty}$. If $\mathcal{O}$ corresponds to a one-dimensional interval, in chapter 4 of Giné and Nickl [2016], a covering result is shown. Using usual extension arguments, one obtains
    $N(\mathcal{F},\left\|\cdot\right\|_{\infty},\rho)\leq\exp\left(c(\alpha,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha}}\right),$
    which is also used in Nickl [2023]. Now, let $\eta\in\mathbb{R}_{>0}\setminus\mathbb{N}$. Following Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$ is compact, and by Edmunds and Triebel [1996], we further have
    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})},\rho)\leq\exp\left(c\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-\eta}}\right).$
    Now let $\eta=k\in\mathbb{N}$. According to Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is compact, as $\alpha>k+\frac{d}{2}$. Thus, for any $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O})$, we have
    $\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta}(\mathcal{O})}\leq c_{\texttt{em}}\cdot\left\|\theta_{1}-\theta_{2}\right\|_{B_{\infty 1}^{k}(\mathcal{O})}.$
    Thus, it sufficies to find an covering of $H^{\alpha}(\mathcal{O},r)$ w.r.t the $\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})}$-norm. Following the theory in Edmunds and Triebel [1996], we have
    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-k}}\right),$
    and hence
    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{C^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{c_{\texttt{em}}r}{\rho}\right)^{\frac{d}{\alpha-k}}\right).$
    The claim follows.
Lemma § A.10 (Properties of regular link functions).
Let $\Psi$ be a regular link function as in Definition 3.1. We have the following properties.
  1. i)
    Let $\theta_{1},\theta_{2}\in C^{0}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\infty}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\infty}.$
  2. i)
    Let $\theta_{1},\theta_{2}\in C^{1}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{C^{1}(\mathcal{O})}\leq c(\Psi)\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{1}(\mathcal{O})}.$
  3. i)
    Let $\theta_{1},\theta_{2}\in\mathbb{L}^{2}(\mathcal{O})$. Then
    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}.$
Proof of 13 (Lemma A.10).
Elementary computations show the claim of all three properties, since by hypothesis $\Psi$ as well as all its derivatives are bounded and hence Lipschitz continuous.

### v3

@@
-Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design
-[…]
-[…]
-Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design
-[…]
-siebel@math.uni-heidelberg.de
-Heidelberg University
-{adjustwidth}
-50pt50pt We consider the statistical inverse problem of recovering a parameter $\theta\in H^{\alpha}$ from data arising from the Gaussian regression problem
+Convergence Rates for the MAP Estimator in PDE-Regression Models with Random Design
+Abstract
+We consider the statistical inverse problem of recovering a parameter $\theta\in H^{\alpha}$ from data arising from the Gaussian regression problem
 $Y=\mathscr{G}(\theta)(Z)+\varepsilon$
-with nonlinear forward map $\mathscr{G}:\mathbb{L}^{2}\to\mathbb{L}^{2}$, random design points $Z$ and Gaussian noise $\varepsilon$.
-The estimation strategy is based on a least squares approach under $\left\|\cdot\right\|_{H^{\alpha}}$-constraints.
+with nonlinear forward map $\mathscr{G}:\mathbb{L}^{2}\to\mathbb{L}^{2}$, random design point $Z$ and Gaussian noise $\varepsilon$.
+The estimation strategy is based on a least squares approach under $\|\cdot\|_{H^{\alpha}}$-constraints.
 We establish the existence of a least squares estimator $\hat{\theta}$ as a maximizer for a given functional under Lipschitz-type assumptions on the forward map $\mathscr{G}$.
@@
 Additionally, we briefly discuss the applicability of the general model to other problems.
-1   1Introduction
-Overview: In this paper, we consider a nonlinear Gaussian regression model with random design, i.e., for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma\in\mathbb{R}_{>0}$, we have access to data $(Y_{i},X_{i})_{i=1}^{N}$ arising from
-$Y_{i}=u_{f}(X_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.1)
+1 Introduction
+In this paper, we consider a nonlinear Gaussian regression model with random design, i.e., for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma^{2}\in\mathbb{R}_{>0}$, we have access to data $(Y_{i},X_{i})_{i=1}^{N}$ arising from
+$Y_{i}=u_{f}(X_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1)
 where $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d.
@@
 The goal is to present an estimation strategy for $f$ by using the data set $(Y_{1},X_{1}),\dots,(Y_{N},X_{N})$, which is an example of a nonlinear statistical inverse problem as $f$ is not observed directly, see, e.g.
-Nickl [2023].
+Nickl (2023).
 Many real-world phenomena, like physical and biological laws, fit into this framework as they are often described by (non-)linear dynamical systems and partial differential equations which exhibit spatial and/or temporal dependencies.
-Examples range from the description of cell movements (see Tokihiro [2021]) and geophysical laws describing the atmosphere (see Kalnay [2002]), to ocean dynamics (see Bennett [2002]), turbulence (see Majda and Harlim [2012]), and fluid flows (see Cotter et al.
-[2009]).
-Moreover, this setting aligns with various frameworks from the perspective of inverse problems: examples include electrical impedance tomography (see Calderón [2006] and Isaacson et al.
-[2004]), photoacoustic tomography and other hybrid imaging problems (see Kuchment and Kunyansky [2015], Bal [2014], and particularly Bal and Uhlmann [2010] for a general framework), inverse scattering (see Colton and Kress [2019] and Hohage and Weidling [2015]), interacting particle models (see Nickl et al.
-[2024]), and reaction diffusion equations (see Nickl [2024]).
-In this work, we consider the so-called Darcy Problem (see, for instance Stuart [2010]), where we refer to Bonito et al.
-[2017] and Hohage and Pricop [2008], and the references therein, for concrete application examples.
+Examples range from the description of cell movements (see Tokihiro (2021)) and geophysical laws describing the atmosphere (see Kalnay (2002)), to ocean dynamics (see Bennett (2002)), turbulence (see Majda and Harlim (2012)), and fluid flows (see Cotter et al.
+(2009)).
+Moreover, this setting aligns with various frameworks from the perspective of inverse problems: examples include electrical impedance tomography (see Calderón (2006) and Isaacson et al.
+(2004)), photoacoustic tomography and other hybrid imaging problems (see Kuchment and Kunyansky (2015); Bal (2014), and particularly Bal and Uhlmann (2010) for a general framework), inverse scattering (see Colton and Kress (2019) and Hohage and Weidling (2015)), interacting particle models (see Nickl et al.
+(2024)), and reaction diffusion equations (see Nickl (2024)).
+In this work, we consider the so-called Darcy Problem (see, for instance Stuart (2010)), where we refer to Bonito et al.
+(2017a) and Hohage and Pricop (2008), and the references therein, for concrete application examples.
 In this situation $u_{f}:\mathcal{O}\to\mathbb{R}$ describes a steady-state heat flow, i.e.
 the unique solution to the elliptic PDE
-$\begin{dcases}\hfill\operatorname{div}(f\cdot\nabla u)=g,\text{ on $\mathcal{O}$},\\ \hfill u=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (1.2)
+$\begin{dcases}\hfill\operatorname{div}(f\cdot\nabla u)&=g,\text{ on $\mathcal{O}$},\\ \hfill u\hskip 2.84526pt&=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (2)
 with $\nabla$ denoting the gradient of a function and $\operatorname{div}$ the divergence operator applied to a vector field, respectively.
-Here $g>0$ is a heat source function on $\mathcal{O}$, which is known in advance, and $f:\mathcal{O}\to(0,\infty)$ describes an unknown conductivity (or diffusion) coefficient function.
+Here, $g>0$ is a heat source function on $\mathcal{O}$, which is known in advance, and $f:\mathcal{O}\to(0,\infty)$ describes an unknown conductivity (or diffusion) coefficient function.
 In real-world observation schemes, idealized observations that exactly match the mathematical formalism of a PDE are typically unavailable, as measurement errors introduce statistical noise, motivating the observation scheme eq.
-1.1 in this work.
+1 in this work.
 More precisely, the data consists of two parts - the exact solution $u_{f}$ of the Darcy problem eq.
-1.2, corrupted by the second part - the Gaussian noise.
+2, corrupted by the second part - the Gaussian noise.
 Nonetheless, the goal remains the same: to reconstruct the underlying unknown coefficient function $f$ that fits the pattern of eq.
-1.2.
+2.
 In the present work, we follow the approaches in Monard et al.
-[2020] and Nickl et al.
-[2020] and embed the regression model Eq.
-1.1 into a more generic setting.
+(2020) and Nickl et al.
+(2020) and embed the regression model eq.
+1 into a more generic setting.
 Specifically, given a probability space $(\mathcal{Z},\mathscr{Z},\zeta)$ and a possibly nonlinear operator (forward map) $\mathscr{G}:\mathbb{L}^{2}(\mathcal{O})\to\mathbb{L}_{\zeta}^{2}(\mathcal{Z})$, we assume to have access to observations $(Y_{i},Z_{i})_{i=1}^{N}$, which arise from the equation
-$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (1.3)
+$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i=1,\dots,N,$ (3)
 where, in this situation, $Z_{1},\dots,Z_{N}$ are i.i.d.
@@
 Note that eq.
-1.1 corresponds to the special case where $\mathscr{G}$ is the nonlinear solution map $f\mapsto u_{f}$ of the Darcy problem.
+1 corresponds to the special case where $\mathscr{G}$ is the nonlinear solution map $f\mapsto u_{f}$ of the Darcy problem.
 The reconstruction of parameters $\theta$ following an indirect and noisy observation scheme as in eq.
-1.3 has a long mathematical history.
+3 has a long mathematical history.
 For an extensive overview of models with deterministic noise and a linear relationship $\theta\mapsto\mathscr{G}(\theta)$, we refer to Engl et al.
-[1996].
+(1996).
 In the last twenty years a wide field has been established to also consider nonlinear and statistical inverse problems.
-We refer to Benning and Burger [2018], Kaltenbacher et al.
-[2008], Dashti et al.
-[2013], Dashti and Stuart [2015], Stuart [2010], Nickl [2023], Arridge et al.
-[2019], Bissantz et al.
-[2007], and Kaipio and Somersalo [2005], as well as to the references therein.
+We refer to Benning and Burger (2018); Kaltenbacher et al.
+(2008); Dashti et al.
+(2013); Dashti and Stuart (2015); Stuart (2010); Nickl (2023); Arridge et al.
+(2019); Bissantz et al.
+(2007), and Kaipio and Somersalo (2005), as well as to the references therein.
 For a novel proof in the Bayesian approach within this PDE-constrained regression setting, we refer to Monard et al.
-[2020] and Monard et al.
-[2021], where statistical guarantees, especially for the posterior mean, are derived under Gaussian process priors on the parameter of interest $\theta$ and analytical assumptions on the forward map $\mathscr{G}$.
+(2020) and Monard et al.
+(2021), where statistical guarantees, especially for the posterior mean, are derived under Gaussian process priors on the parameter of interest $\theta$ and analytical assumptions on the forward map $\mathscr{G}$.
 The applicability to PDE models is illustrated not only for the inversion of non-Abelian X-ray transforms but also for a model as in eq.
-1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the stationary Schrödinger equation.
-Giordano and Nickl [2020] adopt these ideas for the Darcy problem Eq.
-1.2 and provide results on posterior convergence.
-General assumptions to show posterior convergence are formulated in Nickl [2023].
+1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the stationary Schrödinger equation.
+Giordano and Nickl (2020) adopt these ideas for the Darcy problem eq.
+2 and provide results on posterior convergence.
+General assumptions to show posterior convergence are formulated in Nickl (2023).
 In Nickl et al.
-[2020], the authors propose a (generalized) penalized least squares estimator $\hat{\theta}$ for $\theta$ in a Gaussian white noise model, which can be interpreted as minimizer of the quantity
-$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\left\|Y-\mathscr{G}(\theta)\right\|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|^{2}_{H^{\alpha}(\mathcal{O})}\in\mathbb{R}_{\geq 0}.$
+(2020), the authors propose a (generalized) penalized least squares estimator $\hat{\theta}$ for $\theta$ in a Gaussian white noise model, which can be interpreted as minimizer of the quantity
+$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\|Y-\mathscr{G}(\theta)\|^{2}+\mathfrak{r}^{2}\|\theta\|^{2}_{H^{\alpha}(\mathcal{O})}\in\mathbb{R}_{\geq 0}.$
 They make use of M-estimation techniques to derive statistical guarantees for its prediction loss.
 The results are then applied for the special case of considering eq.
-1.1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the Darcy problem Eq.
-1.2 and also for the time-independent Schrödinger equation.
+1 with $u_{f}:\mathcal{O}\to\mathbb{R}$ being the unique solution of the Darcy problem eq.
+2 and also for the time-independent Schrödinger equation.
 In particular, they provide upper bounds for the prediction loss and show that these are in some situations minimax-optimal and coincide with rates known for linear operators.
-For the proof, the mathematical formulation of the observation scheme via a white noise process is very suitable, however, this scheme is rather infeasible for real-world applications.
-Furthermore, the authors discuss the relationship between $\hat{\theta}$ and the Bayesian approach, where $\hat{\theta}$ can also be interpreted as maximizer of the likelihood $\pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$ of the posterior distribution $\Pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$, under a Gaussian process priors $\Pi(\cdot)$ on $\theta$.
-This also results in the naming of $\hat{\theta}$ as maximum a posteriori (MAP) estimator.
-In contrast, we aim to prove results in the more realistic i.i.d.
-regression model Eq.
-1.3.
-For random design regression models described by Eq.
-1.3, Bohr and Nickl [2023] and Nickl and Wang [2024] provide statistical and computational guarantees for a version of the maximum a posteriori estimator (MAP) $\hat{\theta}$, defined as the $D$-dimensional minimizer of the quantity
-$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\frac{1}{N}\sum_{i=1}^{N}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|^{2}+\mathfrak{r}^{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}^{2}\in\mathbb{R}_{>0}.$ (1.4)
+While the white noise model provides a mathematically convenient formulation for theoretical analysis, particularly in deriving statistical guarantees, it is often infeasible in real-world applications.
+More realistic observation schemes involve discrete measurements at deterministic design points, which can be chosen, for instance, to optimally cover the domain under certain regularity assumptions (see Daon (2025) and Teckentrup (2020) for instance).
+In our framework, we instead consider a random design regression model, which is analytically convenient and aligns well with established approaches in nonparametric statistical theory: For random design regression models described by eq.
+3, Bohr and Nickl (2023) and Nickl and Wang (2024) provide statistical and computational guarantees for a Tikhonov-regularized estimator, defined as the $D$-dimensional minimizer of the quantity
+$H^{\alpha}(\mathcal{O})\ni\theta\mapsto\frac{1}{N}\sum_{i=1}^{N}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|^{2}+\mathfrak{r}^{2}\|\theta\|_{H^{\alpha}(\mathcal{O})}^{2}\in\mathbb{R}_{>0}.$ (4)
 Under fairly general assumptions on the forward map $\mathscr{G}$, the authors demonstrate the consistency and stability of $\hat{\theta}$, provided that $\theta$ can be sufficiently well approximated by its orthogonal projection onto a finite-dimensional subspace.
 The key of the proof is an empirical process argument, which implicitly requires the forward map $\mathscr{G}$ to satisfy the ’global’ stability estimates
-$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\mathbb{L}^{2}}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\mathbb{L}^{2}}\text{ and }\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty}$ (1.5)
+$\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\mathbb{L}^{2}}\lesssim\|\theta-\theta^{\prime}\|_{\mathbb{L}^{2}}\text{ and }\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\infty}\lesssim\|\theta-\theta^{\prime}\|_{\infty}$ (5)
 uniformly over the whole parameter space.
-While the solution map of the time-independent Schrödinger equation meets these requirements, there is no known analogous result for the second condition for the solution map of the Darcy problem, which renders the theory of Bohr and Nickl [2023] and Nickl and Wang [2024] inapplicable.
+While the solution map of the time-independent Schrödinger equation meets these requirements, there is no known analogous result for the second condition for the solution map of the Darcy problem, which renders the theory of Bohr and Nickl (2023) and Nickl and Wang (2024) inapplicable.
 In addition, the first requirement leads to slower convergence rates, which we improve by taking the ill-posedness of the forward map $\mathscr{G}$ into account (see Nickl et al.
-[2020] for corresponding results in the white noise model).
+(2020) for corresponding results in the white noise model).
+Furthermore, the authors discuss the relationship between $\hat{\theta}$ and the Bayesian approach for inverse problems, where $\hat{\theta}$ can also be interpreted as maximizer of the likelihood $\pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$ of the posterior distribution $\Pi(\cdot|(Y_{i},Z_{i})_{i=1}^{N})$, under a Gaussian process prior $\Pi(\cdot)$ on $\theta$.
+This also results in the naming of $\hat{\theta}$ as maximum a posteriori (MAP) estimator.
 In section 2.2.
-of Nickl and Wang [2024], the authors propose conditions under which the stochastic gradient algorithm
-$\vartheta_{0}:=\theta_{\operatorname{init}},\hskip 4.30554pt\vartheta_{k+1}=\vartheta_{k}+\rho\nabla\log\tilde{\pi}(\vartheta_{k}|(Y_{i},Z_{i})_{i=1}^{N})$ (1.6)
+of Nickl and Wang (2024), the authors propose conditions under which the stochastic gradient algorithm
+$\vartheta_{0}:=\theta_{\operatorname{init}},\hskip 4.30554pt\vartheta_{k+1}=\vartheta_{k}+\rho\nabla\log\tilde{\pi}(\vartheta_{k}|(Y_{i},Z_{i})_{i=1}^{N})$ (6)
 converges geometrically to the minimizer $\hat{\theta}$ of eq.
-1.4 (and hence to the ground truth $\theta$ after sufficiently many iterations) in polynomial time, if the initial value $\theta_{\operatorname{init}}\in\mathbb{R}^{D}$ is chosen appropriately.
+4 (and hence to the ground truth $\theta$ after sufficiently many iterations) in polynomial time, if the initial value $\theta_{\operatorname{init}}\in\mathbb{R}^{D}$ is chosen appropriately.
 We refer to section 2.2.5.
-(and particularly Theorem 2.8) in Nickl and Wang [2024].
+(and particularly Theorem 2.8) in Nickl and Wang (2024).
 The results are then applied to $\mathscr{G}$ being the solution map $f\mapsto u_{f}$ of the stationary Schrödinger equation, which meets global stability conditions eq.
-1.5.
-Under more general assumptions, Nickl [2023] discusses situations in which the analysis of the local concavity of the associated log-likelihood to Eq.
-1.3 is sufficient to specify a stochastic gradient algorithm, for which $\hat{\theta}$ is computable in polynomial time, if the underlying parameter space is finte-dimensional.
-Under slightly technical changes, the methods also seem to apply to the Darcy problem.
-However, we refrain from discussing the computational aspects in more detail in this work due to the similarities with Nickl [2023] and Nickl and Wang [2024].
+5.
+Under more general assumptions, Nickl (2023) discusses situations in which the analysis of the local concavity of the associated log-likelihood to eq.
+3 is sufficient to specify a stochastic gradient algorithm, for which $\hat{\theta}$ is computable in polynomial time, if the underlying parameter space is finte-dimensional.
+Building on the proof strategy of Nickl and Wang (2024), the authors of Altmeyer (2024) establish nonasymptotic sampling guarantees for a gradient-based MCMC algorithm relative to a surrogate posterior in high-dimensional generalized linear models (see Section 2.4), a framework that also encompasses the Darcy problem discussed in Section 3.4.
+Hence, we refrain from discussing the computational aspects in more detail in this work due to the similarities with Nickl (2023); Nickl and Wang (2024) and Altmeyer (2024).
 A different approach of modeling real-world phenomena utilizes the description of dynamical (time-dependent) systems via stochastic partial differential equations (SPDEs), which inherently take into account the noise in the system dynamics.
@@
 For an overview of recent work of SPDEs we refer to Altmeyer et al.
-[2023], Altmeyer and Reiß [2021] and Gaudlitz and Reiß [2023] as well as the references therein.
+(2023); Altmeyer and Reiß (2021) and Gaudlitz and Reiß (2023) as well as the references therein.
 We emphasize that SPDEs are not the focus of the present work.
 In the present paper, we follow the approach of Monard et al.
-[2020], Monard et al.
-[2021] and Nickl et al.
-[2020] and consider the general nonlinear Gaussian regression model with random design (as Eq.
-1.3) and define a maximum a posteriori estimator, similar to Eq.
-1.4.
+(2020, 2021) and Nickl et al.
+(2020) and consider the general nonlinear Gaussian regression model with random design (as eq.
+3) and define a maximum a posteriori estimator, similar to eq.
+4.
 We introduce conditions on the forward map $\mathscr{G}$, generalizing eq.
-1.5 by replacing the right-hand sides in both stability estimates with $(H^{\kappa})^{*}$- and $C^{\eta}$-norms, respectively, while also allowing polynomial growth.
+5 by replacing the right-hand sides in both stability estimates with $(H^{\kappa})^{*}$- and $C^{\eta}$-norms, respectively, while also allowing polynomial growth.
 We present a general concentration result, where we compare the prediction loss $\mathfrak{d}^{2}(\hat{\theta},\theta)$ with some proxy $\mathfrak{d}^{2}(\tilde{\theta},\theta)$ (interpretable as approximation error) and obtain a corresponding consistency result for the MAP estimate, whenever the proxy approximates the ground truth sufficiently well.
-This result extends the findings of Bohr and Nickl [2023] and Nickl and Wang [2024], as we allow a generic reference function $\tilde{\theta}$, rather than only a finite-dimensional approximation of the ground truth and provide faster convergence rates for the squared prediction loss, which are now upper-bounded by $N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}$, where $\kappa\in\mathbb{R}_{>0}$ indicates the degree of ill-posedness of the underlying inverse problem.
+This result extends the findings of Bohr and Nickl (2023) and Nickl and Wang (2024), as we allow a generic reference function $\tilde{\theta}$, rather than only a finite-dimensional approximation of the ground truth and provide faster convergence rates for the squared prediction loss, which are now upper-bounded by $N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}$, where $\kappa\in\mathbb{R}_{>0}$ indicates the degree of ill-posedness of the underlying inverse problem.
 Under an additional inverse stability assumption on the forward map $\mathscr{G}$, an analogous result for the estimation loss of $\hat{\theta}$ is established.
 For the sake of completeness, we provide an existence result for the MAP-estimator in this random design regression model by exploiting the direct method in the calculus of variations, where we adapt a corresponding proof of Nickl et al.
-[2020] for the white noise model.
+(2020) for the white noise model.
 The general theory is applied on the Darcy problem.
@@
 In this particular situation, we provide consistency for the MAP estimator, as well as corresponding convergence rate for prediction and estimation errors, which fills a gap in existing theory.
+x
 Organisation of the paper: In section 2 we introduce the most essential notation and the statistical model in detail.
@@
 We provide analogously consistency for the MAP-estimator, as well as upper bounds for the prediction and estimation risk.
-In Appendix A, we provide detailed proofs for the results in the main part and further helping results.
-Acknowledgement: This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1-39090098 (the Heidelberg STRUCTURES Cluster of Excellence).
-The author thanks Richard Nickl and Jan Johannes for their time, patience, and fruitful discussions, as well as for the visit to Richard Nickl’s research group in Cambridge in November 2023, which marked the beginning of this work.
-2   2A general statistical inverse Problem
+In section 4, we provide detailed proofs for the results in the main part and further helping results.
+2 A general statistical inverse Problem
 Notation § 2.1 (Preliminary).
-We start with some basic notation that we use throughout the whole paper.
+Throughout the whole paper, we use the following notation.
   1. i)
-    For all $\mathbb{K}\in\{\mathbb{N},\mathbb{R}\}$ and $a\in\mathbb{R}$ set $\mathbb{K}_{\geq a}:=\mathbb{K}\cap[a,\infty)$ and $\mathbb{K}_{>a}:=\mathbb{K}\cap(a,\infty)$.
+    For $\mathbb{K}\in\{\mathbb{N},\mathbb{R}\}$ and $a\in\mathbb{R}$ set $\mathbb{K}_{\geq a}:=\mathbb{K}\cap[a,\infty)$ and $\mathbb{K}_{>a}:=\mathbb{K}\cap(a,\infty)$.
 Analogously, $\mathbb{K}_{\leq a}$ and $\mathbb{K}_{<a}$ are defined.
-In particular, we set $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
+We set $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
   2. ii)
-    For all $a,b\in\mathbb{Z}$ with $a\leq b$ set $\left\llbracket a,b\right\rrbracket:=[a,b]\cap\mathbb{Z}$ and further $\left\llbracket n\right\rrbracket:=\left\llbracket 1,n\right\rrbracket$ for any $n\in\mathbb{N}$.
+    For all $a,b\in\mathbb{Z}$ with $a\leq b$ set $\left\llbracket a,b\right\rrbracket:=[a,b]\cap\mathbb{Z}$, and $\left\llbracket n\right\rrbracket:=\left\llbracket 1,n\right\rrbracket$ for any $n\in\mathbb{N}$.
   3. iii)
@@
   5. v)
-    For two normed spaces $(S_{1},\left\|\cdot\right\|_{S_{1}})$ and $(S_{2},\left\|\cdot\right\|_{S_{2}})$, we write $S_{1}\hookrightarrow S_{2}$, if $S_{1}\subseteq S_{2}$ is continuously embedded into $S_{2}$.
+    For two normed spaces $(S_{1},\|\cdot\|_{S_{1}})$ and $(S_{2},\|\cdot\|_{S_{2}})$, we write $S_{1}\hookrightarrow S_{2}$, if $S_{1}$ is continuously embedded into $S_{2}$.
   6. vi)
-    For $d\in\mathbb{N}$ denote by $\left\|\cdot\right\|_{\mathbb{R}^{d}}$ the usual Euclidean norm on $\mathbb{R}^{d}$.
+    For $d\in\mathbb{N}$ denote by $\|\cdot\|_{\mathbb{R}^{d}}$ the usual Euclidean norm on $\mathbb{R}^{d}$.
   7. vii)
-    For a function $f:\mathcal{S}\to\mathbb{R}$ defined on an arbitrary space $\mathcal{S}$, we denote by $\left\|f\right\|_{\infty}:=\sup_{s\in\mathcal{S}}\left\{|f(s)|\right\}$ the supremum norm of $f$ over $\mathcal{S}$.
-We do not highlight the underlying space, if it is clear from the context.
-  8. viii)
     Universal constants are denoted by $c$.
@@
 If not further mentioned, the value of a constant can change from line to line.
-   Hölder and Sobolev spaces
+2.1 Hölder and Sobolev spaces
 Throughout let $(\mathcal{Z},\mathscr{Z},\zeta)$ be a measure space.
-For $p\in[1,\infty]$ we define the usual space $\mathbb{L}_{\zeta}^{p}(\mathcal{Z}):=\mathbb{L}^{p}(\mathcal{Z},\mathscr{Z},\zeta)$ of (equivalence classes of) functions $f:(\mathcal{Z},\mathscr{Z})\to(\mathbb{R},\mathscr{B}_{\mathbb{R}})$ endowed with the norm $\left\|\cdot\right\|_{\mathbb{L}^{p}_{\zeta}(\mathcal{Z})}$.
-Let $(V,\left|\cdot\right|_{V})$ be a finite-dimensional normed $\mathbb{R}$-vector space with dimension $d_{V}:=\operatorname{dim}_{\mathbb{R}}(V)$.
+For $p\in[1,\infty]$ we define the usual space $\mathbb{L}_{\zeta}^{p}(\mathcal{Z}):=\mathbb{L}^{p}(\mathcal{Z},\mathscr{Z},\zeta)$ of (equivalence classes of) functions $f:(\mathcal{Z},\mathscr{Z})\to(\mathbb{R},\mathscr{B}_{\mathbb{R}})$ endowed with the norm $\|\cdot\|_{\mathbb{L}^{p}_{\zeta}(\mathcal{Z})}$.
+Let $(V,|\cdot|_{V})$ be a finite-dimensional normed $\mathbb{R}$-vector space with dimension $d_{V}:=\operatorname{dim}_{\mathbb{R}}(V)$.
 By identifying $V$ canonically with $\mathbb{R}^{d_{V}}$, we define the space of all vector fields $f=(f_{1},\dots,f_{d_{V}}):\mathcal{Z}\to V$ with the integrable components $f_{i}\in\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$, $i\in\left\llbracket d_{V}\right\rrbracket$, as the tensor space $\mathbb{L}^{p}_{\zeta}(\mathcal{Z},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$.
@@
 where $\cdot_{i}$ indicates the $i$-th component.
+Moreover, for any function $f:\mathcal{Z}\to V$, we denote by $\|f\|_{\infty}:=\sup_{z\in\mathcal{Z}}|f(z)|_{V}$ the uniform norm of $f$ over $\mathcal{Z}$.
+We do not highlight the underlying space, if it is clear from the context.
 Throughout let $d\in\mathbb{N}$.
@@
 By convention, we denote by $C^{0}(\mathcal{O})$ the space of continuous functions on $\mathcal{O}$.
-We extend the definition for any $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by defining $C^{s}(\mathcal{O})$ as the function space of all maps $f:\mathcal{O}\to\mathbb{R}$, such that for all multi-indices $\bm{\beta}\in\mathbb{N}_{0}^{d}$ with $|\bm{\beta}|\leq\lfloor s\rfloor$ the integer part of $s$ the derivative $D^{\bm{\beta}}f$ exists and is $s-\lfloor s\rfloor$-Hölder continuous.
+We extend the definition for any $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by defining $C^{s}(\mathcal{O})$ as the function space of all maps $f:\mathcal{O}\to\mathbb{R}$, such that for all multi-indices $\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}$ with $|\boldsymbol{\beta}|\leq\lfloor s\rfloor$ - the integer part of $s$ - the derivative $D^{\boldsymbol{\beta}}f$ exists and is $s-\lfloor s\rfloor$-Hölder continuous.
 A norm on $C^{s}(\mathcal{O})$ is given by
-$\left\|f\right\|_{C^{s}(\mathcal{O})}:=\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|\leq\lfloor s\rfloor}\left\|D^{\bm{\beta}}f\right\|_{\infty}+\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|=\lfloor s\rfloor}\sup_{x,y\in\mathcal{O}:x\neq y}\left\{\frac{|D^{\bm{\beta}}f(x)-D^{\bm{\beta}}f(y)|}{\left\|x-y\right\|_{\mathbb{R}^{d}}^{s-\lfloor s\rfloor}}\right\}.$
+$\|f\|_{C^{s}(\mathcal{O})}:=\sum_{\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}:|\boldsymbol{\beta}|\leq\lfloor s\rfloor}\|D^{\boldsymbol{\beta}}f\|_{\infty}+\sum_{\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}:|\boldsymbol{\beta}|=\lfloor s\rfloor}\sup_{x,y\in\mathcal{O}:x\neq y}\left\{\frac{|D^{\boldsymbol{\beta}}f(x)-D^{\boldsymbol{\beta}}f(y)|}{\|x-y\|_{\mathbb{R}^{d}}^{s-\lfloor s\rfloor}}\right\}.$
 A function $f:\mathcal{O}\to\mathbb{R}$ is called smooth, if it belongs to $C^{\infty}(\mathcal{O}):=\bigcap_{s\in\mathbb{R}_{>0}}C^{s}(\mathcal{O})$.
@@
 For $s\in\mathbb{N}$, we denote by $H^{s}(\mathcal{O}):=H^{s}(\mathcal{O},\mathbb{R})$ the usual order-$s$ Sobolev space of weakly differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with square integrable weak derivatives up to order $s$.
-By $\left\|\cdot\right\|_{H^{s}(\mathcal{O})}$, we denote its usual norm.
-This definition extends to positive $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by interpolation, see Triebel [1983].
+By $\|\cdot\|_{H^{s}(\mathcal{O})}$, we denote its usual norm.
+This definition extends to positive $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by interpolation, see Triebel (1983).
 By convention, set $H^{0}(\mathcal{O}):=\mathbb{L}^{2}(\mathcal{O})$.
-Further define $H^{s}_{c}(\mathcal{O}):=\overline{C_{c}^{\infty}(\mathcal{O})}^{\left\|\cdot\right\|_{H^{s}(\mathcal{O})}}$ as well as $\bar{H}^{s}(\mathcal{O})$ to be the space of all functions in $H^{s}(\mathcal{O})$, whose support is $\bar{\mathcal{O}}$.
-If $s\neq k+\frac{1}{2}$ for some $k\in\mathbb{N}$, we have $\bar{H}^{s}(\mathcal{O})=H^{s}_{c}(\mathcal{O})$.
+Further define $H^{s}_{c}(\mathcal{O}):=\overline{C_{c}^{\infty}(\mathcal{O})}^{\|\cdot\|_{H^{s}(\mathcal{O})}}$ as well as $\bar{H}^{s}(\mathcal{O})$ to be the space of all functions in $H^{s}(\mathcal{O})$, whose support is $\bar{\mathcal{O}}$.
+If $s\neq k+\frac{1}{2}$ for any $k\in\mathbb{N}$, we have $\bar{H}^{s}(\mathcal{O})=H^{s}_{c}(\mathcal{O})$.
 Negative-order Sobolev spaces are defined as the topological dual of $\bar{H}^{s}(\mathcal{O})$, i.e., $H^{-s}(\mathcal{O}):=H^{-s}(\mathcal{O},\mathbb{R}):=\left(\bar{H}^{s}(\mathcal{O})\right)^{*}$.
-Denoting by $\operatorname{tr}_{s}:H^{s}(\mathcal{O})\to\bigtimes_{l=0}^{s-1}H^{s-l-\frac{1}{2}}(\mathcal{O})$ the higher-order trace operator for any $s\in\mathbb{N}$, we have $H^{s}_{0}(\mathcal{O}):=\left\{f\in H^{s}(\mathcal{O}):\,\operatorname{tr}_{s}[f]=0\right\}=H^{s}_{c}(\mathcal{O})$, where we refer to Lions and Magenes [1972] for detailed definitions and properties.
-All of the previous spaces of functions defined on $\mathcal{O}$ with values in $\mathbb{R}$ can be extended to spaces of the corresponding vector fields $f:\mathcal{O}\to V$ in the same manner as before.
-Indeed if $F(\mathcal{O})$ is one of the above function spaces, we set $F(\mathcal{O},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}F(\mathcal{O})$.
-If $F(\mathcal{O})$ is equipped with some norm $\left\|\cdot\right\|_{F(\mathcal{O})}$, a norm on $F(\mathcal{O},V)$ is given by $\left\|\cdot\right\|_{F(\mathcal{O},V)}^{2}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\left\|\cdot_{i}\right\|_{F(\mathcal{O})}^{2}$.
-Similarly, if $F(\mathcal{O})$ is endowed with a Hilbert space inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O})}$, $F(\mathcal{O},V)$ inherits the Hilbert space structure with the inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O},V)}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{F(\mathcal{O})}$.
-Further, for simplicity, we write for centered balls with radius $r\in\mathbb{R}_{>0}$ in function spaces $F(\mathcal{O})$ and $F(\mathcal{O},V)$ $F(\mathcal{O},r):=\mathbb{B}_{F(\mathcal{O})}(0,r)$ and $F(\mathcal{O},V,r):=\mathbb{B}_{F(\mathcal{O},V)}(0,r)$, respectively.
-   The measurement model
-Throughout let $(W,\left|\cdot\right|_{W})$ be another finite-dimensional and normed $\mathbb{R}$-vector space with dimension $d_{W}:=\operatorname{dim}_{\mathbb{R}}(W)\in\mathbb{N}$, such that we canonically identify $W$ with $\mathbb{R}^{d_{W}}$.
+Denoting by $\operatorname{tr}_{s}:H^{s}(\mathcal{O})\to\bigtimes_{l=0}^{s-1}H^{s-l-\frac{1}{2}}(\mathcal{O})$ the higher-order trace operator for any $s\in\mathbb{N}$, we have $H^{s}_{0}(\mathcal{O}):=\left\{f\in H^{s}(\mathcal{O}):\,\operatorname{tr}_{s}[f]=0\right\}=H^{s}_{c}(\mathcal{O})$, where we refer to Lions and Magenes (1972) for detailed definitions and properties.
+Let $(W,\left|\cdot\right|_{W})$ be another finite-dimensional and normed $\mathbb{R}$-vector space with dimension $d_{W}:=\operatorname{dim}_{\mathbb{R}}(W)\in\mathbb{N}$, such that we canonically identify $W$ with $\mathbb{R}^{d_{W}}$.
+All of the previous spaces of functions defined on $\mathcal{O}$ with values in $\mathbb{R}$ can be extended to spaces of the corresponding vector fields $f:\mathcal{O}\to W$ in the same manner as before.
+Indeed, if $F(\mathcal{O})$ is one of the above function spaces, we set $F(\mathcal{O},W):=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}F(\mathcal{O})$.
+If $F(\mathcal{O})$ is equipped with some norm $\|\cdot\|_{F(\mathcal{O})}$, a norm on $F(\mathcal{O},W)$ is given by $\|\cdot\|_{F(\mathcal{O},W)}^{2}:=\sum_{i\in\left\llbracket d_{W}\right\rrbracket}\|\cdot_{i}\|_{F(\mathcal{O})}^{2}$.
+Similarly, if $F(\mathcal{O})$ is endowed with a Hilbert space inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O})}$, $F(\mathcal{O},W)$ inherits the Hilbert space structure with the inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O},W)}:=\sum_{i\in\left\llbracket d_{W}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{F(\mathcal{O})}$.
+Further, for simplicity, we write for centered balls with radius $r\in\mathbb{R}_{>0}$ in function spaces $F(\mathcal{O})$ and $F(\mathcal{O},W)$, $F(\mathcal{O},r):=\mathbb{B}_{F(\mathcal{O})}(0,r)$ and $F(\mathcal{O},W,r):=\mathbb{B}_{F(\mathcal{O},W)}(0,r)$, respectively.
+2.2 The measurement model
 Assuming that $(\mathcal{Z},\mathscr{Z},\zeta)$ is a probability space, we are interested in a (possibly nonlinear) forward map
@@
 for a given sample size $N\in\mathbb{N}$ and a fixed variance $\sigma^{2}\in\mathbb{R}_{>0}$, we observe $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from the equation
-$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i\in\left\llbracket N\right\rrbracket,\,\theta\in\tilde{\Theta},$ (2.1)
+$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i\in\left\llbracket N\right\rrbracket,\,\theta\in\tilde{\Theta}$, (7)
 where $Z_{1},\dots,Z_{N}$ are i.i.d copies of $Z$, which is drawn from $\zeta$.
@@
 The design variable $Z$ and the error $\varepsilon$ are assumed to be independent.
+Note, we assume that the noise variance $\sigma^{2}$ is known in advance.
+In settings where $\sigma^{2}$ is unknown, it could be estimated from the data, for instance via the sample variance.
+A detailed treatment of a unknown variance setting is beyond the scope of the present work and posponed to future work.
 The distribution of a generic observation $D:=(Y,Z)$ arising from eq.
-2.1 is denoted by $\mathbb{P}_{\theta}$.
+7 is denoted by $\mathbb{P}_{\theta}$.
 The joint product law of $D_{N}$ on $(V\times\mathcal{Z})^{N}$ is denoted as $\mathbb{P}_{\theta}^{N}=\bigotimes_{i\in\left\llbracket N\right\rrbracket}\mathbb{P}_{\theta}$.
@@
 Keeping the identification of $V$ with $\mathbb{R}^{d_{V}}$ in mind, $\mathbb{P}_{\theta}$ has a probability density function $p_{\theta}$ with respect to the product measure $\mathcal{L}^{d_{V}}\otimes\zeta$, namely
-$\forall y\in V,z\in\mathcal{Z}:\,p_{\theta}(y,z)=(2\pi\sigma^{2})^{-\frac{d_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\left|y-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right).$
+$\forall y\in V,z\in\mathcal{Z}:\,p_{\theta}(y,z)=(2\pi\sigma^{2})^{-\frac{d_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}|y-\mathscr{G}(\theta)(z)|_{V}^{2}\right).$
 Consequently, $\mathbb{P}_{\theta}^{N}$ also has a probability density function with respect to $\bigotimes_{i\in\left\llbracket N\right\rrbracket}(\mathcal{L}^{d_{V}}\otimes\zeta)$ denoted by $p_{\theta}^{N}$, which reads for all $(y,z)\in(V\times\mathcal{Z})^{N}$ as
-$p_{\theta}^{N}(y,z)=\prod_{i\in\left\llbracket N\right\rrbracket}p_{\theta}(y_{i},z_{i})=(2\pi\sigma^{2})^{-\frac{Nd_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|y_{i}-\mathscr{G}(\theta)(z_{i})\right|_{V}^{2}\right).$
+$p_{\theta}^{N}(y,z)=\prod_{i\in\left\llbracket N\right\rrbracket}p_{\theta}(y_{i},z_{i})=(2\pi\sigma^{2})^{-\frac{Nd_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|y_{i}-\mathscr{G}(\theta)(z_{i})|_{V}^{2}\right).$
 Given the data vector $D_{N}$, a shifted log-likelihood function $\ell_{N}:\tilde{\Theta}\to\mathbb{R}$ is defined by
-$\ell_{N}(\theta):=\ell_{N}(\theta|D_{N}):=\log p_{\theta}^{N}(D_{N})+\frac{Nd_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
+$\ell_{N}(\theta):=\ell_{N}(\theta|D_{N}):=\log p_{\theta}^{N}(D_{N})+\frac{Nd_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}$
 for each $\theta\in\tilde{\Theta}$. Analogously, given a single datum $D=(Y,Z)$, we define
-$\ell(\theta):=\ell(\theta|D):=\log p_{\theta}(D)+\frac{d_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\left|Y-\mathscr{G}(\theta)(Z)\right|_{V}^{2}.$
+$\ell(\theta):=\ell(\theta|D):=\log p_{\theta}(D)+\frac{d_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}|Y-\mathscr{G}(\theta)(Z)|_{V}^{2}.$
 The results obtained below hold under the frequentist assumption that the data $D_{N}$ from eq.
-2.1 is generated from a fixed ground truth $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ including the law $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$.
-   Maximum a posteriori estimator
-We are now introducing the maximum a posteriori estimator, whose statistical guarantees are presented in the following sections.
+7 is generated from a fixed ground truth $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ including the law $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$.
+2.3 Maximum a posteriori estimator
 Given the data set $D_{N}=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from eq.
-2.1, fix some $\alpha,\mathfrak{r}\in\mathbb{R}_{\geq 0}$.
+7, fix some $\alpha,\mathfrak{r}\in\mathbb{R}_{\geq 0}$.
 We define a penalized least squares functional by
-$\mathscr{J}_{\mathfrak{r},N}:\tilde{\Theta}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\text{ }\mathscr{J}_{\mathfrak{r},N}[\theta]:=-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$ (2.2)
+$\mathscr{J}_{\mathfrak{r},N}:\tilde{\Theta}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\text{ }\mathscr{J}_{\mathfrak{r},N}[\theta]:=-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2}$ (8)
 and set
-$\hat{\theta}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right\}.$ (2.3)
-We call $\hat{\theta}_{\textsc{MAP}}$ maximum a posteriori estimator, or shortly MAP estimator.
+$\hat{\theta}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{\theta\in\tilde{\Theta}}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}.$ (9)
 Remark § 2.2 (Maximum a posteriori estimator).
@@
     Up to now, we do not have formulated any specific conditions on $\tilde{\Theta}$, except of $\tilde{\Theta}\subseteq\mathbb{L}^{2}(\mathcal{O},W)$.
-In Proposition 2.5, a restriction of $\tilde{\Theta}$ is neccessary in order to provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer of the optimization problem Eq.
-2.3.
+In proposition 2.5, we need to restrict ourselves to suitable subsets $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$ (with $\mathscr{H}$ being either $H^{\alpha}(\mathcal{O},W)$ or $H^{\alpha}_{c}(\mathcal{O},W)$) in order to provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer of the optimization problem eq.
+9 over $\Theta$.
 However, the results of the statistical analysis hold true for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$.
 From the application perspective, the definition of $\hat{\theta}_{\textsc{MAP}}$ seems unfeasible as the optimization problem is formulated over a possibly infinite-dimensional space.
-The theory shown in Nickl [2023] and Nickl and Wang [2024] provide general conditions and algorithms, under which $\hat{\theta}_{\textsc{MAP}}$ can be computed in polynomial time, if $\tilde{\Theta}$ is finite dimensional.
+The theory shown in Nickl (2023) and Nickl and Wang (2024) provide general conditions and algorithms, under which $\hat{\theta}_{\textsc{MAP}}$ can be computed in polynomial time, if $\Theta$ is finite dimensional.
   2. ii)
-    The MAP estimator $\hat{\theta}_{\textsc{MAP}}$ introduced in Eq.
-2.3 has two different interpretations: On the one hand, $\hat{\theta}_{\textsc{MAP}}$ is strongly related to the classical theory of Tikhonov-regularized estimators - we refer to Engl et al.
-[1996] for a detailed analysis of the regularization of ill-posed deterministic inverse problems.
-Indeed, $\hat{\theta}_{\textsc{MAP}}$ minimizes the empirical squared distance.
-To exlude unnecessary complex solutions of the problem, the regularity is penalized in terms of the $H^{\alpha}(\mathcal{O},W)$-norm.
-On the other hand, $\hat{\theta}_{\textsc{MAP}}$ corresponds to the framework of Bayesian inverse problem techniques.
-Under Gaussian process priors on the parameter space $\tilde{\Theta}$, the corresponding likelihood of the a posteriori distribution is then maximised by $\hat{\theta}_{\textsc{MAP}}$.
-We refer here to the comments in Nickl et al.
-[2020] and Bohr and Nickl [2023], and further to Dashti et al.
-[2013], Helin and Burger [2015] for more details.
-   A convergence rate result
+    The estimator $\hat{\theta}_{\textsc{MAP}}$ introduced in eq.
+9 admits two different but related interpretations.
+On the one hand, it is closely connected to the classical theory of Tikhonov-regularized estimators, as studied in Engl et al.
+(1996), where the regularization of ill-posed deterministic inverse problems is achieved by penalizing solutions with respect to a suitable Sobolev norm.
+Specifically, $\hat{\theta}_{\textsc{MAP}}$ minimizes the empirical squared loss subject to a regularization term given by the $H^{\alpha}(\mathcal{O},W)$-norm, thereby enforcing smoothness to exclude overly complex solutions.
+On the other hand, $\hat{\theta}_{\textsc{MAP}}$ corresponds to the maximum a posteriori (MAP) estimator in a Bayesian inverse problem framework: Let $\Pi$ be a Gaussian process prior on the parameter $\theta$ with reproducing kernel Hilbert space (RKHS) $\mathscr{H}$ and corresponding prior density $\pi(\theta)\propto\exp(-\frac{\overline{\mathfrak{r}}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2})$ with $\overline{\mathfrak{r}}>0$.
+Under the Gaussian likelihood function of the model eq.
+7, any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\mathscr{H}$ with $\mathfrak{r}=\frac{1}{\sqrt{N}}\overline{\mathfrak{r}}$ has a formal interpretation as maximizer of the log-likelihood $\log\pi(\theta|D_{N})$ of the resulting posterior distributions $\Pi(\cdot|D_{N})$.
+This interpretation has been formalized, for instance, in Remark 4 of Nickl et al.
+(2020).
+We refer also to Bohr and Nickl (2023); Dashti et al.
+(2013), and Helin and Burger (2015) for additional background on MAP estimation in Gaussian Bayesian inverse problems.
+Throughout this work, we refer to $\hat{\theta}_{\textsc{MAP}}$ as the maximum a posteriori (MAP) estimator.
+2.4 A convergence rate result
 The main results of this work - proposition 2.5, theorem 2.6 and corollary 2.7 - are presented in the following.
-Firtsly, we provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a solution to the optimization problem Eq.
-2.3 over a suitable subset $\Theta\subseteq\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$.
+Firstly, we provide existence of $\hat{\theta}_{\textsc{MAP}}$ as a solution to the optimization problem eq.
+9 over a suitable subset $\Theta\subseteq\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$.
 Moreover, we are showing a concentration result controlling the prediction error and the regularity of $\hat{\theta}_{\textsc{MAP}}$ simultanously under $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability.
 Therefore, we define for every $\theta_{1}\in\tilde{\Theta}\cap H^{\alpha}(\mathcal{O},W)$, $\theta_{2}\in\tilde{\Theta}$, and $\mathfrak{r}\in\mathbb{R}_{\geq 0}$ the functional
-$\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}):=\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
+$\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{1},\theta_{2}):=\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
 We start with analytical assumption on the forward map $\mathscr{G}$.
@@
   1. C1)
-    $\mathbb{L}^{2}$-Lipschitz continuity: Let $\alpha,\gamma_{1},\kappa\in\mathbb{R}_{\geq 0}$.
-Set $\mathscr{H}=H^{\alpha}_{c}(\mathcal{O},W),\text{ if $\kappa\geq\frac{1}{2}$ }$ and $\mathscr{H}=H^{\alpha}(\mathcal{O},W),\text{ if $\kappa<\frac{1}{2}$}$.
-The map $\mathscr{G}$ is $(\alpha,\gamma_{1},\kappa)$-regular in $\mathbb{L}^{2}$-norm, if there exists a constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in\tilde{\Theta}\cap\mathscr{H}$
-    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
+[…]
+    $\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq C_{\operatorname{Lip},2}\left(1+\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\hskip-4.2679pt\lor\|\theta_{2}\|_{H^{\alpha}(\mathcal{O},W)}^{\gamma_{1}}\right)\|\theta_{1}-\theta_{2}\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
   2. C2)
     Uniform boundedness: There exists a constant $C_{\mathscr{G},\operatorname{B}}\in\mathbb{R}_{\geq 1}$, such that
-    $\sup_{\theta\in\tilde{\Theta},z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq C_{\mathscr{G},\operatorname{B}}.$
+    $\sup_{\theta\in\tilde{\Theta},z\in\mathcal{Z}}\left\{|\mathscr{G}(\theta)(z)|_{V}\right\}\leq C_{\mathscr{G},\operatorname{B}}.$
   3. C3)
     $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha,\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, such that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$.
-The map $\mathscr{G}$ is $(\alpha,\gamma_{2},\kappa)$-regular in $\mathbb{L}^{\infty}$-norm, if there exists a constant $C_{\operatorname{Lip},\infty}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O},W)$
-    $\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
-Remark § 2.4 ($\mathbb{L}^{\infty}$-Lipschitz continuity).
-The effect that stronger assumptions on the statistical model are needed to prove consistency for the maximum likelihood estimator than to show corresponding concentration results in the Bayesian setting (see Le Cam [1986]) is also apparent in the nonlinear random regression setting considered here: While in Nickl [2023] it suffices to assume condition C1) and C2) (as special case with $\gamma_{1}=0$) to obtain posterior convergence results, in Bohr and Nickl [2023] consistency of the MAP estimator necessitates a $\mathbb{L}^{\infty}$-stability estimate of the forward map $\mathscr{G}$, i.e.,
-$\left\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\right\|_{\infty}\lesssim\left\|\theta-\theta^{\prime}\right\|_{\infty},$
-uniformly over the whole parameter space.
-C3) in 2.3 generalizes this condition by replacing the uniform norm by the stronger $\left\|\cdot\right\|_{C^{\eta_{2}}}$-norm.
-Moreover, we now allow a polynomial increase of order $\gamma_{2}$.
-While the conditions formulated in Bohr and Nickl [2023] (which correspond to the special case $\eta_{1}=\eta_{2}=\gamma_{2}=0$ in the present setting!) are insufficient for the Darcy problem, we see in Section 3 and Appendix A that the associated solution map fulfills C3) with $\eta_{1}=\eta_{2}=1$ and $\gamma_{2}=4$.
-Note that the condition $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$ in C3) ensures that all norms are well-defined as in this situation the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{i}}(\mathcal{O},W)$, $i\in\left\llbracket 2\right\rrbracket$, holds true.
-The first main results addresses the open question of existence of $\hat{\theta}_{\textsc{MAP}}$ as the maximizer of the penalized Tikhonov-regularized functional Eq.
-2.2.
+The map $\mathscr{G}$ is $(\alpha,\gamma_{2},\eta_{1},\eta_{2})$-regular in $\mathbb{L}^{\infty}$-norm, if there exists a constant $C_{\operatorname{Lip},\infty}\in\mathbb{R}_{>0}$, such that for all $\theta_{1},\theta_{2}\in\tilde{\Theta}\cap\mathscr{H}$
+    $\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\infty}\leq C_{\operatorname{Lip},\infty}\left(1+\|\theta_{1}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\|\theta_{2}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\|\theta_{1}-\theta_{2}\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
+Remark § 2.4 ($\mathbb{L}^{2}$- and $\mathbb{L}^{\infty}$-Lipschitz continuity).
+The assumptions C1)–C3) imposed on the forward map $\mathscr{G}$ are abstract but serve as minimal requirements for establishing statistical guarantees for the MAP estimator $\hat{\theta}_{\textsc{MAP}}$.
+Condition C1) quantifies the degree of ill-posedness through the smoothing parameter $\kappa$, which also directly influences the convergence rates obtained in theorem 2.6.
+In contrast, C3) is of a more technical nature: it ensures that entropy numbers of images $\{\mathscr{G}(\theta):\theta\in H^{\alpha}(\mathcal{O},W,r)\}$, $r\in\mathbb{R}_{>0}$, of Sobolev balls under $\mathscr{G}$ are sufficiently controlled and is particularly tailored to the PDE-constrained setting considered here, such as the Darcy flow.
+For this purpose, we weaken the $\mathbb{L}^{\infty}$-stability condition
+$\|\mathscr{G}(\theta)-\mathscr{G}(\theta^{\prime})\|_{\infty}\lesssim\|\theta-\theta^{\prime}\|_{\infty},$ (10)
+as used in Bohr and Nickl (2023), by replacing the uniform norm on the right with the stronger $C^{\eta_{2}}$-norm and additionally allowing polynomial growth of order $\gamma_{2}$.
+The parameters $\eta_{1}$ and $\eta_{2}$ do not appear in the convergence rates obtained in theorem 2.6 explicitly, but the condition $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$ ensures that the relevant embeddings $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{i}}(\mathcal{O},W)$, $i\in\{1,2\}$, hold, so that all norms are well-defined.
+It is also not surprising that stronger assumptions are needed to establish consistency of the MAP estimator than to prove posterior contraction in the Bayesian setting, such as in Giordano and Nickl (2020) for the Darcy problem.
+This phenomenon is well known in classical statistics (see, e.g., Le Cam (1986)) and becomes evident in the general nonlinear random regression framework as well: while Nickl (2023) derives posterior contraction under the comparatively mild conditions C1) and C2) (with $\gamma_{1}=0$), the analysis in Bohr and Nickl (2023) requires the uniform stability condition eq.
+10 to prove consistency of the MAP estimator.
+In our setting, the condition eq.
+10 correspond to the special case $\eta_{1}=\eta_{2}=\gamma_{2}=0$, which is insufficient for the Darcy problem.
+As we demonstrate in section 3 and section 4, the associated solution map fulfills C3) with $\eta_{1}=\eta_{2}=1$ and $\gamma_{2}=4$.
+The first main results address the open question of existence of $\hat{\theta}_{\textsc{MAP}}$ as the maximizer of the penalized Tikhonov-regularized functional eq.
+8.
 For the white noise model, the authors in Nickl et al.
-[2020] provide a proof using the white noise formulation of the regression model.
-We adapt the corresponding proof to the random design regression model, also exploiting the Direct Method of the Calculus of Variations, making reference to Dacorogna [2008].
-The proof can be found in Appendix A.
+(2020) provide a proof using the white noise formulation of the regression model.
+We adapt the corresponding proof to the random design regression model, also exploiting the Direct Method of the Calculus of Variations, making reference to Dacorogna (2008).
+The proof can be found in section 4.
 Proposition § 2.5 (Existence of $\hat{\theta}_{\textsc{MAP}}$).
@@
 for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, some reference parameter $\tilde{\theta}\in\Theta$ (with $\Theta\subseteq\tilde{\Theta}\cap\mathscr{H}$) and some $c_{1}\in\mathbb{R}_{\geq 1}$.
-To shorten notation, we define the family of symbols $\texttt{P}:=\left\{\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma\right\}$ and write $c(\texttt{P})$ for a constant to emphasize the dependency.
+To shorten notation, we define the family of symbols
+$\texttt{P}:=\left\{\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\sigma\right\}$
+and write $c(\texttt{P})$ for a constant to emphasize the dependency.
 Theorem § 2.6 (Concentration inequality).
@@
 Let $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ from eq.
-2.1 with some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$.
+7 with some fixed $\theta_{\mathfrak{o}}\in\tilde{\Theta}$.
 Then the following hold.
@@
     any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ and $\tilde{\theta}\in\Theta$ satisfies
-    $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{2}\cdot\exp(-\bar{c}\cdot NR^{2})$ (2.4)
+    $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{2}\cdot\exp(-\bar{c}\cdot NR^{2})$ (11)
     with $c_{2}=c_{2}(d_{V})\in\mathbb{R}_{>0}$.
@@
     we have for any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$
-    $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}\left[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\right]\leq c_{3}\cdot\inf_{\tilde{\theta}\in\Theta}\left\{\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}(1+\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O})}^{2})+\frac{1}{N}\right\}.$ (2.5)
+    $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}\left[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\right]\leq c_{3}\cdot\inf_{\tilde{\theta}\in\Theta}\left\{\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}^{2}(1+\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O})}^{2})+\frac{1}{N}\right\}.$
 Corollary § 2.7 (Consistency).
-Under the assumptions of Theorem 2.6, assume $\theta_{\mathfrak{o}}\in\Theta$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}\in\mathbb{R}_{>0}$.
+Under the assumptions of theorem 2.6, let $\theta_{\mathfrak{o}}\in\Theta$ with $\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}\in\mathbb{R}_{>0}$.
 Let $\mathfrak{r}\in\mathbb{R}_{>0}$ be given by
@@
 Then, for all $\bar{c}\in\mathbb{R}_{\geq 1}$, there exists $c_{4}=c_{4}(\bar{c},\operatorname{L},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{4}\mathfrak{r}^{2}_{N}\right\}\right)\leq c_{5}\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{4}\mathfrak{r}^{2}_{N}\right)\leq c_{5}\exp(-\bar{c}\cdot N\mathfrak{r}_{N}^{2})$
 with $c_{5}=c_{5}(d_{V})\in\mathbb{R}_{>0}$. In particular, there is $c_{6}=c_{6}(d_{V},\operatorname{L})\in\mathbb{R}_{>0}$, such that
-$\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right]\right\}\leq\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta)]\right\}\leq c_{6}\cdot\mathfrak{r}_{N}^{2}.$
+$\sup_{\theta\in\Theta\cap H^{\alpha}(\mathcal{O},W,\operatorname{L})}\left\{\mathbb{E}^{N}_{\theta}\left[\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta)\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right]\right\}\leq c_{6}\cdot N^{-\frac{2(\alpha+\kappa)}{2(\alpha+\kappa)+d}}.$
 Remark § 2.8 (Consistency and prediction error).
   1. i)
-    Note, that Theorem 2.6 does not require $\theta_{\mathfrak{o}}\in\Theta$ as long as it can be approximated sufficiently by some $\tilde{\theta}\in\Theta$.
-Therefore, the assumption $\theta_{\mathfrak{o}}\in\Theta$ in Corollary 2.7 can also be weakened to the effect that $\mathscr{G}(\tilde{\theta})$ approximates $\mathscr{G}(\theta_{\mathfrak{o}})$ in some sense (see equation 2.23 in Bohr and Nickl [2023] for instance).
+[…]
   2. ii)
     Evidently, corollary 2.7 provides the consistency of the MAP estimator $\hat{\theta}_{\textsc{MAP}}$.
-The rate of convergence for the prediction loss is known to be optimal for $\kappa$-smoothing linear forward maps $\mathscr{G}$.
+The rate of convergence for the prediction loss fits the interpretation of classical inverse problems: The smoothness parameter $\kappa$ of the forward map $\mathscr{G}$, quantifying the ill-posedness of the problem, influences directly the rate of convergence.
+It is known that this rate of convergence is minimax-optimal for $\kappa$-smoothing linear forward maps $\mathscr{G}$ in various settings, see, e.g., Cavalier (2008).
+In contrast, $\eta_{1}$ and $\eta_{2}$ from condition C3) do not appear explicitely in the rate, but ensure by the condition $\alpha>\eta_{1}\lor\eta_{2}+d\lor\frac{d}{2}(1+\gamma_{2})$ that the model is sufficiently regular.
 Without any further assumptions, there is no hope to obtain an analogous upper bound for the estimation risk as well.
-However, the next corollary demonstrates stability estimates under a inverse continuity modulus of $\mathscr{G}^{-1}$.
+However, the next corollary demonstrates stability estimates under an inverse continuity modulus of $\mathscr{G}^{-1}$.
 Condition § 2.9 (Forward Map II).
   1. C7)
-    Inverse continuity modulus: Additionally to 2.3, define for any $\delta\in\mathbb{R}_{>0}$ the set
-    $\Lambda_{\delta}:=\left\{(\theta_{1},\theta_{2})\in\tilde{\Theta}^{2}:\,\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{M},\,\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq\delta\right\}$
+    Inverse continuity modulus: Additionally to 2.3, define for any $\delta\in\mathbb{R}_{>0}$
+    $\Lambda_{\delta}:=\left\{(\theta_{1},\theta_{2})\in\tilde{\Theta}^{2}:\,\hskip-2.84526pt\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}+\|\theta_{2}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{M},\,\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq\delta\right\}$
     for some universal constant $\operatorname{M}\in\mathbb{R}_{>0}$. There exists a constant $C_{\operatorname{M}}\in\mathbb{R}_{>0}$ and some $0<\tau\leq 1$, such that for $\delta$ small enough
-    $\sup\left\{\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}:\,(\theta,\theta^{\prime})\in\Lambda_{\delta}\right\}\leq C_{\operatorname{M}}\cdot\delta^{\tau}.$
+    $\sup\left\{\|\theta_{1}-\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O},W)}:\,(\theta,\theta^{\prime})\in\Lambda_{\delta}\right\}\leq C_{\operatorname{M}}\cdot\delta^{\tau}.$
 Corollary § 2.10 (Stability).
-Under 2.3 and 2.9 and under the conditions formulated in Theorem 2.6, fory every $\bar{c}\in\mathbb{R}_{>0}$, there exists $c_{7}=c_{7}(C_{\operatorname{M}},\bar{c},\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c_{7}\cdot\mathfrak{r}_{N}^{\tau},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{7}\right\}\right)\geq 1-c_{8}\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
+Under 2.3 and 2.9 and under the conditions formulated in theorem 2.6, fory every $\bar{c}\in\mathbb{R}_{>0}$, there exists $c_{7}(C_{\operatorname{M}},\bar{c},\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c_{7}\cdot\mathfrak{r}_{N}^{\tau},\,\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{7}\right)\geq 1-c_{8}\cdot\exp(-\bar{c}\cdot N\mathfrak{r}_{N}^{2})$
 for some $c_{8}=c_{8}(d_{V})\in\mathbb{R}_{>0}$.
-3   3Application: PDE constrained regression models
+3 Application: PDE constrained regression models
 We apply the results from section 2 on the inverse problems described by the PDE eq.
-1.2 and derive corresponding convergence rates for the parameter $f$.
+2 and derive corresponding convergence rates for the parameter $f$.
 We further discuss the applicability of the setup for other PDE models.
@@
 For any $d\in\mathbb{N}$, integer $\alpha>\frac{d}{2}$, a fixed constant $f_{\min}\in\mathbb{R}_{>0}$, and denoting by $\partial_{n}$ the outward pointing normal derivative, define the parameter space
-$\mathcal{F}:=\mathcal{F}_{\alpha,f_{\min}}:=\left\{f\in H^{\alpha}(\mathcal{O}):\,f>f_{\min}\text{ on $\mathcal{O}$},f|_{\partial\mathcal{O}}\equiv 1,\partial^{j}_{n}f\equiv 0,j\in\left\llbracket\alpha-1\right\rrbracket\right\},$ (3.1)
+$\mathcal{F}:=\mathcal{F}_{\alpha,f_{\min}}:=\left\{f\in H^{\alpha}(\mathcal{O}):\,f>f_{\min}\text{ on $\mathcal{O}$},f|_{\partial\mathcal{O}}\equiv 1,\partial^{j}_{n}f\equiv 0,j\in\left\llbracket\alpha-1\right\rrbracket\right\},$ (12)
 and its subclass $\mathcal{F}(\operatorname{L}):=\mathcal{F}\cap H^{\alpha}(\mathcal{O},\operatorname{L})$, $\operatorname{L}\in\mathbb{R}_{>0}$.
 Let $g\in C^{\infty}(\mathcal{O})$ be a known heat source function.
-If $f\in\mathcal{F}_{\alpha,f_{\min}}$ for some $\alpha>\frac{d}{2}+1$ and $f_{\min}\in\mathbb{R}_{>0}$, then well-known PDE theory (see Gilbarg and Trudinger [2001] for instance) provides the existence and uniquess of a classical solution $u\in C^{0}(\overline{\mathcal{O}})\cap C^{2+\nu}(\mathcal{O})$ of the Dirichlet boundary value problem
-$\begin{dcases}\operatorname{div}(f(x)\nabla u(x))&=g(x)\text{ on $\mathcal{O}$,}\\ \hfill u(x)&=0\text{ on $\partial\mathcal{O}$,}\end{dcases}$ (3.2)
+If $f\in\mathcal{F}_{\alpha,f_{\min}}$ for some $\alpha>1+\frac{d}{2}$ and $f_{\min}\in\mathbb{R}_{>0}$, then well-known PDE theory (see Gilbarg and Trudinger (2001) for instance) provides the existence and uniquess of a classical solution $u\in C^{0}(\overline{\mathcal{O}})\cap C^{2+\nu}(\mathcal{O})$ of the Dirichlet boundary value problem
+$\begin{dcases}\hfill\operatorname{div}(f(x)\cdot\nabla u(x))&=g,\text{ on $\mathcal{O}$},\\ \hfill u(x)\hskip 2.84526pt&=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (13)
 for some (small) $\nu\in\mathbb{R}_{>0}$.
-In other words, the solution operator $\mathcal{F}\ni f\mapsto G^{D}(f)=:u_{f}\in\mathbb{L}^{2}(\mathcal{O})$ is always well-defined.
-The nonlinearity of $G^{D}$ becomes more clear when considering the Feynmann-Kac formula, which reads for each $x\in\mathcal{O}$ as
-$u_{f}(x)=-\frac{1}{2}\mathbb{E}^{x}\left[\int_{0}^{\tau_{\mathcal{O}}}g(X_{s}^{f})\mathrm{d}s\right],$
-where $(X_{t}^{f})_{t\in\mathbb{R}_{\geq 0}}$ is a Markov process starting at $x$ with associated exit time $\tau_{\mathcal{O}}$.
-We refer to Bass [1998] for an extensive overview and more details.
+In other words, we can define a non-linear solution map $G^{D}:\mathcal{F}\to\mathbb{L}^{2}(\mathcal{O})$, mapping any $f\in\mathcal{F}$ to the corresponding solution $u_{f}=G^{D}(f)\in\mathbb{L}^{2}(\mathcal{O})$ of eq.
+13.
 Following eq.
-2.1, we then observe
-$Y_{i}=G^{D}(f_{\mathfrak{o}})(X_{i})+\sigma\varepsilon_{i},i\in\left\llbracket N\right\rrbracket,$ (3.3)
+7, we then observe
+$Y_{i}=G^{D}(f_{\mathfrak{o}})(X_{i})+\sigma\varepsilon_{i},i\in\left\llbracket N\right\rrbracket,$ (14)
 where $X_{1},\dots,X_{N}$ are drawn from the uniform law on the domain $\mathcal{O}$, where we assume for simplicity that $\operatorname{vol}(\mathcal{O})=1$.
@@
     A link function $\Psi$ is called regular, if and only if all their derivatives are bounded, i.e.
-    $\sup_{x\in\mathbb{R}}\left\{\left|\Psi^{(k)}(x)\right|\right\}<\infty,$
+    $\sup_{x\in\mathbb{R}}\left\{|\Psi^{(k)}(x)|\right\}<\infty,$
     for all $k\in\mathbb{N}$.
-For examples and typical choices of (regular) link functions in practise, we refer to Nickl et al.
-[2020] and Nickl and Wang [2024].
-In the notation of Section 2, throughout this section we set $\mathcal{Z}=\mathcal{O}$, $\mathscr{Z}:=\mathscr{B}_{\mathcal{O}}$ and $\zeta=\mathcal{L}_{\mathcal{O}}$ (recalling that $\operatorname{vol}(\mathcal{O})=1$!).
+For examples and typical choices of (regular) link functions in practice, we refer to Nickl et al.
+(2020) and Nickl and Wang (2024).
+In the notation of section 2, throughout this section we set $\mathcal{Z}=\mathcal{O}$, $\mathscr{Z}:=\mathscr{B}_{\mathcal{O}}$ and $\zeta=\mathcal{L}_{\mathcal{O}}$ (recalling that $\operatorname{vol}(\mathcal{O})=1$).
 We set $\tilde{\Theta}:=\Theta:=\left\{\Psi^{-1}\circ f:\,f\in\mathcal{F}\right\}=H^{\alpha}_{c}(\mathcal{O})$ (the pulled-back) parameter space, and
@@
 Given the data set $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$, we define a Tikhonov-regularized functional
-$J_{\mathfrak{r},N}:\mathcal{F}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\,f\mapsto-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-G^{D}(f)(X_{i})|^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\Psi^{-1}\circ f\right\|_{H^{\alpha}(\mathcal{O})}^{2}$ (3.4)
+$J_{\mathfrak{r},N}:\mathcal{F}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\,f\mapsto-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-G^{D}(f)(X_{i})|^{2}-\frac{\mathfrak{r}^{2}}{2}\|\Psi^{-1}\circ f\|_{H^{\alpha}(\mathcal{O})}^{2}$ (15)
 for any $\mathfrak{r}\in\mathbb{R}_{>0}$.
 In the notation of eq.
-2.2, we have $J_{\mathfrak{r},N}[f]=\mathscr{J}_{\mathfrak{r},N}[\theta]$ for any $f=f_{\theta}\in\mathcal{F}$, such that maximizing $J_{\mathfrak{r},N}$ over $\mathcal{F}$ is equivalent to maximizing $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta=H^{\alpha}_{c}(\mathcal{O})$.
-Any pairs of maximizer will be denoted by
+8, we have $J_{\mathfrak{r},N}[f]=\mathscr{J}_{\mathfrak{r},N}[\theta]$ for any $f=f_{\theta}\in\mathcal{F}$, such that maximizing $J_{\mathfrak{r},N}$ over $\mathcal{F}$ is equivalent to maximizing $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta=H^{\alpha}_{c}(\mathcal{O})$.
+Any pairs of maximizers will be denoted by
 $\hat{f}_{\textsc{MAP}}:\in\operatorname*{arg\,max}_{f\in\mathcal{F}}\left\{J_{\mathfrak{r},N}[f]\right\}\text{ and }\hat{\theta}_{\textsc{MAP}}=\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\in\operatorname*{arg\,max}_{\theta\in H^{\alpha}_{c}(\mathcal{O})}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}.$
 For a link function $\Psi$, define for $f_{1},f_{2}\in\mathcal{F}$
-$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{1},f_{2}):=\left\|G(f_{1})-G(f_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}^{2}+\mathfrak{r}^{2}\left\|\Psi^{-1}\circ f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2}.$
+$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{1},f_{2}):=\|G(f_{1})-G(f_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}^{2}+\mathfrak{r}^{2}\|\Psi^{-1}\circ f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2}.$
 Identifying $f_{\theta_{1}}$ and $f_{\theta_{2}}$ with $\theta_{1},\theta_{2}\in H^{\alpha}_{c}(\mathcal{O})$, respectively, we see
@@
 Let $\mathcal{F}$ be as in eq.
-3.1 for some $\alpha>1+\frac{5}{2}d$ and $f_{\min}\in\mathbb{R}_{>0}$.
+12 for some $\alpha>1+\frac{5}{2}d$ and $f_{\min}\in\mathbb{R}_{>0}$.
 Let $G^{D}(f)=u_{f}$ denote the unique solution of eq.
-3.2 and let $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$ from Eq.
-3.3 for some $f_{\mathfrak{o}}\in\mathcal{F}$.
+13 and let $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$ from eq.
+14 for some $f_{\mathfrak{o}}\in\mathcal{F}$.
 Let $\Psi$ be a regular link function.
 Let $J_{\mathfrak{r}_{N},N}$ be as in eq.
-3.4 with
+15 with
 $\mathfrak{r}_{N}:=N^{-\frac{\alpha+1}{2(\alpha+1)+d}}.$
@@
 Then there exist constants $c_{1}=c_{1}(d,\sigma,\mathcal{O},f_{\min},g,\Psi,\alpha,\operatorname{L}),c_{2},c_{3}\in\mathbb{R}_{>0}$, such that for any maximizer $\hat{f}_{\textsc{MAP}}$, we have for all $\operatorname{M}\geq c_{1}$
-    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{P}_{f}^{N}\left(\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f)\geq\operatorname{M}\mathfrak{r}_{N}^{2}\right\}\right)\right\}\leq c_{2}\exp\left(-c_{3}\operatorname{M}^{2}N\mathfrak{r}_{N}^{2}\right).$ (3.5)
+    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{P}_{f}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f)\geq\operatorname{M}\cdot\mathfrak{r}_{N}^{2}\right)\right\}\leq c_{2}\exp\left(-c_{3}\cdot\operatorname{M}^{2}N\mathfrak{r}_{N}^{2}\right).$ (16)
   3. iii)
     Assume additionally that $g\geq g_{\min}\in\mathbb{R}_{>0}$.
-For every $\bar{c}\in\mathbb{R}_{>0}$ exists $c_{2}=c_{2}(\alpha,d,\mathcal{O},f_{\min},g_{\min},f_{\mathfrak{o}},\sigma)$ large enough and $c_{3}\in\mathbb{R}_{>0}$, such that
-    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\left\{\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{2}\mathfrak{r}_{N}^{\frac{\alpha-1}{\alpha+1}},\,\left\|\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{2}\right\}\right)\geq 1-c_{3}\exp(-c_{1}N\mathfrak{r}_{N}^{2}).$
+For every $\bar{c}\in\mathbb{R}_{>0}$, there exists a constant $c_{2}=c_{2}(\alpha,d,\mathcal{O},f_{\min},g_{\min},f_{\mathfrak{o}},\sigma)$ large enough and $c_{3}\in\mathbb{R}_{>0}$, such that
+    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{2}\cdot\mathfrak{r}_{N}^{\frac{\alpha-1}{\alpha+1}},\,\|\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O})}\leq c_{2}\right)\\ \geq 1-c_{3}\cdot\exp(-c_{1}\cdot N\mathfrak{r}_{N}^{2}).$
 Theorem § 3.3 (Prediction and estimation risk).
-Grant the conditions formulated in Theorem 3.2.
+Under the conditions formulated in theorem 3.2, we have the following results.
   1. i)
-    (Prediction Error) For every $\operatorname{L}\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$, there exists a finite constant $c_{4}=c_{4}(\alpha,\beta,d,\mathcal{O},f_{\min},L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
-    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f)\right\|_{H^{\beta}(\mathcal{O})}\right]\right\}\leq c_{4}\cdot N^{-\frac{\alpha+1-\beta}{2(\alpha+1)+d}}.$
+    (Prediction Error) For every $\operatorname{L}\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$, there exists a finite constant $c_{4}=c_{4}(\alpha,\beta,d,\mathcal{O},f_{\min},L,\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
+    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f)\|_{H^{\beta}(\mathcal{O})}^{2}\right]\right\}\leq c_{4}\cdot N^{-\frac{2(\alpha+1-\beta)}{2(\alpha+1)+d}}.$
   2. ii)
     (Estimation Error) Let $g\geq g_{\min}\in\mathbb{R}_{>0}$.
-For every $\operatorname{L}\in\mathbb{R}_{>0}$, there exists $c_{5}=c_{5}(g_{\min},f_{\min},\mathcal{O},d,\alpha,L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
-    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\right\}\leq c_{5}\cdot N^{-\frac{\alpha-1}{2(\alpha+1)+d}}.$
-Remark § 3.4 (Rate-optimality).
-The convergence rates for the prediction risk presented in Theorem 3.3 correspond to the usual interpretation of statistical inverse problems: the smoothness of the parameter of interest $f$ is represented by the parameter $\alpha$.
-The ill-posedness of the underlying problem by $\kappa=1$, which corresponds to the smoothing property of $G^{D}$ (Indeed, Lemma 23 in Nickl et al.
-[2020] shows that if $f$ is $\alpha$-smooth, then $G^{D}(f)$ is $\alpha+1$-smooth).
-The prediction risk is considered with respect to the $H^{\beta}$-norm (which can be interpreted as differentiating $\beta$ times), which yields the characteristic numerator $\alpha+1-\beta$ of the rate.
-The rates of convergence here coincide with the rates obtained in Nickl et al.
-[2020] for the Darcy problem.
-There it is also shown that the rate of convergence of the prediction risk is minimax-optimal for the choice $\beta=2$, which also holds true in the random design model, as the proof can be easily adopted.
-The proofs of Theorem 3.2 and Theorem 3.3 are given in detail in Appendix A.
+For every $\operatorname{L}\in\mathbb{R}_{>0}$, there exists a constant $c_{5}=c_{5}(g_{\min},f_{\min},\mathcal{O},d,\alpha,L,\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
+    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\|\hat{f}_{\textsc{MAP}}-f\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\right]\right\}\leq c_{5}\cdot N^{-\frac{2(\alpha-1)}{2(\alpha+1)+d}}.$
+Remark § 3.4 (Prediction and estimation risk).
+  1. i)
+    The convergence rates for the prediction risk presented in theorem 3.3 i) align with the classical interpretation of statistical inverse problems.
+The smoothness of the parameter of interest $f$ is quantified by $\alpha$, while the ill-posedness of the problem is captured by $\kappa=1$, reflecting the smoothing property of the forward map $G^{D}$.
+Indeed, Lemma 23 in Nickl et al.
+(2020) shows that if $f$ is $\alpha$-smooth, then $G^{D}(f)$ is $\alpha+1$-smooth.
+Considering the prediction risk in the $H^{\beta}$-norm, which corresponds to differentiation of order $\beta$, yields the characteristic numerator $\alpha+1-\beta$ in the convergence rate.
+These rates coincide exactly with those obtained in the white noise model setting in Nickl et al.
+(2020) for the Darcy problem.
+Theorem 10 in Nickl et al.
+(2020) in addition shows that for the $H^{2}$-prediction loss the obtained rate of convergence is minimax-optimal.
+This optimality also carries over to the random design model considered here, as the proof techniques can be adapted without difficulty.
+Hence, the results in Nickl et al.
+(2020) provides a rigorous benchmark indicating that the convergence rates derived in these settings are indeed statistically sharp.
+  2. ii)
+    The additional assumption $g\geq g_{\min}>0$ in theorem 3.2 iii) and theorem 3.3 ii) plays a crucial role in the analysis.
+It ensures the validity of a stability estimate for the inverse operator $(G^{D})^{-1}$, which is central to deriving the convergence rate for the estimation error.
+Such a condition is also imposed in Lemma 24 of Nickl et al.
+(2020), which serves as a foundation for the current proof strategy.
+We additionally refer to Richter (1981) and Bonito et al.
+(2017b) for a broader discussion on injectivity and global stability estimates for inverse problems of this type.
+  3. iii)
+    In contrast to the prediction error, the rates of convergence for the estimation error derived in theorem 3.3 ii) do not match those obtained in Giordano and Nickl (2020) for the posterior contraction, which are generally slower.
+This discrepancy is partly due to the lower posterior regularity compared to the MAP estimator, resulting in suboptimal rates for the Bayesian approach.
+However, the optimality of the estimation error rates for $\hat{f}_{\textsc{MAP}}$ established here remains an open question.
+Classical nonparametric theory would suggest a faster rate, such as having exponent $\frac{2\alpha}{2(\alpha+1)+d}$.
+It is unclear to the author whether this gap is a consequence of the estimation methodology and the proof techniques employed—particularly those relying on inverse stability estimates and associated PDE analytic arguments—or whether it reflects an intrinsic structural limitation of the Darcy problem itself.
+Further investigation is needed to clarify this issue.
+The proofs of theorem 3.2 and theorem 3.3 are given in detail in section 4.
 They are based on verifying that the forward map $\mathscr{G}^{D}$ satisfies C1)-C4) formulated in 2.3 and 2.9 for suitable parameters $\alpha,\kappa,\gamma_{1},\gamma_{2},\eta_{1}$ and $\eta_{2}$, with $V=W=\mathbb{R}$, and an application of the results in section 2.
@@
 The general framework presented in section 2 does not only apply to the Darcy problem, it promises to consider various PDE-related problems.
-As already brielfly discussed in Remark 2.4, the results in Bohr and Nickl [2023] concerning the MAP estimate are special cases by setting $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
-Hence, the solution map of the time-independent Schrödinger equation satisfies also meets the requirements of 2.3 and 2.9 taking the results of Nickl et al.
-[2020] also into account.
-Moreover, very analogous and elementary computations show that the solution map of the heat equation studied in Kekkonen [2022] also satisfies the conditions in 2.3 and 2.9 with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
-In very recent work, Nickl and Titi [2024] are considering a non-linear Bayesian data assimilation model for the periodic $d=2$-dimensional Navier Stokes equation with unknown initial condition.
+As already brielfly discussed in remark 2.4, the results in Bohr and Nickl (2023) concerning the MAP estimate are special cases by setting $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
+Hence, the solution map of the time-independent Schrödinger equation satisfies the requirements of 2.3 and 2.9 taking the results of Nickl et al.
+(2020) also into account.
+Moreover, very analogous and elementary computations show that the solution map of the heat equation studied in Kekkonen (2022) also satisfies the conditions in 2.3 and 2.9 with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
+In very recent work, Nickl and Titi (2024) are considering a non-linear Bayesian data assimilation model for the periodic $d=2$-dimensional Navier Stokes equation with unknown initial condition.
 In this setting $u=u_{\theta}:\mathcal{O}\times(0,T)\to\mathbb{R}^{2}$ solves the non-linear evolution equation
@@
 For this type of equation in two dimensions a complete solution theory already exists.
-In particular, if the initial condition $\theta$ is contained in some specific function class $\tilde{\mathcal{F}}$, the solution operator ~$\tilde{\mathcal{F}}\ni\theta\mapsto G^{NS}(\theta)=u_{f}\in\mathbb{L}^{2}(\mathcal{O}\times(0,T))$ is always well-defined.
+In particular, if the initial condition $\theta$ is contained in some specific function class $\tilde{\mathcal{F}}$, the solution operator $\tilde{\mathcal{F}}\ni\theta\mapsto G^{NS}(\theta)=u_{f}\in\mathbb{L}^{2}(\mathcal{O}\times(0,T))$ is always well-defined.
 Based on a data set $D_{N}\in(\mathbb{R}^{2}\times\mathcal{O}\times(0,T))^{N}$ arising from eq.
-2.1, it seems feasible to apply the theory from Section 2.
-Indeed, using the theory presented in Nickl and Titi [2024], elementary computations show that the solution map is $(\alpha,0,0)$-regular in $\mathbb{L}^{2}$-norm and uniformly bounded, thus satisfying C1) and C2) from 2.3.
+7, it seems feasible to apply the theory from section 2.
+Indeed, using the theory presented in Nickl and Titi (2024), elementary computations show that the solution map is $(\alpha,0,0)$-regular in $\mathbb{L}^{2}$-norm and uniformly bounded, thus satisfying C1) and C2) from 2.3.
 C7) from 2.9 can also be shown with $\tau=1$.
-If C3) of 2.3 is additionally satisfied, one would obtain rates of the form $N^{-\frac{\alpha}{\alpha+2}}$ for estimating $\theta$ both in prediction and estimation error.
-Appendix A Appendix
-   Proofs of Section 2
+If C3) of 2.3 is additionally satisfied, one would obtain rates of the form $N^{-\frac{\alpha}{\alpha+2}}$ for estimating $\theta$ - both in prediction and estimation error.
+4 Appendix
+4.1 Proofs of section 2
 In this section, we provide detailed proofs of the main results presented in section 2.
@@
 The proof follows the approach of Nickl et al.
-[2020], where the existence of $\hat{\theta}_{\textsc{MAP}}$ is shown for the white noise formulation of the model.
-The authors make use of the direct method of the calculus of variations (see Dacorogna [2008] for instance) and split the proof into two parts: in one part it is shown that $\hat{\theta}_{\textsc{MAP}}$ exists as a maximizer of the functional $\mathscr{J}_{\mathfrak{r},N}$ over bounded subsets of $\Theta$ (local existence).
-In the other part it is shown that this already suffices to conclude the existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer over the whole parameter space $\Theta$ (global existence).
+(2020), where the existence of $\hat{\theta}_{\textsc{MAP}}$ is shown for the White Noise formulation of the model.
+The authors make use of the direct method of the calculus of variations (see Dacorogna (2008) for instance) and split the proof into two parts: Firstly, it is shown that $\hat{\theta}_{\textsc{MAP}}$ exists as a maximizer of the functional $\mathscr{J}_{\mathfrak{r},N}$ over bounded subsets of $\Theta$ (local existence).
+Secondly, it is shown that this already suffices to conclude the existence of $\hat{\theta}_{\textsc{MAP}}$ as a maximizer over the whole parameter space $\Theta$ (global existence).
 In the present work, we adapt the proof idea to the random design regression setting.
@@
 Let $\mathcal{T}_{w}$ denote the weak topology of $\tilde{\mathscr{H}}$, i.e., the coarsest topology with respect to which all bounded linear functionals $L:\tilde{\mathscr{H}}\to\mathbb{R}$ are continuous.
-With a slight abuse of notation, the subspace topology on subsets of $\tilde{\mathscr{H}}$ is also denoted by $\mathcal{T}_{w}$.
-By standard theory of functional analysis (see, for instance, Megginson [1998], Theorem 2.6.23), on any closed ball ~$\tilde{\mathscr{H}}(r):=\{\theta\in\tilde{\mathscr{H}}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O})}\leq r\}$ of radius $r\in\mathbb{R}_{>0}$, the weak topology $\mathcal{T}_{w}$ is metrizable by some metric $\tilde{d}$.
-Defining analogously :=×$\mathscr{H}:=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\tilde{\mathscr{H}}$, the weak topology $\mathcal{T}_{w}^{W}$ on $\mathscr{H}$ equals the product ×$\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\mathcal{T}_{w}$ (see proof of Theorem 3.10 in Brezis [2011]).
-By a standard fact of functional analysis, on each closed ball $\mathscr{H}(r):=\{\theta\in\mathscr{H}:\,\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq r\}$, the weak topology $\mathcal{T}_{w}^{W}$ is then also metrizable by some metric $d$.
-In the following we claim that $-\mathscr{J}_{\mathfrak{r},N}:\Theta\cap\mathscr{H}(2^{j})\to\mathbb{R}\cup\{\infty\}$ is lower semicontinous with respect to $\mathcal{T}_{w}^{W}$ for any $j\in\mathbb{N}$.
-To prove this, we need the following general result from functional analysis.
-Lemma § A.1 (Compact operators and weak convergence).
-Let $(X,\left\|\cdot\right\|_{X})$ and $(Y,\left\|\cdot\right\|_{Y})$ be two Banach spaces.
-Let $T:X\to Y$ be a linear compact operator.
+With a slight abuse of notation, the corresponding subspace topology on subsets of $\tilde{\mathscr{H}}$ is also denoted by $\mathcal{T}_{w}$.
+By standard theory of functional analysis (see for instance Megginson (1998), Theorem 2.6.23), on any closed ball $\tilde{\mathscr{H}}(r):=\{\theta\in\tilde{\mathscr{H}}:\,\|\theta\|_{H^{\alpha}(\mathcal{O})}\leq r\}$ of radius $r\in\mathbb{R}_{>0}$, the weak topology $\mathcal{T}_{w}$ is metrizable by some metric $\tilde{d}$.
+Defining the product space $\mathscr{H}:=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\tilde{\mathscr{H}}$, the weak topology $\mathcal{T}_{w}^{W}$ on $\mathscr{H}$ equals to $\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}\mathcal{T}_{w}$ (see proof of Theorem 3.10 in Brezis (2011)).
+As above, on each closed ball $\mathscr{H}(r):=\{\theta\in\mathscr{H}:\,\|\theta\|_{H^{\alpha}(\mathcal{O},W)}\leq r\}$, the weak topology $\mathcal{T}_{w}^{W}$ is then also metrizable by some metric $d$.
+x
+In the following we claim that $\mathscr{J}_{\mathfrak{r},N}:\Theta\cap\mathscr{H}(2^{j})\to\mathbb{R}\cup\{\infty\}$ is lower semicontinous with respect to $\mathcal{T}_{w}^{W}$ for any $j\in\mathbb{N}$.
+To this end, we need the following Lemma.
+Lemma § 4.1 (Compact Operators and Weak Convergence).
+Let $(X,\|\cdot\|_{X})$ and $(Y,\|\cdot\|_{Y})$ be two Banach spaces.
+Let $T:X\to Y$ be a linear and compact operator.
 Assume that $(x_{n})_{n\in\mathbb{N}}\subseteq X$ is a weakly convergent sequence with limit $x\in X$.
 Then, $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, i.e.
-$\lim_{n\to\infty}\left\|T[x_{n}]-T[x]\right\|_{Y}=0$.
-Proof of 1 (Lemma A.1).
-Let $(x_{n})_{n\in\mathbb{N}}\subseteq X$ be a weakly convergent sequence with limit $x\in X$, i.e., for all $L\in X^{*}$, we have $\lim_{n\to\infty}L[x_{n}]=L[x]$.
-By an application of the uniform boundness principle, there exists $B\in\mathbb{R}_{>0}$, such that $\left\|x_{n}\right\|_{X}\leq B$ for all $n\in\mathbb{N}$.
+$\lim_{n\to\infty}\|T[x_{n}]-T[x]\|_{Y}=0$.
+Proof of lemma 4.1.
+Let $(x_{n})_{n\in\mathbb{N}}\subseteq X$ be a weakly convergent sequence with limit $x\in X$, i.e., for all $L\in X^{*}$, it holds $\lim_{n\to\infty}L[x_{n}]=L[x]$.
+By an application of the uniform boundedness principle, there exists $B\in\mathbb{R}_{>0}$, such that $\|x_{n}\|_{X}\leq B$ for all $n\in\mathbb{N}$.
+In other words, $(x_{n})_{n\in\mathbb{N}}$ is bounded as well as every sub-sequence of $(x_{n})_{n\in\mathbb{N}}$.
 By hypothesis on $T$ being compact, the set $T((x_{n})_{n\in\mathbb{N}})$ is relative compact in $Y$, i.e.
 $\overline{T((x_{n})_{n\in\mathbb{N}})}$ is compact in $Y$.
-Hence, there exists a subsequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ with $\lim_{m\to\infty}\left\|T[x_{n_{m}}]-v\right\|_{Y}=0$ for some $v\in Y$.
-Further, for any $A\in Y^{*}$, we have that $A\circ T\in X^{*}$.
-Thus, $\lim_{n\to\infty}A[T[x_{n}]]=A[T[x]]$, such that the sequence $(T[x_{n}])_{n\in\mathbb{N}}\subseteq Y$ converges weakly in $Y$ with limit $T[x]\in Y$.
-By uniquess of the weak limit, we deduce $v=T[x]$.
-Overall, we have that the sequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$.
-Now, we see that each subsequence $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}\subseteq Y$ converges strongly in $Y$, which already proves the claim.
-Remark § A.2 (Embeddings and weak convergence).
+Thus, any subsequence $(T[x_{n_{m}}])_{m\in\mathbb{N}}$ of $(T[x_{n}])_{n\in\mathbb{N}}$ has a further strongly converging subsequence $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}$ with limit $y\in Y$, that is $\lim_{l\to\infty}\|T[x_{n_{m_{l}}}]-y\|_{Y}=0$.
+Further, observe that $(T[x_{n}])_{n\in\mathbb{N}}$ converges weakly in $Y$, as for any $A\in Y^{*}$ we have $A\circ T\in X^{*}$ and thus $\lim_{n\to\infty}A[T[x_{n}]]=\lim_{n\to\infty}(A\circ T)[x_{n}]=(A\circ T)[x]=A[T[x]]$.
+In other words, we have shown that $(T[x_{n}])_{n\in\mathbb{N}}$ converges weakly to $T[x]$ in $Y$ and $(T[x_{n_{m_{l}}}])_{l\in\mathbb{N}}$ converges strongly to $y$ in $Y$.
+By uniqueness of the weak limit, we deduce $y=T[x]$ and therefore $T[x_{n}]$ converges strongly to $T[x]$ in $Y$.
+Remark § 4.2 (Embeddings and Weak Convergence).
 For $d\in\mathbb{N}$ let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be a bounded domain with smooth boundary.
-Denoting by $B_{pq}^{s}(\mathcal{O})$ the ususal Besov space with parameters $p,q\in\mathbb{R}_{>0}\cup\{\infty\}$ and $s\in\mathbb{R}$, we have that the embedding
+Denoting by $B_{pq}^{s}(\mathcal{O})$ the ususal Besov space with parameters $p,q\in\mathbb{R}_{>0}\cup\{\infty\}$ and $s\in\mathbb{R}$, the embedding
 $B_{p_{1}q_{1}}^{s_{1}}(\mathcal{O})\hookrightarrow B_{p_{2}q_{2}}^{s_{2}}(\mathcal{O})$
-is compact if $p_{1},p_{2},q_{1},q_{2}\in\mathbb{R}_{>0}\cup\{\infty\}$, $-\infty<s_{2}<s_{1}<\infty$ and $v:=s_{1}-s_{2}-d\max\{\frac{1}{p_{1}}-\frac{1}{p_{2}},0\}\in\mathbb{R}_{>0}$ (see Edmunds and Triebel [1996] for definitions and properties).
+is compact, if $p_{1},p_{2},q_{1},q_{2}\in\mathbb{R}_{>0}\cup\{\infty\}$, $\infty<s_{2}<s_{1}<\infty$ and $s_{1}-s_{2}-d\max\{\frac{1}{p_{1}}-\frac{1}{p_{2}},0\}\in\mathbb{R}_{>0}$ (see Edmunds and Triebel (1996) for definitions and properties).
 Let $\alpha\in\mathbb{R}_{>0}$ and $\eta\in\mathbb{R}_{\geq 0}\setminus\mathbb{N}$.
-Identifying $H^{\alpha}(\mathcal{O})=B_{22}^{\alpha}(\mathcal{O})$, $\mathbb{L}^{2}(\mathcal{O})=B_{22}^{0}(\mathcal{O})$ and $C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$, we have in particular, that the embeddings $H^{\alpha}(\mathcal{O})\hookrightarrow\mathbb{L}^{2}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})$ for $\alpha>\eta+\frac{d}{2}$ are compact.
-Further, following Proposition 4.5 in Triebel [2006], we have for any $k\in\mathbb{N}$ the continous embedding $B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\overline{\mathcal{O}})$.
-Making use of above results, we thus have that the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is also compact, provided that $\alpha>k+\frac{d}{2}$.
+In particular, by identifying $H^{\alpha}(\mathcal{O})=B_{22}^{\alpha}(\mathcal{O})$, $\mathbb{L}^{2}(\mathcal{O})=B_{22}^{0}(\mathcal{O})$ and $C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$, the embeddings $H^{\alpha}(\mathcal{O})\hookrightarrow\mathbb{L}^{2}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})$ for $\alpha>\eta+\frac{d}{2}$ are compact.
+Further, following Proposition 4.5 in Triebel (2006), for any $k\in\mathbb{N}$ the embedding $B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\overline{\mathcal{O}})$ is continuous.
+Hence, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is also compact, provided that $\alpha>k+\frac{d}{2}$.
 If $(W,\left|\cdot\right|_{W})$ denotes a $d_{W}$-dimensional $\mathbb{R}$-vector space, the same results apply, if we replace the function spaces $F(\mathcal{O})$ by their tensor products $F(\mathcal{O},W)$.
-Now let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\mathscr{H}$ be a sequence, such that $\theta_{n}\to\theta\in\mathscr{H}$ with respect to $\mathcal{T}_{w}^{W}$.
-Then, an application of Lemma A.1 yields that $(\theta_{n})_{n\in\mathbb{N}}$ converges strongly in $\mathbb{L}^{2}(\mathcal{O},W)$ as well as in $C^{\eta}(\mathcal{O},W)$, if $\alpha>\eta+\frac{d}{2}$ for $\eta\in\mathbb{R}_{\geq 0}$.
-We derive that the forward map $\mathscr{G}$ is continuous, which is necessary to prove the lower semicontinuity of $-\mathscr{J}_{\mathfrak{r},N}$.
-Lemma § A.3 (Continuity of $\mathscr{G}$).
-For $\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$ let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C3) of 2.3.
+With $\mathscr{H}$ as defined above, let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\mathscr{H}$ be a sequence, such that $\theta_{n}\to\theta\in\mathscr{H}$ with respect to $\mathcal{T}_{w}^{W}$.
+An application of lemma 4.1 then yields that $(\theta_{n})_{n\in\mathbb{N}}$ converges strongly in $\mathbb{L}^{2}(\mathcal{O},W)$ as well as in $C^{\eta}(\mathcal{O},W)$, if $\alpha>\eta+\frac{d}{2}$ for $\eta\in\mathbb{R}_{\geq 0}$.
+Lemma § 4.3 (Continuity of $\mathscr{G}$).
+For $\gamma_{2},\eta_{1},\eta_{2}\in\mathbb{R}_{\geq 0}$, let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C3) of 2.3.
 Let $\Theta_{o}\subseteq\Theta$ be a bounded subset of $\mathscr{H}$ with $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$.
@@
 Then, the map
-$(\Theta_{o},d)\ni\theta\mapsto\mathscr{G}(\theta)(Z)\in(V,\left|\cdot\right|_{V})$
+$(\Theta_{o},d)\ni\theta\mapsto\mathscr{G}(\theta)(Z)\in(V,|\cdot|_{V})$
 is continuous $\mathbb{P}$-almost surely.
-Proof of 2 (Lemma A.3).
+Proof of lemma 4.3.
 Fix some $\omega\in\Omega$ (outside of a $\mathbb{P}$-null-set).
 Let $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta_{o}$ be a convergent sequence with weak limit $\theta\in\Theta_{o}$, i.e., $\lim_{n\to\infty}d(\theta_{n},\theta)=0$.
-Note that $\left\|\theta_{n}\right\|_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}\leq B$ for some $B\in\mathbb{R}_{>0}$ uniformly over all $n\in\mathbb{N}$.
+Note that $\|\theta_{n}\|_{H^{\alpha}(\mathcal{O},W)}\lor\|\theta\|_{H^{\alpha}(\mathcal{O},W)}\leq B$ for some $B\in\mathbb{R}_{>0}$ uniformly over all $n\in\mathbb{N}$.
 The claim is proven, if $(\varphi_{n}(Z(\omega)))_{n\in\mathbb{N}}\subseteq\mathbb{R}$, defined by
-$\varphi_{n}(Z(\omega)):=\left|\mathscr{G}(\theta_{n})(Z(\omega))-\mathscr{G}(\theta)(Z(\omega))\right|_{V}$
+$\varphi_{n}(Z(\omega)):=|\mathscr{G}(\theta_{n})(Z(\omega))-\mathscr{G}(\theta)(Z(\omega))|_{V}$
 is a null-sequence.
-By hypothesis C3) and the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$, there exists a universal constant $c\in\mathbb{R}_{>0}$, such that
-$\varphi_{n}(Z(\omega))\leq\left\|\mathscr{G}(\theta_{n})-\mathscr{G}(\theta)\right\|_{\infty}$ $\leq c(1+B^{\gamma_{2}})\left\|\theta_{n}-\theta\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
-Exploiting Remark A.2, the right hand side converges to zero, which proves the claim.
-Lemma § A.4 (Uniform continuity of $\mathscr{G}$).
-Let $\mathscr{G}$ satisfy C2) and C3) of 2.3.
-Let $D_{N}=((Y_{i},Z_{i}))_{i\in\left\llbracket N\right\rrbracket}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for $N\in\mathbb{N}$.
+By hypothesis and exploiting the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$, there exists a constant $c\in\mathbb{R}_{>0}$ (depending on $\alpha,\eta_{1},\eta_{2},d,C_{\operatorname{Lip},\infty}$), such that
+$\varphi_{n}(Z(\omega))\leq\|\mathscr{G}(\theta_{n})-\mathscr{G}(\theta)\|_{\infty}$ $\leq c\cdot(1+B^{\gamma_{2}})\cdot\|\theta_{n}-\theta\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
+Exploiting remark 4.2, we have $\lim_{n\to\infty}\|\theta_{n}-\theta\|_{C^{\eta_{2}}(\mathcal{O},W)=0}$, which proves the claim. ∎
+Lemma § 4.4 (Uniform Continuity).
+Let $\mathscr{G}:\Theta\to\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)$ satisfy C2) and C3) of 2.3.
+Let $D_{N}=((Y_{i},Z_{i}))_{i\in\left\llbracket N\right\rrbracket}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$ for $N\in\mathbb{N}$ as in theorem 2.6.
 Let $r\in\mathbb{R}_{>0}$.
 Then, the map
-$E:(\Theta\cap\mathscr{H}(r),d)\to(\mathbb{R},|\cdot|),\,\theta\mapsto E[\theta]:=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
+$E:(\Theta\cap\mathscr{H}(r),d)\to(\mathbb{R},|\cdot|),\,\theta\mapsto E[\theta]:=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}$
 is almost surely uniformly continuous.
-Proof of 3 (Lemma A.4).
-For any $\delta\in\mathbb{R}_{>0}$, let $M_{\delta}$ be the modulus of continuity of $E$, i.e.,
-$M_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|E[\theta]-E[\tilde{\theta}]|\right\},$
+Proof of lemma 4.4.
+For any $\delta\in\mathbb{R}_{>0}$, denote by $M_{\delta}$ the modulus of continuity of $E$, i.e.,
+[…]
 which defines a random variable. If $(\Omega,\mathcal{A},\mathbb{P})$ denotes the underlying probability space supporting the law of $Z$ and $\varepsilon$, respectivly, we define the measurable event
-$A:=\left\{\omega\in\Omega:\,\lim_{\delta\to 0}M_{\delta}=0\right\}.$
+$A:=\left\{\omega\in\Omega:\,\lim_{\delta\to 0}M_{\delta}(\omega)=0\right\}.$
 We aim to show $\mathbb{P}(A)=1$. As $M_{\delta}$ is decreasing in $\delta$, it suffices to verify that $\lim_{\delta\to 0}\mathbb{E}[M_{\delta}]=0$. Indeed, elementary computations show that
-$|E[\theta]-E[\tilde{\theta}]|$ $\leq\frac{2C_{\mathscr{G},\operatorname{B}}}{N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}+\frac{1}{N\sigma}\sum_{i\in\left\llbracket N\right\rrbracket}|\langle\varepsilon_{i},\mathscr{G}(\tilde{\theta})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\rangle_{V}|.$
-Analogously, define for each $\delta\in\mathbb{R}_{>0}$ and $Z\sim\zeta$ the modulus of continuity of $\mathscr{G}$ by
-$R_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{\left|\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}\right\},$
-which in turns defines a random variable.
-In Lemma A.3, we have shown that $\mathscr{G}$ is almost sure continuous as a mapping from $\Theta\cap\mathscr{H}(r)$ to $V$.
+[…]
+Define for each $\delta\in\mathbb{R}_{>0}$ and $Z\sim\zeta$ the modulus of continuity of $\mathscr{G}$ by
+$R_{\delta}:=\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)|_{V}\right\},$
+which in turn defines also a random variable.
+In lemma 4.3, we have shown that $\mathscr{G}$ is almost surely continuous as a mapping from $\Theta\cap\mathscr{H}(r)$ to $V$.
 As $\Theta\cap\mathscr{H}(r)$ is a compact metric space, $\mathscr{G}$ is in fact almost surely uniformly continuous, and hence $\lim_{\delta\to 0}R_{\delta}=0$ almost surely.
 Moreover, since $Z$ and $\varepsilon\sim N(0,\operatorname{Id}_{V})$ are independent by hypothesis, we have
-$\mathbb{E}\left[\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\langle\varepsilon,\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\rangle_{V}|\right\}\right]$ $\leq\mathbb{E}[\left|\varepsilon\right|_{V}R_{\delta}]=\sqrt{d_{V}}\cdot\mathbb{E}[R_{\delta}].$
+$\mathbb{E}\left[\sup_{\theta,\tilde{\theta}\in\Theta\cap\mathscr{H}(r):d(\theta,\tilde{\theta})\leq\delta}\left\{|\langle\varepsilon,\mathscr{G}(\tilde{\theta})(Z)-\mathscr{G}(\theta)(Z)\rangle_{V}|\right\}\right]$ $\leq\mathbb{E}[|\varepsilon|_{V}R_{\delta}]=\sqrt{d_{V}}\cdot\mathbb{E}[R_{\delta}].$
 Now by exploiting C2), $(R_{\delta})_{\delta\in\mathbb{R}_{>0}}$ is uniformly integrable and hence, the last expectation is converging to zero for $\delta\to 0$. Overall, we obtain
 $\mathbb{E}[M_{\delta}]\leq\left(\frac{2C_{\mathscr{G},\operatorname{B}}}{\sigma^{2}}+\frac{\sqrt{d_{V}}}{\sigma}\right)\mathbb{E}[R_{\delta}]\overset{\delta\to 0}{\longrightarrow}0,$
-which proves the claim.
-We are now providing the proof of Proposition 2.5.
-Proof of 4 (Proposition 2.5).
-Grant the assumptions of Proposition 2.5. We follow along the lines of Nickl et al. [2020] and split the proof of existence in two parts.
+which shows the claim. ∎
+Proof of proposition 2.5.
+Grant the assumptions of proposition 2.5. We follow along the lines of Nickl et al. (2020) and split the proof of existence in two parts.
   1. i)
@@
 We show now that $\mathscr{J}_{\mathfrak{r},N}$ has a maximizer over $\Theta\cap\mathscr{H}(2^{j})$.
-As $\Theta$ is assumed to be weakly closed and $\mathscr{H}(2^{j})$ is weakly sequentially compact, due to the separability and reflexivity of Sobolev spaces and by an applycation of the Banach-Alaoglu theorem, any sequence $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta\cap\mathscr{H}(2^{j})$ has a weakly convergent sub-sequence $(\theta_{n_{k}})_{k\in\mathbb{N}}$ with weak limit $\theta\in\Theta\cap\mathscr{H}(2^{j})$.
-Thus, following the Direct Method of the Calculus of Variations, it suffices to show that $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ is almost surely weakly lower semicontinuous.
-Rewrite $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ as
-    $-\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}(\theta)=E_{1}(\theta)+E_{2}(\theta)$ (A.1)
+As $\Theta$ is assumed to be weakly closed and $\mathscr{H}(2^{j})$ is weakly sequentially compact, due to the separability and reflexivity of Sobolev spaces and by an application of the Banach-Alaoglu theorem, any sequence $(\theta_{n})_{n\in\mathbb{N}}\subseteq\Theta\cap\mathscr{H}(2^{j})$ has a weakly convergent sub-sequence $(\theta_{n_{k}})_{k\in\mathbb{N}}$ with weak limit $\theta\in\Theta\cap\mathscr{H}(2^{j})$.
+Thus, following the Direct Method of the Calculus of Variations, it suffices to show that $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ is almost surely weakly lower semicontinuous.
+Rewrite $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}$ as
+    $\mathscr{J}_{\mathfrak{r},N}|_{\Theta\cap\mathscr{H}(2^{j})}(\theta)=E_{1}(\theta)+E_{2}(\theta)$ (17)
     with
-    $E_{1}(\theta):=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\text{ and }E_{2}(\theta):=\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
+    $E_{1}(\theta):=\frac{1}{2N\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\text{ and }E_{2}(\theta):=\frac{\mathfrak{r}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2}.$
     $E_{2}$ is weakly lower semicontinuous, which is a standard result in functional analysis concerning norms on Banach spaces.
-In Lemma A.4, we have derived that $E_{1}$ is almost surely uniformly continuous as a map between the metric spaces $(\Theta\cap\mathscr{H}(2^{j}),d)$ and $(\mathbb{R},|\cdot|)$.
+In lemma 4.4, we have derived that $E_{1}$ is almost surely uniformly continuous as a map between the metric spaces $(\Theta\cap\mathscr{H}(2^{j}),d)$ and $(\mathbb{R},|\cdot|)$.
 In particular, $E_{1}$ is weakly lower semicontinuous.
-All in all, local existence of $\hat{\theta}_{\textsc{MAP}}$ follows immediately by an application of the Direct Methods of Calculus of Variations (see Dacorogna [2008]).
+All in all, local existence of $\hat{\theta}_{\textsc{MAP}}$ follows immediately by an application of the Direct Methods of Calculus of Variations (see Dacorogna (2008)).
   2. ii)
@@
     $A_{j}:=\left\{\omega\in\Omega:\,\hat{\theta}_{\textsc{MAP}}(\omega)\notin\Theta\cap\mathscr{H}(2^{j})\right\}.$
-    Then, $(A_{j})_{j\in\mathbb{N}}$ satisfies $\lim_{j\to\infty}\mathbb{P}[A_{j}]=0$ according to Theorem 2.6. Thus, choosing $j\in\mathbb{N}$ large enough, we have
-    $\sup_{\theta\in\Theta\cap\mathscr{H}(2^{j})}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}=\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[F]\right\}$
+    Then, $(A_{j})_{j\in\mathbb{N}}$ satisfies $\lim_{j\to\infty}\mathbb{P}[A_{j}]=0$ according to the concentration inequality in theorem 2.6 i). Thus, choosing $j\in\mathbb{N}$ large enough, we have
+    $\sup_{\theta\in\Theta\cap\mathscr{H}(2^{j})}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}=\sup_{\theta\in\Theta}\left\{\mathscr{J}_{\mathfrak{r},N}[\theta]\right\}$
     with probability as close to one as desired.
-We now approaching the proof of Theorem 2.6. Therefore, we need some notation and further helping lemma.
-Notation § A.5 (Empirical Measures).
-For $N\in\mathbb{N}$ let $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ be a data set.
-For any $(y,z)\in V\times\mathcal{Z}$ denote by $\delta_{(y,z)}(\cdot)$ the Dirac measure on $V\times\mathcal{Z}$.
-The empirical measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ based on $D_{N}$ is then defined by
-^$\widehat{\mathbb{P}}_{N}(A):=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}\delta_{(Y_{i},Z_{i})}(A)$
-for any $A\in\mathscr{B}_{V\times\mathcal{Z}}$. Evidently, we have for any measurable function $h:V\times\mathcal{Z}\to\mathbb{R}$
+We are now approaching the proof of theorem 2.6. Therefore, we need some notations and further helping lemma.
+Notation § 4.5 (Empirical Measures).
+For $N\in\mathbb{N}$, let $D_{N}:=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ be given data.
+For any $(y,z)\in V\times\mathcal{Z}$, denote by $\delta_{(y,z)}(\cdot)$ the Dirac measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ in $(y,z)$.
+The empirical measure on $(V\times\mathcal{Z},\mathscr{B}_{V\times\mathcal{Z}})$ based on $D_{N}$ is then for any $A\in\mathscr{B}_{V\times\mathcal{Z}}$ defined by
+$\widehat{\mathbb{P}}_{N}(A):=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}\delta_{(Y_{i},Z_{i})}(A).$
+Evidently, we have for any measurable function $h:V\times\mathcal{Z}\to\mathbb{R}$
 $\int_{V\times\mathcal{Z}}h(y,z)\mathrm{d}\widehat{\mathbb{P}}_{N}(y,z)=\frac{1}{N}\sum_{i\in\left\llbracket N\right\rrbracket}h(Y_{i},Z_{i}).$
-The following helping lemma provides M-estimation techniques, which are used to prove Theorem 2.6.
-The proof follows along the lines of van de Geer [2001], Nickl et al.
-[2020] and Bohr and Nickl [2023].
-Lemma § A.6 (M-estimation techniques).
-With the notation of Theorem 2.6 and the lines above it, we have for any $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ and $\tilde{\theta}\in\Theta$, $\mathfrak{r}\in\mathbb{R}_{>0}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$
-$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$ (A.2)
+The following Lemma provides M-estimation techniques, which are used to prove theorem 2.6. The proof follows along the lines of van de Geer (2001), Nickl et al. (2020) and Bohr and Nickl (2023).
+Lemma § 4.6 (M-Estimation Techniques).
+With the notation of theorem 2.6 and the lines above it, for any $\theta_{\mathfrak{o}}\in\tilde{\Theta}$, $\tilde{\theta}\in\Theta$, $\mathfrak{r}\in\mathbb{R}_{>0}$ and any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$, we have
+$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$ (18)
 $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-almost surely. Moreover, for any $R,\mathfrak{r}\in\mathbb{R}_{>0}$ and some fixed constant $c_{0}\in\mathbb{R}_{\geq 2}$ define the measurable event
 $\Xi_{\mathfrak{r},R}:=\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{0}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}.$
-On $\Xi_{\mathfrak{r},R}$, we have further
-$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}\cdot R^{2}$ (A.3)
+On $\Xi_{\mathfrak{r},R}$, we further have
+$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}\cdot R^{2}$ (19)
 for some $c_{1}\in\mathbb{R}_{\geq 1}$, as well as
-$\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})-\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})\geq\frac{1}{6}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}).$ (A.4)
-Proof of 5 (Lemma A.6).
+[…]
+Proof of lemma 4.6.
 The proof of eq.
-A.2 is elementary and follows along the lines of Lemma 4.10 in Nickl and Wang [2024] (replacing the finite-dimensional approximation of $\theta_{\mathfrak{o}}$ by some arbitrary $\tilde{\theta}\in\Theta$).
+18 is elementary and follows along the lines of Lemma 4.10 in Nickl and Wang (2024) (replacing the finite-dimensional approximation of $\theta_{\mathfrak{o}}$ by some arbitrary $\tilde{\theta}\in\Theta$).
 We continue with the proof of eq.
-A.3.
+19.
 On the event $\Xi_{\mathfrak{r},R}$, we have
 $c_{0}\cdot R^{2}$ $\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})-c_{0}\cdot\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})$
-$\leq 2\left(\frac{1}{2}\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\ -\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right)+\mathfrak{r}^{2}\left(\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right)$
-$\leq 2\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}-\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right)^{2}+2\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-\mathfrak{r}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\mathfrak{r}^{2}\left\|\tilde{\theta}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$
-$\leq 2\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}),$
+$\leq 2\left(\frac{1}{2}\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}-\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\right)$
+$\hskip 170.71652pt+\mathfrak{r}^{2}\left(\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O},W)}^{2}\right)$
+$\leq 2\left(\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}-\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right)^{2}$
+$\hskip 113.81102pt+2\mathfrak{r}^{2}\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-\mathfrak{r}^{2}\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}-2\mathfrak{r}^{2}\|\tilde{\theta}\|_{H^{\alpha}(\mathcal{O},W)}^{2}$
+[…]
 where we have used $\frac{1}{2}a^{2}-b^{2}\leq(a-b)^{2}$ for two real numbers $a,b\in\mathbb{R}$ as well as the reversed triangle inequality.
 Rearranging, the last inequality shows eq.
-A.3.
+19.
 eq.
-A.4 follows by the same elementary computations as in van de Geer [2001] (see pp.
+20 follows by the same elementary computations as in van de Geer (2001) (see pp.
 3-4).
-Proof of 6 (Theorem 2.6).
-The proof follows the key ideas of Bohr and Nickl [2023].
-Since we are allowing a more general $\tilde{\theta}\in\Theta$, we also need to adapt the ideas of Nickl et al.
-[2020].
+∎
+Proof of theorem 2.6.
+The proof follows the key ideas of Bohr and Nickl (2023).
+Since we are allowing a general $\tilde{\theta}\in\Theta$, we also need to adapt the ideas of Nickl et al.
+(2020).
 We start with the exponential inequality eq.
-2.4.
+11.
 Fix $\tilde{\theta}\in\Theta$.
-According to Lemma A.6, it sufficies to controll the $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability of the event
+According to lemma 4.6, it sufficies to control the $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$-probability of the event
 $\tilde{\Xi}_{\mathfrak{r},R}:=\Xi_{\mathfrak{r},R}\cap\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\leq 2\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\hat{\theta}_{\textsc{MAP}}}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)+\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})\right\}.$
@@
 $\mathcal{W}_{N}(\theta):=\int_{V\times\mathcal{Z}}\log\left(\frac{p_{\theta}(y,z)}{p_{\tilde{\theta}}(y,z)}\right)\mathrm{d}(\widehat{\mathbb{P}}_{N}-\mathbb{P}_{\theta_{\mathfrak{o}}})(y,z)$
-and obtain with Eq. A.3 and Eq. A.4 from Lemma A.6
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\tilde{\Xi}_{\mathfrak{r},R})\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}^{2}R^{2},\,\frac{1}{12}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\leq|\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})|\right\}\right).$
+and obtain with eq. 19 and eq. 20 from lemma 4.6
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\tilde{\Xi}_{\mathfrak{r},R})\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(|\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})|\geq\frac{1}{12}\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta}),\,\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\tilde{\theta})\geq c_{1}^{2}R^{2},\right).$
 for some $c_{1}\in\mathbb{R}_{\geq 1}$.
-In the following step, we apply a so-called peeling device, where we refer to van de Geer [2000] for a generic explanatation.
+In the following step, we apply a so-called peeling device, where we refer to van de Geer (2000) for a generic explanation.
 This method was already used in a similar way in Nickl et al.
-[2020] and Bohr and Nickl [2023] to controll the process $\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})$ over disjoint slices of a subset of $\Theta$.
+(2020) and Bohr and Nickl (2023) to control the process $\mathcal{W}_{N}(\hat{\theta}_{\textsc{MAP}})$ over disjoint slices of $\Theta$.
 In detail, for all $l\in\mathbb{N}$ and $c_{1}$ sufficiently large, define the family of slices
-[…]
+$\Theta_{l}:=\left\{\theta\in\Theta:\,2^{2(l-1)}R^{2}c_{1}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})<2^{2l}R^{2}c_{1}^{2}\right\}\subseteq H^{\alpha}(\mathcal{O},W,2^{l}R\mathfrak{r}^{-1}c_{1}),$
 where the inclusion follows from the definition of $\mathfrak{d}_{\mathfrak{r}}^{2}$.
@@
 We obtain
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right).$
-We follow the approach of Bohr and Nickl [2023] and decompose $\mathcal{W}_{N}$ into two empirical processes. Indeed, for each $\theta\in\Theta$, simple computations show
-$N\mathcal{W}_{N}(\theta)$ $=\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})],$
-which evidently defines a centered process. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, we further decompose the log-likelihood difference of the last display by
-$\ell_{N}(\theta)-\ell_{N}(\tilde{\theta})$ $=\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|Y_{i}-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}$
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right).$
+We follow the approach of Bohr and Nickl (2023) and decompose $\mathcal{W}_{N}$ into two empirical processes. Indeed, for each $\theta\in\Theta$, simple computations show
 […]
+which evidently defines a centered stochastic process. Given the data $D_{N}\sim\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$, we further decompose the log-likelihood difference of the last display by
+[…]
+$\overset{d}{=}\frac{1}{\sigma}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V}$
+$\phantom{==}+\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\right\}.$
 Defining for all $\theta\in\Theta$
-$T_{1}(\theta):=\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V},$ (A.5)
+$T_{1}(\theta):=\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\langle\varepsilon_{i},\mathscr{G}(\theta)(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\rangle_{V},$ (21)
 and
-$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\},$ (A.6)
-respectively, we have that $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{1}(\theta)]=0$, as $Z$ and $\varepsilon$ are independent and $\varepsilon$ is centered. In other words, we have derived so far
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)$ $\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right\}\right)$
-$\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)+\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right).$
-We will discuss the two different probabilties separately and starting with the the multiplier-type process defined in Eq.
-A.5.
-To apply Lemma A.7, we denote for each $j\in\left\llbracket d_{V}\right\rrbracket$ by $\mathscr{G}(\cdot)_{j}:\Theta\to\mathbb{L}^{2}(\mathcal{O})$ the $j$-th coordinate of the forward map $\mathscr{G}:\Theta\to\mathbb{L}^{2}(\mathcal{O},V)$, where we again identify $V$ with $\mathbb{R}^{d_{V}}$.
+$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}\right\},$ (22)
+respectively, we have that $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{1}(\theta)]=0$, as $Z$ and $\varepsilon$ are independent and $\varepsilon$ is centered. With these, we have derived so far
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\Xi_{\mathfrak{r},R}\right)$ $\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|\mathcal{W}_{N}(\theta)|\geq\frac{2^{2l}}{48}R^{2}c_{1}^{2}\right)$
+$\leq\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)$
+$\phantom{==}+\sum_{l\in\mathbb{N}}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right).$
+We will discuss the two different probabilities separately and starting with the multiplier-type process defined in eq.
+21.
+In order to apply lemma 4.7, we denote for each $j\in\left\llbracket d_{V}\right\rrbracket$ by $\mathscr{G}(\cdot)_{j}:\Theta\to\mathbb{L}^{2}(\mathcal{O})$ the $j$-th coordinate of the forward map $\mathscr{G}:\Theta\to\mathbb{L}^{2}(\mathcal{O},V)$, where we again identify $V$ with $\mathbb{R}^{d_{V}}$.
 We define a class of measurable functions by
-$\mathcal{H}_{l,j}:=\left\{\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}:\mathcal{Z}\to\mathbb{R}:\,\theta\in\Theta_{l}\right\}.$ (A.7)
-To shorten notation in the subsequent, we write $h_{\theta,j}:=\mathscr{G}(\theta_{\mathfrak{o}})_{j}-\mathscr{G}(\theta)_{j}$. Thus, for any $\theta\in\Theta$, we have
-[…]
-where $\varepsilon_{ij}\overset{i.i.d.}{\sim}\operatorname{N}(0,1)$ for $(i,j)\in\left\llbracket N\right\rrbracket\times\left\llbracket d_{V}\right\rrbracket$. Conclusevly, we see
+$\mathcal{H}_{l,j}:=\left\{\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}:\mathcal{Z}\to\mathbb{R}:\,\theta\in\Theta_{l}\right\}.$ (23)
+To shorten notation in the subsequent, we write $h_{\theta,j}:=\mathscr{G}(\tilde{\theta})_{j}-\mathscr{G}(\theta)_{j}$. Thus, for any $\theta\in\Theta$, we have
+$T_{1}(\theta)$ $\overset{d}{=}-\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}(\mathscr{G}(\tilde{\theta})_{j}(Z_{i})-\mathscr{G}(\theta)_{j}(Z_{i}))\varepsilon_{ij}$
+$=-\frac{1}{\sigma\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}h_{\theta,j}(Z_{i})\varepsilon_{ij}=:-\frac{1}{\sigma}\sum_{j\in\left\llbracket d_{V}\right\rrbracket}Z_{N,j}^{(1)}(\theta),$
+where $\varepsilon_{ij}\overset{i.i.d.}{\sim}\operatorname{N}(0,1)$ for $(i,j)\in\left\llbracket N\right\rrbracket\times\left\llbracket d_{V}\right\rrbracket$. Conclusively, we see
 $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|T_{1}(\theta)|\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)\leq d_{V}\max_{j\in\left\llbracket d_{V}\right\rrbracket}\left\{\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{2^{2l}}{96d_{V}}R^{2}c_{1}^{2}\sigma\right)\right\}.$
-We now verify that the function class $\mathcal{H}_{l,j}$ defined in Eq. A.7 satisfies the assumptions formulated in Lemma A.7. Indeed, by C2) in 2.3, we have uniformly for all $\theta\in\Theta_{l}$
-$\sup_{z\in\mathcal{Z}}\left\{|h_{\theta,j}(z)|\right\}=\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})_{j}(z)-\mathscr{G}(\theta)_{j}(z)|\right\}\leq\sup_{z\in\mathcal{Z}}\left\{\left|\mathscr{G}(\tilde{\theta})(z)-\mathscr{G}(\theta)(z)\right|_{V}\right\}\leq 2C_{\mathscr{G},\operatorname{B}}.$
+We now verify that the function class $\mathcal{H}_{l,j}$ defined in eq. 23 satisfies the assumptions formulated in lemma 4.7. Indeed, by C2) in 2.3, we have uniformly for all $\theta\in\Theta_{l}$
+$\sup_{z\in\mathcal{Z}}\left\{|h_{\theta,j}(z)|\right\}=\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})_{j}(z)-\mathscr{G}(\theta)_{j}(z)|\right\}\leq\sup_{z\in\mathcal{Z}}\left\{|\mathscr{G}(\tilde{\theta})(z)-\mathscr{G}(\theta)(z)|_{V}\right\}\leq 2C_{\mathscr{G},\operatorname{B}}.$
 Moreover, for $Z\sim\zeta$ and all $\theta\in\Theta_{l}$, $l\in\mathbb{N}$, we have
-$\mathbb{E}^{Z}[h_{\theta,j}^{2}(Z)]\leq\sum_{j\in\left\llbracket d_{V}\right\rrbracket}\int_{\mathcal{Z}}\left|\mathscr{G}(\theta)_{j}(Z)-\mathscr{G}(\tilde{\theta})_{j}(Z)\right|_{V}^{2}\mathrm{d}\zeta(z)=\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}.$
+$\mathbb{E}^{Z}[h_{\theta,j}^{2}(Z)]\leq\sum_{j\in\left\llbracket d_{V}\right\rrbracket}\int_{\mathcal{Z}}|\mathscr{G}(\theta)_{j}(z)-\mathscr{G}(\tilde{\theta})_{j}(z)|_{V}^{2}\mathrm{d}\zeta(z)=\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}.$
 Observing that $\theta\in\Theta_{l}$, the last expression is further upper bounded by
-$\left\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\leq 2^{2l}R^{2}c_{1}^{2}=:v_{l}^{2}.$
+$\|\mathscr{G}(\theta)-\mathscr{G}(\tilde{\theta})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq\mathfrak{d}_{\mathfrak{r}}^{2}(\theta,\tilde{\theta})\leq 2^{2l}R^{2}c_{1}^{2}=:v_{l}^{2}.$ (24)
 Additionally, we need to upper bound the $\mathbb{L}^{2}$- and $\mathbb{L}^{\infty}$-metric entropy integrals $\mathcal{J}_{2}$ and $\mathcal{J}_{\infty}$.
@@
 $\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right),$
-where $d_{2}$ and $d_{\infty}$ are defined according to Lemma A.7. Starting with the first one, let $\theta_{1},\theta_{2}\in\Theta_{l}$ be arbitrarily chosen. Exploiting C1) in 2.3, we have
-$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\cdot\left(1+\left\|\theta_{1}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\lor\left\|\theta_{2}\right\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
-By definition of $\Theta_{l}$, $\theta_{1}$, $\theta_{2}$ are contained in a $H^{\alpha}$-ball, such that the polynomial growth in the last display can be upper bounded.
-It suffices to find a $\rho$-covering of $\Theta_{l}$ with respect to the dual $(H^{\kappa}(\mathcal{O},W))^{*}$-norm.
-Applying Lemma A.9, there exists a constant $c_{2}=c_{2}(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that for all $\rho\in\mathbb{R}_{>0}$
-$N(\Theta_{l},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}\right).$
-Consequently, for every $\rho\in\mathbb{R}_{>0}$ there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with $N$ upper bounded by the line above.
-Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
-$d_{2}(\theta,\theta^{(i_{\circ})})=\left\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)\cdot\rho=:m_{1}\cdot\rho.$
-Thus,
+where $d_{2}$ and $d_{\infty}$ are defined according to lemma 4.7. Starting with the first one, let $\theta_{1},\theta_{2}\in\Theta_{l}$ be arbitrarily chosen. Exploiting C1) in 2.3, we have
+$d_{2}(\theta_{1},\theta_{2})=\|h_{\theta_{1},j}-h_{\theta_{2},j}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}\hskip-2.15277pt\leq C_{\operatorname{Lip},2}\left(1+\|\theta_{1}\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\hskip-4.30554pt\lor\|\theta_{2}\|^{\gamma_{1}}_{H^{\alpha}(\mathcal{O},W)}\right)\hskip-2.15277pt\|\theta_{1}-\theta_{2}\|_{(H^{\kappa}(\mathcal{O},W))^{*}}.$
+By definition of $\Theta_{l}$, $\theta_{1}$, $\theta_{2}$ are contained in a $H^{\alpha}$-ball, such that the polynomial growth in the last display can be upper bounded by the corresponding radius.
+Hence, it suffices to find a $\rho$-covering of $\Theta_{l}$ with respect to the dual $(H^{\kappa}(\mathcal{O},W))^{*}$-norm.
+Applying lemma 4.9 i), there exists a constant $c_{2}=c_{2}(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that for all $\rho\in\mathbb{R}_{>0}$
+$N(\Theta_{l},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}\right).$
+As a conseqence, for every $\rho\in\mathbb{R}_{>0}$, there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with respect to $\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}}$ with $N=N(\Theta_{l},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)$.
+Hence, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
+$d_{2}(\theta,\theta^{(i_{\circ})})=\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$ $\leq C_{\operatorname{Lip},2}\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)\cdot\rho=:m_{1}\cdot\rho$
+and we deduce
 $\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)\leq c_{2}\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}m_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}=c_{3}\cdot\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha+\kappa}}$
@@
 Indeed, for any $\theta_{1},\theta_{2}\in\Theta_{l}$, exploiting C3) in 2.3, we have
-$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1},j}-h_{\theta_{2},j}\right\|_{\infty}\leq C_{\operatorname{Lip},\infty}\cdot\left(1+\left\|\theta_{1}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\left\|\theta_{2}\right\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
-Using that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$, the polynomial growth can be bounded due to the embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$.
-Thus, it suffices to find an covering for $\Theta_{l}$ with respect to the $C^{\eta_{2}}$-norm.
-Indeed, by an application of Lemma A.9, we have
-$N(\Theta_{l},\left\|\cdot\right\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)\leq\exp\left(c(\alpha,\eta_{2},d)\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\right).$
-Now, for $\rho\in\mathbb{R}_{>0}$ fixed, let $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ be a covering of $\Theta_{l}$ w.r.t.
-$\left\|\cdot\right\|_{C^{\eta_{2}}(\mathcal{O},W)}$ with $N$ bounded by the line above.
+$d_{\infty}(\theta_{1},\theta_{2})=\|h_{\theta_{1},j}-h_{\theta_{2},j}\|_{\infty}\leq C_{\operatorname{Lip},\infty}\left(1+\|\theta_{1}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\lor\|\theta_{2}\|_{C^{\eta_{1}}(\mathcal{O},W)}^{\gamma_{2}}\right)\|\theta_{1}-\theta_{2}\|_{C^{\eta_{2}}(\mathcal{O},W)}.$
+Using that $\alpha>\eta_{1}\lor\eta_{2}+\frac{d}{2}$, the polynomial growth can be bounded due to the continuous embedding $H^{\alpha}(\mathcal{O},W)\hookrightarrow C^{\eta_{1}}(\mathcal{O},W)$.
+Hence, it suffices to find a covering for $\Theta_{l}$ with respect to the $C^{\eta_{2}}$-norm.
+By an application of lemma 4.9 ii), we have
+$N(\Theta_{l},\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)\leq\exp\left(c(\alpha,\eta_{2},d)\cdot d_{W}\left(\frac{2^{l}R\sqrt{d_{W}}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\right).$
+As a conseqence, for every $\rho\in\mathbb{R}_{>0}$, there exists a $\rho$-covering $\{\theta^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}$ of $\Theta_{l}$ with respect to $\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}}$ with $N=N(\Theta_{l},\|\cdot\|_{C^{\eta_{2}(\mathcal{O},W)}},\rho)$.
 Thus, for all $\theta\in\Theta_{l}$, there exists some $i_{\circ}\in\left\llbracket N\right\rrbracket$ with
-$d_{\infty}(\theta,\theta^{(i_{\circ})})\leq c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right)\cdot\rho.$
-Thus, there exists a constant $c_{4}=c_{4}(\alpha,\eta_{2},d,d_{W})\in\mathbb{R}_{>0}$, such that
-$\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\leq c_{4}\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta}}\text{ with}m_{2}:=c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right).$
-We are now able to upper bound the corresponding entropy integrals. Indeed, as $2(\alpha+\kappa)>d$, we have
-$\mathcal{J}_{2}(\mathcal{H}_{l,j})=\int_{0}^{4v_{l}}\sqrt{\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)}\mathrm{d}\rho\leq\sqrt{c_{3}}\cdot\int_{0}^{4v_{l}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{2(\alpha+\kappa)}}\mathrm{d}\rho=\sqrt{c_{3}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}},$
+$d_{\infty}(\theta,\theta^{(i_{\circ})})\leq{\color[rgb]{0,0,0}\|h_{\theta,j}-h_{\theta^{(i_{\circ})},j}\|_{\infty}}\leq c(C_{\operatorname{Lip},\infty},\gamma_{2})\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{2}}\right)\cdot\rho{\color[rgb]{0,0,0}=:m_{2}\cdot\rho}$
+and we deduce
+$\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\leq c_{4}\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta}}$
+for some constant $c_{4}=c_{4}(\alpha,\eta_{2},d,d_{W})\in\mathbb{R}_{>0}$. We are now able to upper bound the corresponding entropy integrals. Indeed, as $2(\alpha+\kappa)>d$, we have
+$\mathcal{J}_{2}(\mathcal{H}_{l,j})$ $=\int_{0}^{4v_{l}}\sqrt{\log N\left(\mathcal{H}_{l,j},d_{2},\rho\right)}\mathrm{d}\rho$
+$\leq\sqrt{c_{3}}\cdot\int_{0}^{4v_{l}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{2(\alpha+\kappa)}}\mathrm{d}\rho=\sqrt{c_{3}}\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}},$
 where the numerical value of constant $c_{3}$ changed without any further dependencies.
@@
 We have
-$\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}}$ $=c_{1}^{\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)^{\tau_{1}}\cdot 2^{l\tau_{1}}\cdot R^{\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}}\cdot v_{l}^{-1-\tau_{1}}$
+$\left(\frac{2^{l}Rm_{1}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{2(\alpha+\kappa)}}v_{l}^{1-{\frac{d}{2(\alpha+\kappa)}}}\hskip-4.30554pt$ $=c_{1}^{\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot\left(1+\left(\frac{2^{l}Rc_{1}}{\mathfrak{r}}\right)^{\gamma_{1}}\right)^{\tau_{1}}\cdot 2^{l\tau_{1}}\cdot R^{\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}}\cdot v_{l}^{-1-\tau_{1}}$
 $\leq c_{1}^{\tau_{1}+\gamma_{1}\tau_{1}}\cdot C_{\operatorname{Lip},2}^{\tau_{1}}\cdot v_{l}^{2}\cdot 2^{\tau_{1}}\cdot 2^{l\tau_{1}+l\tau_{1}\gamma_{1}}\cdot R^{\tau_{1}+\gamma_{1}\tau_{1}}\cdot\mathfrak{r}^{-\tau_{1}-\tau_{1}\gamma_{1}}\cdot v_{l}^{-1-\tau_{1}}$
@@
 $\mathcal{J}_{2}(\mathcal{H}_{l,j})\leq c_{5}\cdot c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot v_{l}^{2}\cdot\sqrt{N}.$
-Evidently, the upper bound of $\mathcal{J}_{2}(\mathcal{H}_{l,j})$ can be as small as desired by choosing $c_{1}$ accordingly large. Analogously, as $\alpha-\eta>d$, we have
-$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})=\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\mathrm{d}\rho\leq c_{4}\cdot\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\mathrm{d}\rho\leq c(\alpha,\eta_{2},d,d_{W},C_{\mathscr{G},\operatorname{B}})\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{\alpha-\eta_{2}}}.$
-Using Eq. C5 with similar arguments as above, one can verify analogously verify
+Analogously, as $\alpha-\eta_{2}>d$, we have
+$\mathcal{J}_{\infty}(\mathcal{H}_{l,j})$ $=\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\log N\left(\mathcal{H}_{l,j},d_{\infty},\rho\right)\mathrm{d}\rho$
+$\leq c_{4}\cdot\int_{0}^{8C_{\mathscr{G},\operatorname{B}}}\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}\rho}\right)^{\frac{d}{\alpha-\eta_{2}}}\mathrm{d}\rho\leq c(\alpha,\eta_{2},d,d_{W},C_{\mathscr{G},\operatorname{B}})\cdot\left(\frac{2^{l}Rm_{2}c_{1}}{\mathfrak{r}}\right)^{\frac{d}{\alpha-\eta_{2}}}.$
+Using eq. C5 with similar arguments as above, one can verify analogously
 $\mathcal{J}_{\infty}(\mathcal{H}_{l,j})\leq c_{6}v_{l}^{2}N$
 with some $c_{6}=c_{6}(\alpha,\eta_{1},\eta_{2},\gamma_{2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},c_{1})\in\mathbb{R}_{>0}$, which can be also as small as desired by choosing $c_{1}$ sufficiently large.
-Now in the setting of Lemma A.7, for arbitrary $\bar{c}\in\mathbb{R}_{>0}$, such that $x:=\bar{c}N2^{2l}R^{2}\in\mathbb{R}_{\geq 1}$, we have for some constant $L\in\mathbb{R}_{>0}$
-$L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\leq L\left(c_{5}\cdot v_{l}^{2}\sqrt{N}+v_{l}\sqrt{\bar{c}N2^{2l}R^{2}}+N^{-\frac{1}{2}}(c_{6}\cdot 2^{2l}R^{2}N+8C_{\mathscr{G},\operatorname{B}}\bar{c}N2^{2l}R^{2})\right)$
-$\leq\frac{\sqrt{N}v_{l}^{2}}{96\cdot d_{V}}\left(c_{5}\cdot c_{1}^{-(1-\gamma_{1}\tau_{1})}\cdot d_{V}\cdot 96\cdot L+L\cdot 96\cdot d_{V}\frac{\sqrt{\bar{c}}}{\sqrt{c_{1}}}+c_{6}\cdot 96\cdot d_{V}\cdot L+\frac{8C_{\operatorname{Lip},2}\bar{c}}{c_{1}^{2}}\cdot 96\cdot d_{V}\cdot L\right)\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}},$
-which is attainable for $c_{1}=c_{1}(\alpha,\eta_{2},\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ large enough.
-Hence an applycation of Lemma A.7 yields
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{v_{l}^{2}\sigma}{96d_{V}}\right)$ $\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\right)$
-$\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
+Now in the setting of lemma 4.7, for arbitrary $\bar{c}\in\mathbb{R}_{\geq 1}$, such that $x:=\bar{c}2^{2l}NR^{2}\in\mathbb{R}_{\geq 1}$, we have for some universal constant $L\in\mathbb{R}_{>0}$
+$L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)$
+$\hskip 28.45274pt\leq L\left(c_{5}c_{1}^{-(1-\gamma_{1}\tau_{1})}v_{l}^{2}\sqrt{N}+v_{l}\sqrt{\bar{c}}\sqrt{N}2^{l}R+\frac{1}{\sqrt{N}}(c_{6}v_{l}^{2}N+8\bar{c}N2^{2l}R^{2}C_{\mathscr{G},\operatorname{B}})\right)$
+$\hskip 28.45274pt\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}}\left(96Ld_{V}\sigma^{-1}c_{5}c_{1}^{-(1-\gamma_{1}\tau_{1})}+96Ld_{V}\frac{\sqrt{\bar{c}}}{\sigma c_{1}}+96Ld_{V}c_{6}+768Ld_{V}C_{\mathscr{G},\operatorname{B}}\frac{\bar{c}}{c_{1}^{2}}\right)$
+$\hskip 28.45274pt\leq\sqrt{N}\frac{v_{l}^{2}\sigma}{96\cdot d_{V}},$
+where in the last step follows after choosing $c_{1}=c_{1}(\texttt{P})\in\mathbb{R}_{>0}$ sufficiently large. Hence an application of lemma 4.7 yields
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq\sqrt{N}\frac{v_{l}^{2}\sigma}{96d_{V}}\right)$
+$\hskip 42.67912pt\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}|Z_{N,j}^{(1)}(\theta)|\geq L\left(\mathcal{J}_{2}(\mathcal{H}_{l,j})+v_{l}\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H}_{l,j})+8C_{\mathscr{G},\operatorname{B}}x)\right)\right)$
+$\hskip 42.67912pt\leq 2\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
 As the last upper bound is independent of the choice of $j\in\left\llbracket d_{V}\right\rrbracket$, we have overall
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{v_{l}^{2}}{96}\right)\leq 2d_{V}\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{1}(\theta)|\right\}\geq\sqrt{N}\frac{v_{l}^{2}}{96}\right)\leq 2d_{V}\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
 It remains to study the process $T_{2}$. By writing
-$T_{2}(\theta):=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}\left\{\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\tilde{\theta})(Z_{i})\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z_{i})-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}\right\}=:\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i}),$
-we need to controll
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{\left|\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[h_{\theta}(Z_{i})])\right|\right\}\geq\sqrt{N}\frac{2^{2l}}{48}R^{2}c_{1}^{2}\sigma^{2}\right\}\right).$
+$T_{2}(\theta)=\frac{1}{2\sigma^{2}\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i})$
+with $h_{\theta}(\cdot)=|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)|_{V}^{2}$, we need to control
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{\left|\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}(h_{\theta}(Z_{i})-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[h_{\theta}(Z_{i})])\right|\right\}\geq\sqrt{N}\frac{2^{2l}}{48}R^{2}c_{1}^{2}\sigma^{2}\right).$
 Similiar as before, we define for each $l\in\mathbb{N}$ a class of measurable functions
-$\mathcal{H}_{l}:=\left\{h_{\theta}:=\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)\right|_{V}^{2}:\mathcal{Z}\to\mathbb{R}\,:\,\theta\in\Theta_{l}\right\},$
-and verify the assumptions of Lemma A.7. Indeed, note that for each $\theta\in\Theta_{l}$, we have
-$\sup_{z\in\mathcal{Z}}\left\{\left|h_{\theta}(z)\right|\right\}$ $=\sup_{z\in\mathcal{Z}}\left\{\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\tilde{\theta})(z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right|\right\}\leq 8C_{\mathscr{G},\operatorname{B}}$
+$\mathcal{H}_{l}:=\left\{h_{\theta}:=|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\tilde{\theta})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta)(\cdot)|_{V}^{2}:\mathcal{Z}\to\mathbb{R}\,:\,\theta\in\Theta_{l}\right\},$
+and verify the assumptions of lemma 4.7. Indeed, note that for each $\theta\in\Theta_{l}$, we have
+$\sup_{z\in\mathcal{Z}}\left\{\left|h_{\theta}(z)\right|\right\}$ $=\sup_{z\in\mathcal{Z}}\left\{\left||\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\tilde{\theta})(z)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(z)-\mathscr{G}(\theta)(z)|_{V}^{2}\right|\right\}\leq 8C_{\mathscr{G},\operatorname{B}}$
 due to C2) in 2.3. Moreover, for $Z\sim\zeta$ and all $\theta\in\Theta_{l}$, $l\in\mathbb{N}$, we have
-$\mathbb{E}^{Z}[h_{\theta}^{2}(Z)]$ $=\mathbb{E}^{Z}\left[\left|\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\tilde{\theta})(Z)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\theta)(Z)\right|_{V}^{2}\right|^{2}\right]\leq 16C_{\mathscr{G},\operatorname{B}}^{2}\left\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta)\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq 16C_{\mathscr{G},\operatorname{B}}^{2}v_{l}^{2},$
-where we have applied C2) from 2.3 and used the notation as before. In order to bound the entropy integrals, observe that for any $\theta_{1},\theta_{2}\in\Theta_{l}$
-$d_{2}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)},$
+$\mathbb{E}^{Z}[h_{\theta}^{2}(Z)]$ $=\mathbb{E}^{Z}\left[\left||\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\tilde{\theta})(Z)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(Z)-\mathscr{G}(\theta)(Z)|_{V}^{2}\right|^{2}\right]$
+$\leq 16C_{\mathscr{G},\operatorname{B}}^{2}\|\mathscr{G}(\tilde{\theta})-\mathscr{G}(\theta)\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq 16C_{\mathscr{G},\operatorname{B}}^{2}v_{l}^{2},$
+with $v_{l}$ defined in eq. 24, where we have applied C2) from 2.3. In order to bound the entropy integrals, observe that for any $\theta_{1},\theta_{2}\in\Theta_{l}$
+$d_{2}(\theta_{1},\theta_{2})$ $=\|h_{\theta_{1}}-h_{\theta_{2}}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}=\||\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)|_{V}^{2}\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}$
+$\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)},$
 as well as
-$d_{\infty}(\theta_{1},\theta_{2})=\left\|h_{\theta_{1}}-h_{\theta_{2}}\right\|_{\infty}=\left\|\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)\right|_{V}^{2}-\left|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)\right|_{V}^{2}\right\|_{\infty}$ $\leq 4C_{\mathscr{G},\operatorname{B}}\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\infty}.$
+$d_{\infty}(\theta_{1},\theta_{2})$ $=\|h_{\theta_{1}}-h_{\theta_{2}}\|_{\infty}=\||\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{2})(\cdot)|_{V}^{2}-|\mathscr{G}(\theta_{\mathfrak{o}})(\cdot)-\mathscr{G}(\theta_{1})(\cdot)|_{V}^{2}\|_{\infty}$
+$\leq 4C_{\mathscr{G},\operatorname{B}}\cdot\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\infty}.$
 Evidently, we obtain the same upper bounds for the metric entropies
-$\log N\left(\mathcal{H}_{l},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l},d_{\infty},\rho\right),$
-as already seen for the process $T_{1}$, by a change of constant.
-Thus, following the argumentation as before, we can choose $c_{1}:=(\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},d_{V},\bar{c},\sigma)\in\mathbb{R}_{>0}$ sufficiently large to obtain
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right\}\right)\leq 2\exp\left(-\bar{c}N2^{2l}R^{2}\right).$
+$\log N\left(\mathcal{H}_{l},d_{2},\rho\right)\text{ and }\log N\left(\mathcal{H}_{l},d_{\infty},\rho\right)$
+as already derived before (up to a change of constants). Thus, following the same argumentation as above, we can choose $c_{1}=c_{1}(\texttt{P},d_{V})\in\mathbb{R}_{>0}$ sufficiently large to obtain
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\sup_{\theta\in\Theta_{l}}\left\{|T_{2}(\theta)-\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[T_{2}(\theta)]|\right\}\geq\sqrt{N}\frac{2^{2l}}{96}R^{2}c_{1}^{2}\right)\leq 2\exp\left(-\bar{c}2^{2l}NR^{2}\right).$
 Overall, we have derived the existence of constants $c_{7}=c_{7}(d_{V})\in\mathbb{R}_{>0}$, such that
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{7}\cdot\sum_{l\in\mathbb{N}}\exp(-\bar{c}N2^{2l}R^{2})\leq c_{7}\exp(-\bar{c}NR^{2}),$
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}(\Xi_{\mathfrak{r},R})\leq c_{7}\cdot\sum_{l\in\mathbb{N}}\exp(-\bar{c}2^{2l}NR^{2})\leq c_{7}\exp(-\bar{c}NR^{2}),$
 if $c_{1}\in\mathbb{R}_{\geq 1}$ in the definition of $\Xi_{\mathfrak{r},R}$ is choosen sufficiently large, observing that the numerical values of the constants may changed again in the last line without any further dependencies.
@@
 C6 and fix some $\bar{c}\in\mathbb{R}_{\geq 1}$.
-We define the map $T:(\mathfrak{r}^{2},\infty)\to\mathbb{R}_{>0},\,t\mapsto c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)$ and set $\mathfrak{m}:=c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+\mathfrak{r}^{2})$, with $c_{1}$ from the first part of the theorem.
+We define the map $T:(\mathfrak{r}^{2},\infty)\to\mathbb{R}_{>0},\,t\mapsto c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)$ and set $\mathfrak{m}:=c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+\mathfrak{r}^{2})$, with $c_{1}\in\mathbb{R}_{>0}$ choosen sufficiently large according to i).
 Applying the transformation theorem, we obtain
 $\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot\int_{(\mathfrak{r}^{2},\infty)}\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+t)\right)\mathrm{d}\mathcal{L}(t).$
-Now observe that $t\in(\mathfrak{r}^{2},\infty)$ satisfies $\sqrt{t}\geq\mathfrak{r}$ and hence also Eq.
-C4 and Eq.
-C5 by an application of Eq.
-C6.
-Thus, by an application of the first part of the theorem
-$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot c(d_{V})\cdot\int_{(\mathfrak{r}^{2},\infty)}\exp(-\bar{c}Nt)\mathrm{d}\mathcal{L}(t)\leq\mathfrak{m}+\frac{c_{1}\cdot c(d_{V})}{\bar{c}N},$
-where we have chosen $c_{1}=c_{1}(\bar{c},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\sigma)$ sufficiently large.
-This shows the claim.
-Proof of 7 (Corollary 2.7).
+Now observe that $t\in(\mathfrak{r}^{2},\infty)$ satisfies $\sqrt{t}\geq\mathfrak{r}$ and hence also eq. C4 and eq. C5 by an application of eq. C6. Thus, utilizing the concentration inequality yields
+$\mathbb{E}_{\theta_{\mathfrak{o}}}^{N}[\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})]$ $\leq\mathfrak{m}+c_{1}\cdot c(d_{V})\cdot\int_{(\mathfrak{r}^{2},\infty)}\exp(-\bar{c}Nt)\mathrm{d}\mathcal{L}(t)\leq\mathfrak{m}+\frac{c_{1}\cdot c(d_{V})}{\bar{c}N}.$
+The claim then follows by taking the infimum over all admissible $\tilde{\theta}$. ∎
+Proof of corollary 2.7.
 To prove the claim, we aim to apply theorem 2.6.
@@
 Moreover, observe for any $c_{1}\in\mathbb{R}_{>0}$ that
-$c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+R^{2})=c_{1}\cdot(\mathfrak{r}^{2}\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}+\mathfrak{r}^{2})\leq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}.$
-Thus, for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\bar{c},\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ sufficiently large, that
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}\right)\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq 2(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
+$c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+R^{2})=c_{1}\cdot(\mathfrak{r}^{2}\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}+\mathfrak{r}^{2})\leq c_{1}\cdot(\operatorname{L}^{2}+1)\cdot\mathfrak{r}^{2}.$
+Thus, for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\texttt{P},L)\in\mathbb{R}_{\geq 1}$ sufficiently large, that
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\operatorname{L}^{2}+1)\cdot\mathfrak{r}^{2}\right)\\ \leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq 2(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$ (25)
 This shows the first part of the claim. Moreover, according to the second part of theorem 2.6, we have
-$\mathbb{E}_{\theta_{0}}^{N}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|^{2}_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right]$ $\leq c_{3}\cdot\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{3}\cdot\left((\operatorname{L}+1)\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{4}\cdot\mathfrak{r}_{N}^{2}$
+$\mathbb{E}_{\theta_{0}}^{N}\left[\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|^{2}_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right]\leq c_{3}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{3}\left((\operatorname{L}^{2}+1)\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{4}\mathfrak{r}_{N}^{2}$
 for some constant $c_{4}=c_{4}(\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{>0}$.
-This proves the claim.
-Proof of 8 (Corollary 2.10).
-In Corollary 2.7, we have seen that for each $\bar{c}\in\mathbb{R}_{\geq 1}$ we have
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot\mathfrak{r}^{2}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}),$
-for a constant $c_{1}=c_{1}(\bar{c},,\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ large enough, which is equivalent with
-$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}_{N}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|^{2}_{H^{\alpha}(\mathcal{O},W)}\leq c_{1}\mathfrak{r}_{N}^{2}\right)\geq 1-c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
-Thus, to prove the claim, it suffices to restrict ourselves on the measurable event
-$T_{N}:=\left\{\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\mathfrak{r}_{N}^{2},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}\right\}.$
-As $\theta_{\mathfrak{o}}\in H^{\alpha}(\mathcal{O},W)$ is fixed, there exists some constant $\operatorname{L}=\operatorname{L}(\theta_{\mathfrak{o}})\in\mathbb{R}_{>0}$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}$ and thus some $c_{2}=c_{2}(c_{1},\operatorname{L})\in\mathbb{R}_{>0}$, such that on the event $T_{N}$, we have
-$\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq c_{2}\cdot\mathfrak{r}_{N}^{2}\text{ as well as }\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{2}.$
-By an application of C7) in 2.9, we have on the event $T_{N}$
-$\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c(C_{\operatorname{M}},c_{2})\cdot\mathfrak{r}_{N}^{\tau},$
-which shows the claim.
-   Proofs of Section 3
+Taking the supremum over all admissible $\theta_{\mathfrak{o}}$ on both sides shows the claim.
+∎
+Proof of corollary 2.10.
+Firstly, for $c_{1}\in\mathbb{R}_{>0}$ define the event
+$A_{N}:=\left\{\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}_{N}^{2}\|\hat{\theta}_{\textsc{MAP}}\|^{2}_{H^{\alpha}(\mathcal{O},W)}\leq c_{1}\mathfrak{r}_{N}^{2}\right\}.$
+By an application of corollary 2.7, we can choose for each $\bar{c}\in\mathbb{R}_{\geq 1}$ the constant $c_{1}=c_{1}(\bar{c},,\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)})\in\mathbb{R}_{\geq 1}$ large enough to obatin
+$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(A_{N}\right)\geq 1-c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
+We observe that
+$A_{N}\subset B_{N}:=\left\{\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\mathfrak{r}_{N}^{2},\,\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}\right\}.$
+As $\theta_{\mathfrak{o}}\in H^{\alpha}(\mathcal{O},W)$ is fixed, there exists some constant $\operatorname{L}=\operatorname{L}(\theta_{\mathfrak{o}})\in\mathbb{R}_{>0}$ with $\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}$ and thus on the event $B_{N}$, we have
+$\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\cdot\mathfrak{r}_{N}^{2}\text{ as well as }\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}+\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}+L.$
+By applying the inverse continuity modulus formulated in 2.9, we have on the event $B_{N}$
+$\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c(C_{\operatorname{M}},c_{1})\cdot\mathfrak{r}_{N}^{\tau}.$
+With the estimations above, the claim follows.
+4.2 Proofs of section 3
 […]
-   Darcy Problem - Proofs
-Let $f\in C^{1}(\bar{\mathcal{O}})$ with $f\geq f_{\min}\in\mathbb{R}_{>0}$. We define the differential operator
 $L^{D}_{f}:H_{0}^{2}(\mathcal{O})\to\mathbb{L}^{2}(\mathcal{O})\text{, }u\mapsto L_{f}^{D}[u]:=\operatorname{div}(f\nabla u),$
 where $\nabla$ denotes the usual gradient and $\operatorname{div}$ the divergence operator.
-Refering on Evans [2010], it is well known from elliptic PDE theory, that $L^{D}$ is a linear and bounded operator, which posses a linear inverse, denoted by
+From classical elliptic PDE theory (see for instance Evans (2010) Theorem 4 and the Remark after it in Chapter 6.3), it is well-known, that $L^{D}$ is a linear and bounded operator, which possess a linear inverse, we denote by
 $V_{f}^{D}:\mathbb{L}^{2}(\mathcal{O})\to H^{2}_{0}(\mathcal{O})\text{, }h\mapsto V_{f}^{D}[h].$
@@
 $\begin{dcases}L_{f}[w_{f,h}]&=h\text{ on $\mathcal{O}$,}\\ \hfill w_{f,h}&=0\text{ on $\partial\mathcal{O}$.}\end{dcases}$
-Utilizing the Feynman-Kac formula, Lemma 20 in Nickl et al. [2020] demonstrates that
-$\left\|V^{D}_{f}[h]\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|h\right\|_{\infty},$ (A.8)
-for all $h\in C^{\eta}(\mathcal{O})$, $\eta\in\mathbb{R}_{>0}$, if additionally $f\in C^{2}(\mathcal{O})$.
+Utilizing the corresponding Feynman-Kac formula, Lemma 20 in Nickl et al. (2020) demonstrates that
+$\|V^{D}_{f}[h]\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\|h\|_{\infty}$ (26)
+holds for all $h\in C^{\eta}(\mathcal{O})$, $\eta\in\mathbb{R}_{>0}$, if additionally $f\in C^{2}(\mathcal{O})$.
 Keeping the notation of section 3 in mind, we now verify that the forward map $\mathscr{G}^{D}$ satisfies the conditions C1) - C3) formulated in 2.3.
-As already suggested in the previous section, C3) is the cruicial part for the Darcy problem, as it fails to satisfy the assumption with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
+The cruicial part is to verify condition C3) for some $\eta_{1}$, $\eta_{2}$ and $\gamma_{2}$.
   1. C1)
@@
 The forward map $\mathscr{G}^{D}:H^{\alpha}_{c}(\mathcal{O})\to\mathbb{L}^{2}(\mathcal{O})$ is $(\alpha,4,1)$-regular, that is for all $\theta_{1},\theta_{2}\in H^{\alpha}_{c}(\mathcal{O})$, we have
-    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq C_{\operatorname{Lip},2}\cdot(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{4})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{1}(\mathcal{O}))^{*}}$
-    with some universal constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$. The proof can be found in Nickl et al. [2020].
-  2. C1)
-    Uniform boundedness: As described in Section 3, we identify with each $\theta\in H^{\alpha}_{c}(\mathcal{O})$ some $f=f_{\theta}\in\mathcal{F}$, such that $\mathscr{G}^{D}(\theta)=G^{D}(f_{\theta})=V_{f_{\theta}}^{D}[g]$.
+    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}\leq C_{\operatorname{Lip},2}\cdot(1+\|\theta_{1}\|_{H^{\alpha}(\mathcal{O})}^{4}\lor\|\theta_{2}\|_{H^{\alpha}(\mathcal{O})}^{4})\cdot\|\theta_{1}-\theta_{2}\|_{(H^{1}(\mathcal{O}))^{*}}$
+    with some universal constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$. The proof can be found in Nickl et al. (2020).
+  2. C2)
+    Uniform boundedness: As described in section 3, we identify with each $\theta\in H^{\alpha}_{c}(\mathcal{O})$ some $f=f_{\theta}=\Psi(\theta)\in\mathcal{F}$, such that $\mathscr{G}^{D}(\theta)=G^{D}(f_{\theta})=V_{f_{\theta}}^{D}[g]$.
 Thus, with eq.
-A.8, we obtain
-    $\sup_{\theta\in H^{\alpha}_{c}(\mathcal{O}),x\in\mathcal{O}}\left\{|\mathscr{G}^{D}(\theta)(x)|\right\}$ $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|G^{D}(f)(x)|\right\}=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|V_{f}^{D}[g](x)|\right\}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|g\right\|_{\infty}.$
-  3. C1)
+26, we obtain
+    $\sup_{\theta\in H^{\alpha}_{c}(\mathcal{O}),x\in\mathcal{O}}\left\{|\mathscr{G}^{D}(\theta)(x)|\right\}$ $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|G^{D}(f)(x)|\right\}$
+    $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|V_{f}^{D}[g](x)|\right\}\leq c(d,\mathcal{O},f_{\min})\cdot\|g\|_{\infty}.$
+  3. C3)
     $\mathbb{L}^{\infty}$-Lipschitz continuity: Let $\alpha>2+\frac{d}{2}$.
@@
     where the right-hand side defines a function in $C^{\nu}(\mathcal{O})$ for some small $\nu\in\mathbb{R}_{>0}$.
-In other words, applying the linear inverse $V_{f_{\theta_{1}}}^{D}$ on both sides of the equation, we obtain
-    $u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}=V_{f}^{D}[\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})].$
-    Thus, by an applycation of Eq. A.8
-    $\left\|u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}.$
-    Moreover, by an applycation of the product rule
-    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}$ $\leq\left\|\nabla(f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}\right\|_{\infty}+\left\|(f_{\theta_{2}}-f_{\theta_{1}})\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$
-    $\leq\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}\cdot\left(\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}+\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\right).$
+Thus, applying the linear inverse $V_{f_{\theta_{1}}}^{D}$ on both sides of the equation, we obtain
+    $u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}=V_{f}^{D}[\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})],$
+    and hence, by an application of eq. 26
+    $\|u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}.$
+    Utilizing the product rule, the right-hand side can be further upper bounded by
+    $\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}$ $\leq\|\nabla(f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}\|_{\infty}+\|(f_{\theta_{2}}-f_{\theta_{1}})\Delta u_{f_{\theta_{2}}}\|_{\infty}$
+    $\leq\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}\cdot\left(\|u_{f_{\theta_{2}}}\|_{C^{1}(\mathcal{O})}+\|\Delta u_{f_{\theta_{2}}}\|_{\infty}\right).$
     We study the two remaining terms in the last display seperately, starting with the second one. We have
-    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$ $=\left\|f_{\theta_{2}}^{-1}\cdot(g-\nabla u_{f_{\theta_{2}}}\cdot\nabla f_{\theta_{2}})\right\|_{\infty}\leq\left\|f_{\theta_{2}}^{-1}\right\|_{\infty}\left(\left\|g\right\|_{\infty}+\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}\right).$
+    $\|\Delta u_{f_{\theta_{2}}}\|_{\infty}$ $=\|f_{\theta_{2}}^{-1}\cdot(g-\nabla u_{f_{\theta_{2}}}\cdot\nabla f_{\theta_{2}})\|_{\infty}$
+    $\leq\|f_{\theta_{2}}^{-1}\|_{\infty}\cdot\left(\|g\|_{\infty}+\|u_{f_{\theta_{2}}}\|_{C^{1}(\mathcal{O})}\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}\right)$
 […]
@@
 […]
-[…]
-    Thus, elementary computation show that
-    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3}).$
-    Hence, by an further application of Lemma 22 in Nickl et al. [2020], we have
-    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
-    Overall, we have
-    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}^{3}\lor\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
-    By an application of Lemma 29 in Nickl et al. [2020] and Lemma A.10, we obtain finally
-    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})}^{4})\cdot\left\|\theta_{2}-\theta_{1}\right\|_{C^{1}(\mathcal{O})}.$
-    Thus, the forward map $\mathscr{G}^{D}$ satisfies C3) in 2.3 with $\gamma_{2}=4$, and $\eta_{1}=\eta_{2}=1$.
+    and hence
+    $\|\Delta u_{f_{\theta_{2}}}\|_{\infty}\leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3}).$
+    Another application of Lemma 22 in Nickl et al. (2020) yields
+    $\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}\leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3})\cdot\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})},$
+    which leads overall to
+    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\infty}\\ \leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}^{3}\lor\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3})\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}.$
+    Now utilizing the properties of the regular link function $\Psi$ formulated in lemma 4.10 and Lemma 29 in Nickl et al. (2020), it follows
+    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\infty}\\ \leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|\theta_{1}\|_{C^{1}(\mathcal{O})}^{4}\lor\|\theta_{2}\|_{C^{1}(\mathcal{O})}^{4})\cdot\|\theta_{2}-\theta_{1}\|_{C^{1}(\mathcal{O})}.$
+    Hence, the forward map $\mathscr{G}^{D}$ satisfies C3) in 2.3 with $\gamma_{2}=4$, and $\eta_{1}=\eta_{2}=1$.
 After prooving the conditions C1) - C3) for the forward map $\mathscr{G}^{D}$ formulated in 2.3, we are showing an analogous inverse continuity modulus for the solution operator $G^{D}$ of the Darcy Problem as assumed in C7) in 2.9.
 The proof is based on a interpolation inequality for Sobolev spaces, which reads for all $\alpha\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$
-$\left\|u\right\|_{H^{\beta}(\mathcal{O})}\leq c(\alpha,\beta,\mathcal{O})\cdot\left\|u\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|u\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$ (A.9)
+$\|u\|_{H^{\beta}(\mathcal{O})}\leq c(\alpha,\beta,\mathcal{O})\cdot\|u\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\|u\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}},$ (27)
+which is proven for instance in Theorem 1 in Section 4.3.1 in Triebel (1978) observing that $H^{\beta}(\mathcal{O})=B_{22}^{\beta}(\mathcal{O})$.
   1. C’7)
     Inverse Continuity Modulus: Let $f_{1},f_{2}\in\mathcal{F}$ for some $\alpha>2+\frac{d}{2}$.
-Then there exists a constant $\operatorname{B}\in\mathbb{R}_{>0}$, such that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq\operatorname{B}$.
+Assume there exists a constant $\operatorname{B}\in\mathbb{R}_{>0}$, such that $\|f_{1}\|_{C^{1}(\mathcal{O})}\lor\|G^{D}(f_{1})\|_{C^{2}(\mathcal{O})}\leq\operatorname{B}$ and additionally that the heat source function $g$ satisfies $\inf_{x\in\mathcal{O}}g(x)\geq g_{\min}$.
 Lemma 24 in Nickl et al.
-[2020] provides the existence of a finite constant $c_{1}=c_{1}(g_{\min},f_{\min},\operatorname{B},\mathcal{O},d)\in\mathbb{R}_{>0}$, such that
-    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{1}\cdot\left\|f_{2}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}.$
-    By an application of Eq. A.9, we have further
-    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}\leq c_{2}(\alpha,\mathcal{O})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
-    Moreover, by an application of Lemma 23 in Nickl et al. [2020], we have
-    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\leq c_{3}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})$
-    with a finite constant $c_{3}=c_{3}(\alpha,d,\mathcal{O},f_{\min},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. Thus, we have
-    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{4}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}$
-    for a constant $c_{4}=c_{4}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{B},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$.
-Now, let $\operatorname{M}\in\mathbb{R}_{>0}$ be arbitrary and assume that $\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}+\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}\leq\operatorname{M}$, such that $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq\delta$ for $\delta\in\mathbb{R}_{>0}$ small enough.
-By observing that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq c(\operatorname{M},\alpha,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})$, we obtain
-    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{5}\cdot\delta^{\frac{\alpha-1}{\alpha+1}}$
-    with $c_{5}=c_{5}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{M},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. This shows the claim with $\tau=\frac{\alpha-1}{\alpha+1}\in(0,1)$.
-Proof of 9 (Theorem 3.2).
+(2020) then provides the existence of a finite constant $c_{1}=c_{1}(g_{\min},f_{\min},\operatorname{B},\mathcal{O},d)\in\mathbb{R}_{>0}$, such that
+    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{1}\cdot\|f_{2}\|_{C^{1}(\mathcal{O})}\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{2}(\mathcal{O})}.$
+    By an application of eq. 27 with $\beta=2$, we further have
+    $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{2}(\mathcal{O})}\\ \leq c_{2}(\alpha,\mathcal{O})\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
+    Moreover, by an application of Lemma 23 in Nickl et al. (2020), we have
+    $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\leq c_{3}\cdot(1+\|f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\|f_{2}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})$
+    with a finite constant $c_{3}=c_{3}(\alpha,d,\mathcal{O},f_{\min},\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$. Thus, we have
+    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{4}\cdot(1+\|f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\|f_{2}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}$
+    for a constant $c_{4}=c_{4}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{B},\|g\|_{\alpha-1})\in\mathbb{R}_{>0}$.
+    Now, let $\operatorname{M}\in\mathbb{R}_{>0}$ be arbitrary and let $f_{1}$ and $f_{2}$ be chosen in such a way that $\|f_{1}\|_{H^{\alpha}(\mathcal{O})}+\|f_{2}\|_{H^{\alpha}(\mathcal{O})}\leq\operatorname{M}$ and $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}\leq\delta$ for $\delta\in\mathbb{R}_{>0}$ small enough.
+By an application of Lemma 23 in Nickl et al.
+(2020) and the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{1}(\mathcal{O})$, we have $\|f_{1}\|_{C^{1}(\mathcal{O})}\lor\|G^{D}(f_{1})\|_{C^{2}(\mathcal{O})}\leq c(\alpha,d,\mathcal{O},f_{\min},\|g\|_{H^{\alpha-1}(\mathcal{O})})$.
+By above computations we derive
+    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{5}\cdot\delta^{\frac{\alpha-1}{\alpha+1}}$
+    with $c_{5}=c_{5}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{M},\|g\|_{\alpha-1})\in\mathbb{R}_{>0}$.
+Taking the supremum over all admissible $f_{1}$ and $f_{2}$ as in the spirit of 2.9 shows the claim with $\tau=\frac{\alpha-1}{\alpha+1}\in(0,1)$.
+We are now providing the proof of theorem 3.2.
+Proof of theorem 3.2.
 Firstly, note that the Darcy problem is formulated as special case of section 2 with $V=W=\mathbb{R}$.
@@
     The existence of a maximizer $\hat{f}_{\textsc{MAP}}$ follows directly by the equivalent formulation presented in section 3 and an application of theorem 2.6 under considerations that the forward map $\mathscr{G}^{D}$ satisfies all assumption of theorem 2.6.
-  2. i)
+  2. ii)
     As the forward map $\mathscr{G}^{D}$ satisfies C1) - C3) in 2.3, we apply theorem 2.6 directly with $\tilde{\theta}=\theta_{\mathfrak{o}}$ by observing that
     $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})=\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}}).$
-    Thus, for every $\bar{c}\in\mathbb{R}_{>0}$ there exists a constant $c_{1}=c_{1}(\alpha,f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O},\sigma)\in\mathbb{R}_{>0}$ sufficiently large, such that
-    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\right)\leq c_{2}\exp\leq\left(-\bar{c}m^{2}N\right)$
+    Thus, for every $\bar{c}\in\mathbb{R}_{>0}$ there exists a constant $c_{1}=c_{1}(\alpha,f_{\min},\|g\|_{\infty},d,\mathcal{O},\sigma)\in\mathbb{R}_{>0}$ sufficiently large, such that
+    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq c_{1}(\mathfrak{r}_{N}^{2}\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\right)\leq c_{2}\exp\leq\left(-\bar{c}m^{2}N\right)$
     for all $m\geq\mathfrak{r}_{N}$ satisfying eq.
@@
 C5.
-We identify $f_{\mathfrak{o}}\in\mathcal{F}$ with some $\theta_{\mathfrak{o}}=\Psi^{-1}\circ f_{\mathfrak{o}}\in H^{\alpha}_{c}(\mathcal{O})$ and conclude $\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{3}(\Psi,\operatorname{L})\in\mathbb{R}_{>0}$ uniformly over all $f_{\mathfrak{o}}\in\mathcal{F}(\operatorname{L})$.
-Choosing $\operatorname{M}\geq 1\lor c_{3}$ and $m:=\operatorname{M}\mathfrak{r}_{N}\geq\mathfrak{r}_{N}$, such that $c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\leq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}$.
-Thus,
-    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}\right)\leq c_{2}\exp\leq\left(-\bar{c}M^{2}N\mathfrak{r}_{N}^{2}\right).$
+We identify $f_{\mathfrak{o}}\in\mathcal{F}$ with some $\theta_{\mathfrak{o}}=\Psi^{-1}\circ f_{\mathfrak{o}}\in H^{\alpha}_{c}(\mathcal{O})$ and conclude $\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}\leq c_{3}(\Psi,\operatorname{L})\in\mathbb{R}_{>0}$ uniformly over all $f_{\mathfrak{o}}\in\mathcal{F}(\operatorname{L})$.
+Choose $\operatorname{M}\geq 1\lor c_{3}$ and $m:=\operatorname{M}\mathfrak{r}_{N}\geq\mathfrak{r}_{N}$, such that $c_{1}(\mathfrak{r}_{N}^{2}\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\leq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}$.
+Then,
+    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}\right)\leq c_{2}\exp\leq\left(-\bar{c}M^{2}N\mathfrak{r}_{N}^{2}\right).$
     Updating constants yields the claim.
+  3. iii)
+    As we have shown an analogoue inverse continuity modulus C’7), the claim follows directly following along the lines of the proof of corollary 2.10 and is hence omitted.
+We are now providing the proof of theorem 3.3.
+Proof of theorem 3.3.
+The proof follows the lines of the proofs of analogous results in the White Noise Model setting derived in Nickl et al.
+(2020) (see Proof of Theorem 9).
+Hence, in this work we only provide a sketch of the proof and ommit the details, which are left for the interested reader.
   1. i)
-    As we have shown an analogoue inverse continuity modulus C’7), the claim follows directly following along the lines of the proof of Corollary 2.10 and is hence omitted.
-Proof of 10 (Theorem 3.3).
-Observing that we have shown the analogue results to the random design regression model, the proof of this theorem follows exactly along the lines of corresponding proofs in Nickl et al.
-[2020].
-Hence, we are here only providing a scetch of the proof and ommit the details, which are let for the interested reader.
-  1. i)
-    By an applycatio of Eq. A.9, we obtain
-    $\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\leq c_{1}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$
-    While the second factor can be upper bounded by regularity estimates for $G^{D}$, the first factor is already indicating the correct exponent.
-Indeed, the first factor can be upper bounded by $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$.
+    By an application of eq. 27, we obtain
+    $\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\\ \leq c_{1}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2\beta}{\alpha+1}}.$
+    The first factor can be directly upper bounded by $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$.
+The second factor can be further upper bounded by firstly applying Lemma 23 in Nickl et al.
+(2020) and then by considering the definition of $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$.
 Decomposing the underlying probability space into slices
-    $A_{0}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})<\delta_{N}\right\}\text{ and }A_{j}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\in(2^{j-1}\delta_{N},2^{j}\delta_{N}]\right\}$
-    for some $\delta_{N}\approx\mathfrak{r}_{N}$, the corresponding expectation can be upper bounded by
-    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\sum_{j\in\mathbb{N}_{0}}\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\mathds{1}_{A_{j}}\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\frac{\alpha+1-\beta}{\alpha+1}}\left(1+\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\alpha^{2}\beta}\right)\right].$
-    Using the upper bound of $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$ on each event $A_{j}$, as well as the concentration inequality of Theorem 3.2, we obtain
-    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
+    $A_{0}:=\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})<\delta_{N}^{2}\right\}\text{ and }A_{j}:=\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\in(2^{2(j-1)}\delta_{N}^{2},2^{2j}\delta_{N}^{2}]\right\}$
+    for some $\delta_{N}\propto\mathfrak{r}_{N}$, the corresponding expectation can be upper bounded by
+    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\right]\\ \lesssim\sum_{j\in\mathbb{N}_{0}}\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\mathds{1}_{A_{j}}\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\frac{\alpha+1-\beta}{\alpha+1}}\left(1+(\mathfrak{r}_{N}^{-2}\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}}))^{\alpha^{2}\beta}\right)\right]\\ \lesssim\delta_{N}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\bigg(1+\sum_{j\in\mathbb{N}_{0}}2^{2j\frac{\alpha+1-\beta}{\alpha+1}}(1+2^{2j\alpha^{2}\beta})\mathbb{P}_{f_{\mathfrak{o}}}^{N}(A_{j})\bigg).$
+    Using the concentration inequality of theorem 3.2, we obtain
+    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\right]\lesssim\delta_{N}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
     which shows the claim.
-  2. i)
-    Similiarly, as in the proof of C’7) before, we can upper bound
-    $\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\lesssim\left\|\hat{f}_{\textsc{MAP}}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}.$
-    Thus, by an application of Eq. A.9 with $\beta=2$, we obtain
-    $\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}\lesssim\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{1+\alpha}(\mathcal{O})}^{\frac{2}{1+\alpha}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
-    The proof now follows the same idea as for the prediction error, to obtain an upper bound of the form
-    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha-1}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
+  2. ii)
+    Similiarly, as in the proof of C’7) before, we upper bound
+    $\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\lesssim\|\hat{f}_{\textsc{MAP}}\|_{C^{1}(\mathcal{O})}^{2}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{2}(\mathcal{O})}^{2}.$
+    Thus, by an application of eq. 27 with $\beta=2$, we obtain
+    $\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\|_{H^{2}(\mathcal{O})}^{2}\\ \lesssim\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\|_{H^{1+\alpha}(\mathcal{O})}^{\frac{4}{1+\alpha}}\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{2(\alpha-1)}{\alpha+1}}.$
+    Using the same techniques as for the prediction error, we obtain the bound
+    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\right]\lesssim\delta_{N}^{\frac{2(\alpha-1)}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
     which proves the claim.
-   Auxiliary results
-The following lemma was first proven in Nickl and Wang [2024] (see Lemma 3.12) and is a key tool in the proof of Theorem 2.6.
-Lemma § A.7 (Chaining for an empirical process).
+4.3 Auxiliary results
+The following lemma was proven in Nickl and Wang (2024) (see Lemma 3.12) and is a key tool in the proof of theorem 2.6.
+Lemma § 4.7 (Chaining for an Empirical Process).
 Let $\Theta$ be a countable set. Let $(\Omega,\mathcal{A},\mathbb{P}^{Z})$ be a probability space. Suppose a class of real-valued measurable functions
-:=$\mathcal{H}:=\left\{h_{\theta}:\Omega\to\mathbb{R}:\,\theta\in\Theta\right\}.$
+$\mathcal{H}:=\left\{h_{\theta}:\Omega\to\mathbb{R}:\,\theta\in\Theta\right\}.$
 Assume that $\mathcal{H}$ is uniformly bounded and has a variance envelope, namely
@@
     $\sup_{\theta\in\Theta}\left\{\sup_{\omega\in\Omega}\left\{|h_{\theta}(\omega)|\right\}\right\}\leq C_{\operatorname{B}},$
-  2. i)
+  2. ii)
     there exists $\sigma^{2}\in\mathbb{R}_{>0}$, such that
@@
 For i.i.d.
-copies $Z_{1},\dots,Z_{N}$ of $X\sim\mathbb{P}^{Z}$ and $\varepsilon_{1},\dots\varepsilon_{N}$ of $\varepsilon\sim\operatorname{N}(0,1)$ independent of $Z$, consider empirical processes either of the form
+copies $Z_{1},\dots,Z_{N}$ of $Z\sim\mathbb{P}^{Z}$ and $\varepsilon_{1},\dots\varepsilon_{N}$ of $\varepsilon\sim\operatorname{N}(0,1)$ independent of $Z$, consider empirical processes either of the form
 $T_{N}(\theta):=\frac{1}{\sqrt{N}}\sum_{i\in\left\llbracket N\right\rrbracket}h_{\theta}(Z_{i})\varepsilon_{i},\,\theta\in\Theta$
@@
 $\mathbb{P}\left(\sup_{\theta\in\Theta}\left\{|T_{N}(\theta)|\right\}\geq L\left(\mathcal{J}_{2}(\mathcal{H})+\sigma\sqrt{x}+N^{-\frac{1}{2}}(\mathcal{J}_{\infty}(\mathcal{H})+C_{\operatorname{B}}x)\right)\right)\leq 2\exp(-x).$
-Similarly to Nickl and Wang [2024], whenever we apply Lemma A.7 in this work with an uncountable set $\Theta$, one can verify that the supremum can be realized as one over a countable subset of it.
-Lemma § A.8 (Covering Numbers of Product Spaces).
-Let $(\tilde{S},\left\|\cdot\right\|_{\tilde{S}})$ be a normed space.
-Assume there exists a function $\psi:\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to\mathbb{R}$, such that for any $r,\rho\in\mathbb{R}_{>0}$ the covering number $N(\mathbb{B}_{\tilde{S}}(0,r),\left\|\cdot\right\|_{\tilde{S}},\rho)=\psi(r,\rho)$ is finite.
-For $K\in\mathbb{N}$ define $S:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\tilde{S}$ equipped with the norm $\left\|s\right\|_{S}:=\left(\sum_{i\in\left\llbracket K\right\rrbracket}\left\|s_{i}\right\|_{\tilde{S}}^{2}\right)^{\frac{1}{2}}$ for any $s=(s_{1},\dots,s_{K})\in S$.
+Similarly to Nickl and Wang (2024), whenever we apply lemma 4.7 in this work with an uncountable set $\Theta$, one can verify that the supremum can be realized as one over a countable subset of it.
+Lemma § 4.8 (Covering Numbers of Product Spaces).
+Let (S~,S~)($\tilde$$\|\cdot\|_{\tilde{S}}$) be a normed space.
+Assume there exists a function $\psi:\mathbb{R}_{>0}\times\mathbb{R}_{>0}\to\mathbb{R}$, such that for any $r,\rho\in\mathbb{R}_{>0}$ the covering number $N(\mathbb{B}_{\tilde{S}}(0,r),\|\cdot\|_{\tilde{S}},\rho)=\psi(r,\rho)$ is finite.
+For $K\in\mathbb{N}$ define $S:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\tilde{S}$ equipped with the norm $\|s\|_{S}:=\left(\sum_{i\in\left\llbracket K\right\rrbracket}\|s_{i}\|_{\tilde{S}}^{2}\right)^{\frac{1}{2}}$ for any $s=(s_{1},\dots,s_{K})\in S$.
 We then have
-$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\rho)\leq\psi\left(r,\frac{\rho}{\sqrt{K}}\right)^{K}.$
-Proof of 11 (Lemma A.8).
+$N(\mathbb{B}_{S}(0,r),\|\cdot\|_{S},\rho)\leq\psi\left(r,\frac{\rho}{\sqrt{K}}\right)^{K}.$
+Proof of lemma 4.8.
 By hypothesis, for all $\rho,r\in\mathbb{R}_{>0}$ there exists a covering $\mathcal{C}_{\tilde{S}}:=\{\tilde{s}^{(i)}\}_{i\in\left\llbracket N\right\rrbracket}\subseteq\tilde{S}$ of $\mathbb{B}_{\tilde{S}}(0,r)$ with $N:=\psi(r,\rho)$, i.e.
-for all $\tilde{s}\in\mathbb{B}_{\tilde{S}}(0,r)$ there exists $i_{\circ}=i_{\circ}(\tilde{s})\in\left\llbracket N\right\rrbracket$ with
-$\left\|\tilde{s}-\tilde{s}^{(i_{\circ})}\right\|_{\tilde{S}}\leq\rho.$
+for all $\tilde{s}\in\mathbb{B}_{\tilde{S}}(0,r)$ there exists $i_{\circ}=i_{\circ}(\tilde{s})\in\left\llbracket N\right\rrbracket$ with $\|\tilde{s}-\tilde{s}^{(i_{\circ})}\|_{\tilde{S}}\leq\rho$.
 Now let $s=(s_{1},\dots,s_{K})\in\mathbb{B}_{S}(0,r)$ be arbitrary chosen, such that
-$\left\|s\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}\right\|_{\tilde{S}}^{2}\leq r^{2}.$
-Thus, for all $j\in\left\llbracket K\right\rrbracket$, we have $s_{j}\in\mathbb{B}_{\tilde{S}}(0,r)$.
+$\|s\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\|s_{j}\|_{\tilde{S}}^{2}\leq r^{2},$
+and thus, we have $s_{j}\in\mathbb{B}_{\tilde{S}}(0,r)$ for all $j\in\left\llbracket K\right\rrbracket$.
 Hence, for each $s_{j}$, $j\in\left\llbracket K\right\rrbracket$, there exists $i=i_{j}\in\left\llbracket N\right\rrbracket$, such that $s_{j}\in\mathbb{B}_{\tilde{S}}(\tilde{s}^{(i_{j})},\rho)$.
 Defining $s^{(i_{1},\dots,i_{K})}:=(\tilde{s}^{(i_{1})},\dots,\tilde{s}^{(i_{K})})\in S$, we have
-$\left\|s-s^{(i_{1},\dots,i_{K})}\right\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\left\|s_{j}-\tilde{s}^{(i_{j})}\right\|_{\tilde{S}}^{2}<K\rho^{2}.$
+$\|s-s^{(i_{1},\dots,i_{K})}\|_{S}^{2}=\sum_{j\in\left\llbracket K\right\rrbracket}\|s_{j}-\tilde{s}^{(i_{j})}\|_{\tilde{S}}^{2}<K\rho^{2}.$
 Hence, $\mathcal{C}_{S}:=\bigtimes_{i\in\left\llbracket K\right\rrbracket}\mathcal{C}_{\tilde{S}}$ forms a $\sqrt{K}\varepsilon$-covering of $\mathbb{B}_{S}(0,r)$ and thus
-$N(\mathbb{B}_{S}(0,r),\left\|\cdot\right\|_{S},\sqrt{K}\rho)\leq N^{K}=\psi(r,\rho)^{K}.$
-This shows the claim.
-In the following lemma we use Lemma A.8 to derive an upper bound for the covering number of Sobolev balls.
-Lemma § A.9 (Covering Numbers of Sobolev Spaces).
+$N(\mathbb{B}_{S}(0,r),\|\cdot\|_{S},\sqrt{K}\rho)\leq N^{K}=\psi(r,\rho)^{K}.$
+This shows the claim. ∎
+In the following lemma we use lemma 4.8 to derive an upper bound for the covering number of Sobolev balls.
+Lemma § 4.9 (Covering Numbers of Sobolev Spaces).
 Let $d\in\mathbb{N}$.
@@
 Then there exists a constant $c=c(\alpha,\kappa,d)\in\mathbb{R}_{>0}$, such that
-    $N(\mathcal{F},\left\|\cdot\right\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha+\kappa}}\right)$
+    $N(\mathcal{F},\|\cdot\|_{(H^{\kappa}(\mathcal{O},W))^{*}},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha+\kappa}}\right)$
     for any $\rho\in\mathbb{R}_{>0}$.
-  2. i)
+  2. ii)
     Let $\eta\in\mathbb{R}_{\geq 0}$ , $r\in\mathbb{R}_{>0}$.
@@
 Then there exists a constant $c=c(\alpha,\eta,d)\in\mathbb{R}_{>0}$, such that
-    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O},W)},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha-\eta}}\right)$
+    $N(\mathcal{F},\|\cdot\|_{C^{\eta}(\mathcal{O},W)},\rho)\leq\exp\left(cd_{W}\left(r\sqrt{d_{W}}\rho^{-1}\right)^{\frac{d}{\alpha-\eta}}\right)$
     for any $\rho\in\mathbb{R}_{>0}$.
-Proof of 12 (Lemma A.9).
+Proof of lemma 4.9.
   1. i)
-    Restricting ourselfes to the case $d_{W}=1$, the claim is already shown in Nickl et al. [2020]. For general $d_{W}\in\mathbb{N}$, the claim follows by an application of Lemma A.8.
-  2. i)
+    Restricting ourselfes to the case $d_{W}=1$, the claim is already shown in the proof of Lemma 19 in Nickl et al.
+(2020), where a differentiation between the cases $\kappa\geq\frac{1}{2}$ and $\kappa<\frac{1}{2}$ is necessary to find proper $(H^{\kappa}(\mathcal{O}))^{*}$-coverings of $H^{\alpha}_{c}(\mathcal{O})$ and $H^{\alpha}(\mathcal{O})$, respectively.
+For general $d_{W}\in\mathbb{N}$, the claim follows by an application of lemma 4.8.
+  2. ii)
     We restrict ourselves to the case $d_{W}=1$.
-The general claim follows then by Lemma A.8.
-First, let $\eta=0$, such that $\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})}=\left\|\cdot\right\|_{\infty}$.
-If $\mathcal{O}$ corresponds to a one-dimensional interval, in chapter 4 of Giné and Nickl [2016], a covering result is shown.
-Using usual extension arguments, one obtains
-    $N(\mathcal{F},\left\|\cdot\right\|_{\infty},\rho)\leq\exp\left(c(\alpha,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha}}\right),$
-    which is also used in Nickl [2023].
+The general claim follows then by lemma 4.8.
+First, let $\eta=0$, such that $\|\cdot\|_{C^{\eta}(\mathcal{O})}=\|\cdot\|_{\infty}$.
+For $d=1$, if $\mathcal{O}$ equals $[0,1]$, in Giné and Nickl (2016) a covering result is shown, see Corollary 4.3.38 for instance.
+Using usual extension arguments, one obtains for the general case
+    $N(\mathcal{F},\|\cdot\|_{\infty},\rho)\leq\exp\left(c(\alpha,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha}}\right),$
+    which is also used in Nickl (2023).
 Now, let $\eta\in\mathbb{R}_{>0}\setminus\mathbb{N}$.
-Following Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$ is compact, and by Edmunds and Triebel [1996], we further have
-    $N(\mathcal{F},\left\|\cdot\right\|_{C^{\eta}(\mathcal{O})},\rho)\leq\exp\left(c\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-\eta}}\right).$
+Following remark 4.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{\eta}(\mathcal{O})=B_{\infty\infty}^{\eta}(\mathcal{O})$ is compact, and by Edmunds and Triebel (1996) (Section 3.3) and Triebel (1978) (Section 4.10.3), respectively, we further have
+    $N(\mathcal{F},\|\cdot\|_{C^{\eta}(\mathcal{O})},\rho)\leq\exp\left(c\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-\eta}}\right).$
     Now let $\eta=k\in\mathbb{N}$.
-According to Remark A.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is compact, as $\alpha>k+\frac{d}{2}$.
+According to remark 4.2, the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow B_{\infty 1}^{k}(\mathcal{O})\hookrightarrow C^{k}(\mathcal{O})$ is compact, as $\alpha>k+\frac{d}{2}$.
 Thus, for any $\theta_{1},\theta_{2}\in H^{\alpha}(\mathcal{O})$, we have
-    $\left\|\theta_{1}-\theta_{2}\right\|_{C^{\eta}(\mathcal{O})}\leq c_{\texttt{em}}\cdot\left\|\theta_{1}-\theta_{2}\right\|_{B_{\infty 1}^{k}(\mathcal{O})}.$
-    Thus, it sufficies to find an covering of $H^{\alpha}(\mathcal{O},r)$ w.r.t the $\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})}$-norm.
-Following the theory in Edmunds and Triebel [1996], we have
-    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{B_{\infty 1}^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-k}}\right),$
+    $\|\theta_{1}-\theta_{2}\|_{C^{\eta}(\mathcal{O})}\leq c_{\texttt{em}}\cdot\|\theta_{1}-\theta_{2}\|_{B_{\infty 1}^{k}(\mathcal{O})}.$
+    Thus, it sufficies to find an covering of $H^{\alpha}(\mathcal{O},r)$ w.r.t the $\|\cdot\|_{B_{\infty 1}^{k}(\mathcal{O})}$-norm.
+Following the theory in Edmunds and Triebel (1996) (Section 3.3) and Triebel (1978) (Section 4.10.3), respectively, we have
+    $N(H^{\alpha}(\mathcal{O},r),\|\cdot\|_{B_{\infty 1}^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{r}{\rho}\right)^{\frac{d}{\alpha-k}}\right),$
     and hence
-    $N(H^{\alpha}(\mathcal{O},r),\left\|\cdot\right\|_{C^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{c_{\texttt{em}}r}{\rho}\right)^{\frac{d}{\alpha-k}}\right).$
+    $N(H^{\alpha}(\mathcal{O},r),\|\cdot\|_{C^{k}(\mathcal{O})},\rho)\leq\exp\left(c(\alpha,k,d)\cdot\left(\frac{c_{\texttt{em}}r}{\rho}\right)^{\frac{d}{\alpha-k}}\right).$
     The claim follows.
-Lemma § A.10 (Properties of regular link functions).
+Lemma § 4.10 (Properties of Regular Link Functions).
 Let $\Psi$ be a regular link function as in definition 3.1. We have the following properties.
   1. i)
-    Let $\theta_{1},\theta_{2}\in C^{0}(\mathcal{O})$. Then
-    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\infty}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\infty}.$
-  2. i)
-    Let $\theta_{1},\theta_{2}\in C^{1}(\mathcal{O})$. Then
-    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{C^{1}(\mathcal{O})}\leq c(\Psi)\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{C^{1}(\mathcal{O})}.$
-  3. i)
-    Let $\theta_{1},\theta_{2}\in\mathbb{L}^{2}(\mathcal{O})$. Then
-    $\left\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c(\Psi)\cdot\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}.$
-Proof of 13 (Lemma A.10).
-Elementary computations show the claim of all three properties, since by hypothesis $\Psi$ as well as all its derivatives are bounded and hence Lipschitz continuous.
+    Let $\theta_{1},\theta_{2}\in C^{0}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
+    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{\infty}\leq c(\Psi)\cdot\|\theta_{1}-\theta_{2}\|_{\infty}.$
+  2. ii)
+    Let $\theta_{1},\theta_{2}\in C^{1}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
+    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{C^{1}(\mathcal{O})}\leq c(\Psi)\cdot(1+\|\theta_{1}\|_{C^{1}(\mathcal{O})}\lor\|\theta_{2}\|_{C^{1}(\mathcal{O})})\cdot\|\theta_{1}-\theta_{2}\|_{C^{1}(\mathcal{O})}.$
+  3. iii)
+    Let $\theta_{1},\theta_{2}\in\mathbb{L}^{2}(\mathcal{O})$. Then there exists a constant $c=c(\Psi)\in\mathbb{R}_{>0}$, such that
+    $\|\Psi\circ\theta_{1}-\Psi\circ\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c(\Psi)\cdot\|\theta_{1}-\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}.$
+Proof of lemma 4.10.
+Elementary computations show all three statements of the Lemma, since by hypothesis $\Psi$ as well as all its derivatives are bounded and hence Lipschitz continuous. ∎
+Acknowledgements
+The author thanks Richard Nickl and Jan Johannes for their time, patience, and fruitful discussions, as well as for the visit to Richard Nickl’s research group in Cambridge in November 2023, which marked the beginning of this work.
+This work is funded by Deutsche Forschungsgemeinschaft (DFG, German Re- search Foundation) under Germany’s Excellence Strategy EXC-2181/1-39090098 (the Heidelberg STRUCTURES Cluster of Excellence).